### v1

# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

Abstract
Conformal prediction builds marginally valid prediction intervals which cover the unknown outcome of a randomly drawn new test point with a prescribed probability. In practice, a common scenario is that, after seeing the test unit(s), practitioners decide which test unit(s) to focus on in a data-driven manner, and wish to quantify the uncertainty for the focal unit(s). In such cases, marginally valid prediction intervals for these focal units can be misleading due to selection bias. This paper presents a general framework for constructing a prediction set with finite-sample exact coverage conditional on the unit being selected. Its general form works for arbitrary selection rules, and generalizes Mondrian Conformal Prediction to multiple test units and non-equivariant classifiers. We then work out computationally efficient implementation of our framework for a number of realistic selection rules, including top-K selection, optimization-based selection, selection based on conformal p-values, and selection based on properties of preliminary conformal prediction sets. The performance of our methods is demonstrated via applications in drug discovery and health risk prediction.
1 Introduction
Conformal prediction is a versatile framework for quantifying the uncertainty of any black-box prediction model, by issuing a prediction set that covers an unknown outcome with a prescribed probability. Formally, suppose the task is to predict an outcome $Y\in{\mathcal{Y}}$ based on features $X\in{\mathcal{X}}$. Given a set of calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the features of a new test point $X_{n+1}$, conformal prediction builds upon a given prediction model and delivers a prediction set $\widehat{C}_{\alpha,n+1}\subseteq{\mathcal{Y}}$ at level $\alpha\in(0,1)$, which obeys
$\mathbb{P}\big{(}Y_{n+1}\in\widehat{C}_{\alpha,n+1}\big{)}\geq 1-\alpha,$ (1)
as long as $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are exchangeable (e.g., when they are i.i.d. samples). The probability in (1) is over both the calibration data and the test point (Vovk et al.,, 2005; Lei et al.,, 2018).
With this finite-sample, distribution-free guarantee, the conformal prediction set $\widehat{C}_{\alpha,n+1}$ describes a range of plausible values the unknown outcome $Y_{n+1}$ can take, thereby expected to inform downstream decision-making based on any black-box prediction model. With such a promise, methods for issuing marginally-valid – in the sense of (1) – prediction sets have been developed for various problems; see e.g., Angelopoulos and Bates, (2021) for a review.
In many downstream applications, however, people are often only interested in a selective subset of units. For example, practitioners may act upon a unit if it exhibits an interesting property (Olsson et al.,, 2022; Sokol et al.,, 2024; Levitskaya,, 2023); in some cases, practitioners may want to focus on a subset suggested by some complicated deta-dependent decision process such as resource optimization (Svensson et al.,, 2018; Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). It would be misleading if the prediction sets fail to deliver the promised coverage guarantee for the selected units. Let us discuss a few applications where such cases may arise.
  • In drug discovery, an important task is to predict the binding affinity of a drug candidate to a target to inform drug prioritization (Laghuvarapu et al.,, 2024). Scientists may only focus on drug candidates with highest predicted affinites, or selected by an false discovery rate (FDR)-controlling procedure (Jin and Candès, 2023b, ), or whose prediction sets only cover high values (Svensson et al.,, 2017), or that optimize resource usage (Svensson et al.,, 2018). It will be misleading if more than an $\alpha$-fraction of prediction sets for the selected drugs fail to cover the actual binding affinities.
  • In business decision-making, companies may take different inventory decisions based on whether a conformal prediction set suggests a strong demand or a weak demand (Levitskaya,, 2023). Similarly, it will be problematic if the average coverage among strong-demand predictions is below $1-\alpha$.
  • In disease diagnosisOlsson et al., (2022) suggest human intervention if the prediction set for a disease status is too large (this implicitly declares confidence in small prediction sets; see similar ideas in Ren et al., (2023); Sokol et al., (2024)). However, it would be concerning if more than an $\alpha$-fraction of small-sized prediction sets (“approved” as confident) miss the true disease status.
  • In healthcare management, clinical decision makers may send patients to different healthcare categories after optimizing a performance measure (such as waiting time) subject to certain constraints, such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). It is thus natural to only include a data-dependent subset of patients in downstream uncertainty quantification.
In all these cases, it is highly desirable that a prediction set should cover the unknown outcome for a unit of interest with a prescribed probability. This motivates a stronger, selection-conditional guarantee. Supposing there are $m\geq 1$ test units $\{X_{n+j}\}_{j=1}^{m}$, we shall aim for
$\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\big{)}\geq 1-\alpha,$ (2)
where ${\mathcal{D}}_{\textnormal{calib}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is the calibration data, ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ is the test data, and ${\mathcal{S}}(\cdot,\cdot)\subseteq[m]:=\{1,\dots,m\}$ is a data-driven process to decide which units are of interest. In particular, we focus on the case where $\widehat{C}_{\alpha,n+j}$ is constructed after the selection set ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is determined. This is similar to post-selection inference (Lee et al.,, 2016; Markovic et al.,, 2017; Tibshirani et al.,, 2018), yet we study a predictive inference problem; see Section 1.2 for more discussion.
A prediction set with marginal validity (1) does not necessarily cover an unknown outcome conditional on being of interest as (2). Such selection issue has caught recent attention in the literature of predictive inference. Through analysis of a real drug discovery dataset, Jin and Candès, 2023b (, Section 1) demonstrate that more than $30\%$ of seemingly promising prediction sets (in the sense that $\widehat{C}_{\alpha}(X_{n+1})=\{1\}$) miss the actual outcomes for $\alpha=0.01$, i.e., with a marginal coverage of $99\%$. We shall see more examples of this issue in our numerical experiments under various selection processes. Focused on downstream decisions, Jin and Candès, 2023b ; Jin and Candès, 2023a develop methods that select units with large outcomes while guranteeing control of FDR. For uncertainty quantification, Bao et al., (2024) propose methods to construct prediction sets for selected test units with (approximately) false coverage rate (FCR) control. Their methods only achieve exact guarantees for a limited class of selection rules such as top-$K$ selection, while we are to present a general framework that works for arbitrary selection rules. In addition, FCR control and selection-conditional coverage (2) are related but distinct error notions, which we discuss in Section 2.2.
We also note that the selection-conditional coverage (2) may be implied by other stronger notions such as conditional coverage: $\mathbb{P}(Y_{n+1}\in\widehat{C}_{\alpha,n+1}{\,|\,}X_{n+1}=x)\geq 1-\alpha$ for $\mathbb{P}$-almost all $x\in{\mathcal{X}}$. Conditional coverage implies (2) if $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are mutually independent and the selection rule only depends on the features. However, it is not achievable by finite-length prediction intervals in the absence of distributional assumptions (Foygel Barber et al.,, 2021), and the underlying selection rule may be more complicated. In contrast, we may view (2) as a “relaxed” version of conditional coverage that is both achievable and relevant for practical use.
1.1 Overview of contributions
In this paper, we propose a unified framework, named JOint Mondrian Conformal Inference (JOMI), that builds valid prediction sets for focal units picked by any data-dependent selection process. To this end, we make the following contributions.
In Section 3, we introduce the general formulation of JOMI, whose crux is to find a data-dependent “reference” set of calibration data that are exchangeable conditional on the selection event. In this sense, our method can be viewed as a generalization of Mondrian Conformal Prediction (Vovk et al.,, 2005). We prove that for exchangeable data, the prediction set $\widehat{C}_{\alpha,n+j}$ produced by JOMI covers $Y_{n+j}$ with probability at least $1-\alpha$ conditional on a selection event. The selection event can be “test unit $j$ is selected”, which leads to (2). It can also be more granular, such as “test unit $j$ is selected, and there are $k$ selected test units”. Our framework works for any selection rule ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that is permutation-invariant to ${\mathcal{D}}_{\textnormal{calib}}$, without any modeling assumptions on the data generating process.
In Section 4, we work out efficient implementation of JOMI for a number of selection processes that may be of practical interest, including:
  • Covariate-dependent selection. When the selection rule do not involve $\{Y_{i}\}_{i=1}^{n}$, the computation complexity of our method is (at most) $O(mn)$ times that of the selection rule. This imples efficient implementation for a wide range of problems, including various forms of top-K selection and selecting based on complicated constrained optimization programs. As a special case, we recover the method of Bao et al., (2024) for top-K selection among test data, and yield valid solutions to other ranking-based selection methods attempted in their work.
  • Conformal p-value-based selection. We also derive efficient implementation for selection rules based on conformal p-values with both fixed and data-dependent thresholds. For instance, our prediction set is finite-sample exactly valid for units selected by Conformalized Selection (Jin and Candès, 2023b, ), which is studied in Bao et al., (2024) with approximate guarantee.
  • Selection based on preliminary conformal prediction set. In addition, we derive a general, efficient instantiation when unit(s) whose marginal prediction sets demonstrate an interesting property are selected, such as being of a short length, or having a lower bound above a threshold. Our methods can be useful for re-calibrating the uncertainty quantification of such seemingly promising units.
Finally, we demonstrate the applications of our methods via a few realistic selection rules that may occur in drug discovery (Section 5) and health risk prediction (Section 6). Our results show that marginally valid prediction sets may undercover or overcover, while our methods lead to valid coverage.
1.2 Related work
Our methods extend conformal prediction, which is a general framework for building marginally valid prediction intervals (Vovk et al.,, 2005; Lei et al.,, 2018; Romano et al.,, 2019, 2020). Compared with vanilla conformal prediction, we aim at a new goal, which is to achieve selection-conditional coverage for a unit of interest, instead of marginally for an “average” unit. Particularly related to our framework is Mondrian Conformal Prediction (Vovk et al.,, 2003, MCP), which uses a subset of calibration data to build a prediction interval for coverage conditional on a pre-determined or permutation-invariant class membership. However, achieving selection-conditional coverage is a new goal with distinct motivations, and our techniques and scopes of application are significantly different (see Remark 1).
The selection-conditional guarantee we seek for is closely related to post-selection inference (POSI) literature (Berk et al.,, 2013; Tibshirani et al.,, 2016; Chernozhukov et al.,, 2015) and in particular conditional selective inference (Lee et al.,, 2016; Markovic et al.,, 2017; Reid et al.,, 2017; Tibshirani et al.,, 2018; Kivaranovic and Leeb,, 2020; McCloskey,, 2024). In POSI, one important task is to build confidence intervals for a selected subset of model parameters conditional on the selection event, which is close to our goal of selection-conditional coverage guarantee. Methods for this goal usually leverage specific problem structure such as linearity and distribution of the estimators to derive the confidence intervals (Lee et al.,, 2016; Tian and Taylor,, 2017; Andrews et al.,, 2024; Liu,, 2023). As we focus on predictive inference (we have units instead of parameters), we leverage the exchangeability across units to achieve valid selection-conditional coverage, hence our methods are very different.
Another related error notion in the selective inference literature is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005); we compare them in more details in Section 2.2. Methods for achieving FCR control have been proposed for various settings (Benjamini and Yekutieli,, 2005; Weinstein et al.,, 2013; Weinstein and Ramdas,, 2020; Zhao and Cui,, 2020). For conformal prediction, Bao et al., (2024) propose prediction sets for selected units while achieving (approximate) FCR control. However, their methods only work for multiple test units. In addition, they achieve exact guarantee only for selection rules that are invariant to permutations of calibration and test units and top-$K$ selection (where $K$ test units with the highest predictions are selected); we will show that when applying JOMI to these problems, we yield the same prediction sets as theirs. Meanwhile, our framework leads to exact guarantee for some other problems studied in Bao et al., (2024) with approximate FCR control. Our framework also tackles many more settings with complicated selection rules that cannot be handled by theirs.
Selective inference issues in conformal prediction or predictive inference have caught recent attention. Most of existing works study multiple testing, i.e., selecting individuals that obey some properties with false discovery rate (FDR) control, including selecting outliers (Bates et al.,, 2023; Marandon et al.,, 2022; Liang et al.,, 2022) and selecting high-quality data (Jin and Candès, 2023b, ; Jin and Candès, 2023a, ). The construction of conformal prediction sets is an orthogonal direction, and our methods also work for selection based on such procedures (we develop an efficient method in Section 4.2.2).
Under a similar name, Sarkar and Kuchibhotla, (2023) proposed “post-selection inference for conformal prediction”, but they mainly focus on selecting the confidence level $\alpha\in(0,1)$ instead of selecting the units. In addition, they fall within the “simultaneous inference” strand in POSI so that their inference is valid simultaneously for all confidence levels, other than conditional on a selection event. For these reasons, our methods and guarantees are quite different.
2 Problem setup
We are interested in inferring the unknown outcome $Y$ of certain “focal” units selected in a data-driven manner. Following the split conformal prediction framework (Vovk et al.,, 2005; Lei et al.,, 2018), our prediction set is based on a fitted prediction model obtained from a training set ${\mathcal{D}}_{\textnormal{train}}$. Recall the calibration data ${\mathcal{D}}_{\textnormal{calib}}=\{Z_{i}\}_{i=1}^{n}$, $Z_{i}=(X_{i},Y_{i})$, and test points $\{Z_{n+j}\}_{j=1}^{m}$, $Z_{n+j}=(X_{n+j},Y_{n+j})$. The outcomes $\{Y_{n+j}\}_{j=1}^{m}$ are unobserved and of interest. We write ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ for brevity. We assume ${\mathcal{D}}_{\textnormal{train}}$ is independent of ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, and we condition on ${\mathcal{D}}_{\textnormal{train}}$, treating the fitted models as fixed.
2.1 Selection-conditional coverage
We call $Z_{n+j}$ a focal unit if $j\in\widehat{\mathcal{S}}$, where $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\subseteq[m]$ is obtained from a selection rule ${\mathcal{S}}$ that may depend on the observe data. Its potential dependence on ${\mathcal{D}}_{\textnormal{train}}$ is clear since we treat ${\mathcal{D}}_{\textnormal{train}}$ as fixed. Without loss of generality, we posit that ${\mathcal{S}}$ is a determinististic function; when the selection rule is randomized, one can condition on its randomness and follow our framework.
For each test unit $j$, we wish to construct a prediction set $\widehat{C}_{\alpha,n+j}\subseteq{\mathcal{Y}}$ such that
$\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\geq 1-\alpha,$ (3)
where $\alpha\in(0,1)$ is the confidence level, and $\mathfrak{S}\subseteq 2^{[m]}$ is some pre-specified collection of subsets of $[m]$. We call $\mathfrak{S}$ the selection taxonomy in what follows.
Equation (3) is the most general form of the selection-conditional coverage we aim for in this paper, where difference choices of $\mathfrak{S}$ leads to a spectrum of granularity in the conditioning event. For instance, taking $\mathfrak{S}=\{S_{0}\}$ for some $S_{0}\subseteq[m]$ achieves coverage conditional on selecting a specific set $S_{0}$; this is similar to the coverage guarantee in (Lee et al.,, 2016) for high-dimensional inference. Taking $\mathfrak{S}=\{S\subseteq[m]\colon|S|=r\}$ for some $r\leq m$ achieves coverage conditional on getting a specific number of selected units. Finally, taking $\mathfrak{S}=2^{[m]}$ puts no restrictions on the selection set, giving rise to the guarantee (2) introduced in the beginning, coverage conditional only on a unit being selected.
2.2 Relations among notions of selective coverage
Before introducing our methods, we take a moment here to compare different notions of selective coverage. Readers who are more interested in our methodology may skip the remaining of this section.
Our first observation is that (3) implies (2) under appropriate conditions. The proof of the next Proposition is in Appendix B.1.
Proposition 1.
If $\widehat{C}_{\alpha,n+j}$ satisfies (3) for a set of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, then it satisfies (2).
To distinguish the selection-conditional coverage in (3) and that in (2), we refer to the former as strong selection-conditional coverage and the latter as weak selection-conditional coverage.
Another widely used post-selection-type guarantee is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), defined as the expected proportion of selected units missed by the prediction set:
$\text{FCR}\,:=\,\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]},$ (4)
where $a\vee b=\max\{a,b\}$ for any $a,b\in\mathbb{R}$. As introduced earlier, Bao et al., (2024) mainly focuses on constructing prediction sets with FCR control. However, with $m=1$, it holds that $\text{FCR}=\mathbb{P}(j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j}))\leq\alpha$ for any marginally valid prediction set. Thus, the FCR is not very meaningful when we consider one test point.
We claim that the strong selection-conditional coverage implies FCR control with a proper choice of the selection taxonomy. The proof of the following proposition can be found in Appendix B.2.
Proposition 2.
[…]
As a side note, we find that the weak selection-conditional coverage does not necessarily imply FCR control, although in some special cases both can be true (see, e.g., Bao et al., (2024) for such examples). We put this as a proposition below, with a counterexample given in Appendix B.3.
Proposition 3.
There exists an instance and prediction sets $\widehat{C}_{\alpha}$ such that $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage at level $\alpha$ but violates the FCR control at level $\alpha$.
We close this part by introducing several notations that will be used throughout the paper.
Notations.
We denote the index sets ${\mathcal{I}}_{{\textnormal{calib}}}=\{1,\dots,n\}$ and ${\mathcal{I}}_{\textnormal{test}}=\{n+1,\dots,n+m\}$. Let $S:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ be the score function for selection, and $V:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ the nonconformity score function for constructing the prediction set. For example, $S(x,y)=y$ and $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$. For any $j\in{\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, define $S_{i}=S(X_{i},Y_{i})$ and $V_{j}=V(X_{j},Y_{j})$.
Let ${\mathcal{D}}_{j}={\mathcal{D}}_{\textnormal{calib}}\cup\{Z_{n+j}\}$, and ${\mathcal{I}}_{j}={\mathcal{I}}_{\textnormal{calib}}\cup\{n+j\}$ be the corresponding indices; define also ${\mathcal{D}}^{c}_{j}={\mathcal{D}}_{\textnormal{test}}\backslash\{Z_{n+j}\}$ and ${\mathcal{I}}^{c}_{j}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$. For any $j\in[m]$, we define $Z_{n+j}(y)=(X_{n+j},y)$, the value of $Z_{n+j}$ if $Y_{n+j}$ were to take a hypothesized value $y\in{\mathcal{Y}}$.
3 JOMI: A unified framework
3.1 Warm-up: split conformal prediction
To warm up, we briefly summarize the split conformal prediction (SCP) method (Vovk et al.,, 2005; Lei et al.,, 2018) and the ideas for its finite-sample coverage due to exchangeability.
We begin with a nonconformity score function $V\colon{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ determined by the training set, so that $V(x,y)$ informs how well a hypothetical value $y\in{\mathcal{Y}}$ conforms to a machine prediction. For instance, one may set $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$. Other popular choices include CQR (Romano et al.,, 2019) for regression and APS (Romano et al.,, 2020) for classification. We compute $V_{i}=V(X_{i},Y_{i})$ for $i\in[n]$. The split conformal prediction set for a test point $j\in[m]$ is then
$\widehat{C}_{\alpha,n+j}^{\text{SCP}}=\Big{\{}y\colon V(X_{n+j},y)\leq\text{Quantile}\big{(}1-\alpha;\{V_{i}\}_{i=1}^{n}\cup\{+\infty\}\big{)}\Big{\}},$ (5)
where $\text{Quantile}(1-\alpha;\cdot)$ is the empirical $(1-\alpha)$-th quantile of the set in the second argument. When $(Z_{1},\dots,Z_{n},Z_{n+j})$ are exchangeable, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ achieves valid marginal coverage (1(Vovk et al.,, 2005).
It helps motivate our approach to first introduce the ideas behind the validity of $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$. Note that
$\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\text{SCP}}\big{)}=\mathbb{P}\Big{(}V_{n+j}\leq\text{Quantile}\big{(}1-\alpha;\{V_{i}\}_{i=1}^{n}\cup\{V_{n+j}\}\big{)}\Big{)},$ (6)
where $V_{n+j}=V(X_{n+j},Y_{n+j})$. Define the unordered set $[Z]_{j}=[Z_{1},\dots,Z_{n},Z_{n+j}]$. Due to exchangeability, conditional on the event $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, the probability of $Z_{n+j}$ taking on each value in $z_{1},\dots,z_{n+j}$ is equal. The right-handed side of (6) is thus lower bounded by $1-\alpha$.
The key challenge of achieving selection-conditional coverage like (3) is that, conditional on the selection event, the data points $\{Z_{1},\dots,Z_{n},Z_{n+j}\}$ are not exchangeable any more. More specifically, given selection, and further conditional on $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, it is unclear how $Z_{n+j}$ is distributed over $z_{1},\dots,z_{n+j}$. As such, a natural idea is to find a subset of calibration data that are “exchangeable” with the new test point even given the selection event. We introduce our methods in the next.
3.2 Conformal inference via locally exchangeable subgroups
Fix a test unit $j\in[m]$. Recall that $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the focal set, and $Z_{n+j}(y)=(X_{n+j},y)$ is the imputed data point with a hypothesized response $y\in{\mathcal{Y}}$. For any calibration data $i\in{\mathcal{I}}_{\textnormal{calib}}$, we define the “swapped” calibration data ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)$ and test data ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}(y)$ to be
${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)=(Z_{1}^{{\textnormal{swap}(i,j)}}(y),Z_{2}^{{\textnormal{swap}(i,j)}}(y),\dots,Z_{n}^{{\textnormal{swap}(i,j)}}(y)),$
${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}=(X_{n+1}^{{\textnormal{swap}(i,j)}},X_{n+2}^{{\textnormal{swap}(i,j)}},\dots,X_{n+m}^{{\textnormal{swap}(i,j)}}),$
where
$Z_{\ell}^{{\textnormal{swap}(i,j)}}(y)=\begin{cases}Z_{n+j}(y)&\ell=i,\\ Z_{\ell}&\ell\neq i.\end{cases}\quad X_{n+\ell}^{{\textnormal{swap}(i,j)}}=\begin{cases}X_{i}&\ell=j,\\ X_{n+\ell}&\ell\neq j.\end{cases}$
[…]
Figure 1: Graphical illustration of ${\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}(y)$ and ${\mathcal{D}}_{\textnormal{test}}^{{\textnormal{swap}(i,j)}}$.
That is, they are the calibration and test data if we treat $Z_{i}$ as the $j$-th test point, and $Z_{n+j}(y)$ as the $i$-th calibration point. Figure 1 is an illustration of the swapping operation. Applying the same selection rule, the selection set after swapping and with the hypothesized $y$ is then
$\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y)={\mathcal{S}}\big{(}{\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y),{\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}\big{)}.$
Then, we define the “reference set” for achieving (3) with taxonomy $\mathfrak{S}$ as
^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}\colon j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y),\text{ and }\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}\big{\}}.$ (7)
In words, the reference set is the collection of calibration points $i$ such that, after swapping unit $i$ and $n+j$, the test point $j$ remains in the focal set and the focal set remains in $\mathfrak{S}$. We use the notation ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ to emphasize that ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is a data-dependent mapping from ${\mathcal{Y}}$ to subsets of ${\mathcal{I}}_{{\textnormal{calib}}}$.
With all the preparation, we define our prediction set via
$\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\cup\{+\infty\}\big{)}\Big{\}}.$ (8)
As we shall show shortly, our prediction set (8) achieves near-exact coverage when the nonconformity score is continuous and the reference set is of moderate size. In fact, it is possible to achieve exact coverage by introducing extra randomness:
$\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}$
=$\quad=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}(y)|}\leq 1-\alpha\bigg{\}},$ (9)
where $U_{1},\ldots,U_{m}$ are i.i.d. random variables drawn from $\textrm{Unif}[0,1]$ which are independent of the data.
Remark 1 (Connection to MCP).
Our framework is closely related to Mondrian Conformal Prediction (MCP(Vovk et al.,, 2003) which applies to one test unit ($m=1$). Given $K$ categories $[K]$, they assume access to a classifier ${\mathcal{K}}$ that gives the class label for all units: $(\kappa_{1},\kappa_{2},\dots,\kappa_{n+1})={\mathcal{K}}(Z_{1},Z_{2},\dots,Z_{n+1}),$ where $\kappa_{i}\in[K]$ is the class label for unit $i$. The classifier ${\mathcal{K}}$ is required to be invariant to any permutation $\pi$ of $[n+1]$, i.e., $(\kappa_{\pi(1)},\kappa_{\pi(2)},\dots,\kappa_{\pi(n+1)})={\mathcal{K}}(Z_{\pi(1)},Z_{\pi(2)},\dots,Z_{\pi(n+1)}).$ The goal of MCP is to construct a prediction set for the test unit with cateogory-wise coverage, i.e., $\mathbb{P}(Y_{n+1}\in\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}{\,|\,}\kappa_{n+1}=k)$ for each $k\in[K]$. This is achieved by finding a subset of calibration data for each hypothesized $y\in{\mathcal{Y}}$,
~$\widetilde{{\mathcal{R}}}(y)=\{i\in{\mathcal{I}}_{\textnormal{calib}}:{\mathcal{K}}({\mathcal{D}}_{{\textnormal{calib}}},(X_{n+1},y))_{[i]}=k\},$ (10)
and then constructing the prediction set as
$\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+1},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widetilde{{\mathcal{R}}}(y)}\cup\{\infty\}\big{)}\Big{\}}.$
While the setups seem different, JOMI effectively reduces to MCP when ${\mathcal{K}}$ is a binary classifier and ${\mathcal{S}}$ is equivariant under permutations of $[n+1]$. To see this, one can view the selection rule ${\mathcal{S}}$ as categorizing test points into two classes (selected and not selected), and selection-conditional coverage (3) as coverage given the “selected” category. In this way, one can check that our reference set ^$\widehat{{\mathcal{R}}}_{n+1}(y)$ coincides with ~$\widetilde{{\mathcal{R}}}(y)$ defined in (10), and thus the prediction sets also coincide.
With a distinct goal of selection-conditional coverage, we additionally address the challenge that realistic selection rules are often asymmetric to calibration and (multiple) test points. To this end, JOMI generalizes MCP via a more delicate but still general framework for finding exchangeable subgroups. As we jointly consider the calibration and test points when defining the reference set, we call our method JOint Mondrain Conformal Inference (JOMI).
3.3 Theoretical guarantee
Theorem 1 confirms the validity of our prediction sets (8) and (3.2).
Theorem 1.
Suppose ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is invariant to permutations of ${\mathcal{D}}_{\textnormal{calib}}$. Suppose $\{Z_{i}\}_{i=1}^{n}\cup\{Z_{n+j}\}$ are exchangeable conditional on $\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$ for any $j\in[m]$. Then, for any selection taxonomy $\mathfrak{S}$,
  1. (a)
    $\widehat{C}_{\alpha,n+j}$ defined in (8) obeys
    $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\geq 1-\alpha.$ (11)
    Furthermore, if ties between $V_{1},\ldots,V_{n},V_{n+j}$ occur with probability zero, then
    $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\leq 1-\alpha+\mathbb{E}\bigg{[}\frac{1}{1+|\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})|}\bigg{]}.$ (12)
  2. (b)
    The randomized prediction set $\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}$ defined in (3.2) satisfies
    $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S})=1-\alpha.$
We defer the detailed proof of Theorem 1 to Appendix B.4. Let us provide some intuitions for the validity results here. Similar to the ideas of SCP in Section 3.1, the key fact we rely on is that, plugging in the true value $y=Y_{n+j}$, data in the reference set and the new test point are still exchangeable conditional on the event that $j$ is selected. To be specific, imagine that we know the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]=[z_{1},\dots,z_{n},z_{n+j}]$ and the values of the remaining test points $\{X_{n+\ell}\}_{\ell\neq j}$. The only randomness is thus in the ordering of $Z_{1},\dots,Z_{n+j}$. Meanwhile, our reference set ^$\widehat{\mathcal{R}}_{j}(\cdot)$ is constructed in a delicate way such that the (unordered) scores in the reference set $[V_{i}\colon i\in\widehat{\mathcal{R}}_{j}(Y_{n+j})]$ is fully determined by the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]$ if ${\mathcal{S}}(\cdot,\cdot)$ is invariant to permutations of the “calibration” data in the first argument. Then, by the exchangeability of $Z_{1},\dots,Z_{n},Z_{n+j}$, the probability of $Y_{n+j}$ taking on any value in $\widehat{R}$ is equal given the unordered set, leading to the validity in Theorem 1.
4 Computationally tractable instances
Section 3 provides a general framework for constructing valid prediction sets conditional on selection events, but computing the prediction sets according to their definitions requires looping over all possible values of $y$, often a computationally intractable task.
In this section, we instantiate the generic procedure through concrete examples, in which special structures allow us to efficiently compute the prediction sets (8) and (3.2). We focus on three classes of selection rules: the selection rules using solely the covariates, those based on conformal p-values, and those based on conformal prediction sets.
4.1 Covariate-dependent selection rules
When the selection rule depends solely on the covariates, the reference sets can be efficiently computed by definition, since ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ no longer depends on $y$ (we shall also suppress the dependence of ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ and $S(x,y)$ on $y$ throughout Section 4.1). The complete procedure for constructing $\widehat{C}_{\alpha,n+j}$’s and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ with an arbitrary covariate-dependent selection rule is summarized in Algorithm 1, and the computational complexity the procedure is $O(mn\cdot C_{\mathcal{S}})$, with $C_{\mathcal{S}}$ being the complexity of the selection process.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$;
[…]
for $j=1,\ldots,m$ do
       /* Constructing reference sets */
       Initialize ^$\widehat{{\mathcal{R}}}_{n+j}=\emptyset$;
       for $i=1,\ldots,n$ do
             ^$\widehat{{\mathcal{R}}}_{n+j}\leftarrow\widehat{{\mathcal{R}}}_{n+j}\cup\{i\}$   if $j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}$ and $\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}$;
       end for
       /* Constructing prediction sets */
       if form = dtm then
             $\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}}\cup\{+\infty\}\big{)}\Big{\}}.$
       end if
      if form = rand then
             Sample $U_{j}\sim\textnormal{Unif}[0,1]$;
             $\widehat{C}_{\alpha,n+j}=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}|}\leq 1-\alpha\bigg{\}}.$
       end if
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 1 JOMI for arbitrary covariate-dependent selection rules
By Theorem 1, the output of Algorithm 1 is valid as long as the selection rule is invariant to the permutation of the calibration points. This includes many commonly-used selection rules:
  1. (1)
    Top-K selection: the $K$ test units with the highest scores $S(X_{i})$ are selected.
  2. (2)
    Selection based on joint quantiles: unit $i$ is selected if its score $S(X_{i})$ is surpasses the $q$-th quantile of both the calibration and test scores.
  3. (3)
    Selection based on calibration quantiles: a test unit $i$ is selected if its score $S(X_{i})$ surpasses the $q$-th quantile of calibration scores.
  4. (4)
    Selection with constraints: each unit $i$ has a score $S(X_{i})$ and a cost $C(X_{i})$. The selection rule amounts to finding a subset $\widehat{{\mathcal{S}}}$ of test units that maximizes $\sum_{j\in\widehat{{\mathcal{S}}}}S(X_{n+j})$ subject to $\sum_{j\in\widehat{{\mathcal{S}}}}C(X_{n+j})\leq C$, where $C$ reflects the total budget.
As we shall show shortly, it turns out that for rules (1)-(3) we can further improve the computational efficiency by explicitly writing down the reference sets.
We also note that rules (1)-(2) were considered in Bao et al., (2024), in which the authors proposed methods achieving the weak selection-conditional coverage in (2) and FCR. The prediction sets proposed therein coincide with ${\mathcal{C}}_{\alpha,n+j}$, and we shall show that they in fact achieve the strong selection-conditional coverage in (3) for free. We provide specially tailored solutions for selection rules (1)-(3) below. Throughout, we let ${\mathcal{T}}=\{S_{i}\}_{i\in{\mathcal{I}}_{\textnormal{calib}}\cup{\mathcal{I}}_{\textnormal{test}}}$.
4.1.1 Top-K selection
In this example, we assume that $S_{n+j}$’s are distinct for all $j\in[m]$ almost surely (otherwise it is unclear how to define the “top $K$” elements). Under the top-K selection rule, the $K$ test units with the highest scores are selected. We can write the top-K selection set as
$\widehat{{\mathcal{S}}}_{\textnormal{topk}}=\{j\in[m]:S_{n+j}>T_{\textnormal{topk}}\},\quad\text{where}\quad T_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{j\in[m]}}\mathds{1}\{S_{n+j}\leq t\}\geq m-K\Big{\}}.$
For this selection rule, the reference set of each $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ takes a unified form:
^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{topk}}\}.$ (13)
Intuitively, $T_{\textnormal{topk}}$ refers to the $(m-K)$-th smallest score among the test units; when $S_{n+j}$ is among the top $K$, replacing $S_{n+j}$ with another score that is greater than $T_{\textnormal{topk}}$ does not change the selection outcome. This intuition is formalized in the following proposition, with the proof deferred to Appendix B.5.
Proposition 4.
For the top-K selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}\in 2^{[m]}$, such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{topk}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ for all $y\in{\mathcal{Y}}$.
Computing ^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ as a universal reference set reduces the computation complexity of the reference construction step in Algorithm 1 to $O(\max(m,n))$. It is also worth mentioning that the definition of reference set here is agnostic to the choice of $\mathfrak{S}$.
4.1.2 Selection based on joint quantiles
In some cases, practitioners may want to select a test unit $n+j$ if its score $S_{n+j}$ greater than the $q$-th quantile of the calibration and test scores. Equivalently, we could write the selection set as
$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\{j\in[m]:S_{n+j}>T_{\textnormal{jq}}\},\quad\text{where}\quad T_{\textnormal{jq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{m+n}}\mathds{1}\{S_{i}\leq t\}\geq q(n+m)\Big{\}}$
By definition, the selection threshold $T_{\textnormal{jq}}$ is invariant to the permutation of the calibration and test scores, and it is straightforward to check that for all $j\in\widehat{{\mathcal{S}}}$, the reference set is in the form of
^$\widehat{{\mathcal{R}}}_{\textnormal{jq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{jq}}\}.$
Using ^$\widehat{{\mathcal{R}}}_{jq}$ reduces the computation complexity of the reference set in Algorithm 1 to $O(\max(m,n))$.
4.1.3 Selection based on calibration quantiles
Consider now a selection rule, where test unit $n+j$ is selected if its score $S_{n+j}$ is greater than the $q$-th quantile of of the calibration scores. Equivalently, we could write the selection set as
$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\big{\{}j\in[m]:S_{n+j}>T_{\textnormal{cq}}\big{\}},\text{ where }T_{\textnormal{cq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{n}}\mathds{1}\{S_{i}\leq t\}\geq qn\Big{\}}.$ (14)
For each $j\in\widehat{{\mathcal{S}}}$, the refernence set consists of all calibration units in the top $q$-th quantile, i.e.,
^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{cq}}\}.$ (15)
The validity of ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ as the referenece set is established below, whose proof is deferred to appendix B.6.
Proposition 5.
For the selection rule (14), for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$, such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\widehat{{\mathcal{R}}}_{n+j}(y)$ defined in (7) for all $y\in{\mathcal{Y}}$.
Similar to the previous two cases, directly computing ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ reduces the computation complexity of the reference set to $O(\max(m,n))$. As a side note, Bao et al., (2024) also considered the selection rule based on calibration quantiles (as an instance of “calibration-assisted selection”), for which they provide asymptoically FCR control guarantees under additional assumptions.
4.2 Selection based on conformal p-values
Suppose we wish to select the test units whose outcomes satisfy certain conditions. For example, Jin and Candès, 2023b studies the problem of selecting test units with large outcomes, where they formulate the task as hypothesis testing problems. Following their formulation, we consider for each $j\in[m]$ the following null and alternative hypotheses:
$H_{j,0}:Y_{n+j}\leq c_{n+j}\text{ v.s. }H_{j,1}:Y_{n+j}>c_{n+j},$ (16)
where $c_{n+j}$ is a (potentially random) threshold. When $H_{j,0}$ is rejected, we select the $j$-th test unit. For testing $H_{j,0}$in and Candès, 2023b devise the conformal p-value $p_{j}$ such that
$\mathbb{P}(p_{j}\leq t,Y_{n+j}\leq c_{j})\leq t,\text{ for any }t\in[0,1].$
In this paper, we follow their strategy, but work with a slightly different form of the conformal p-value that is uniformly more powerful than the original version in Jin and Candès, 2023b . Throughout Section 4.2, we assume the score function $S(x,y)$ to be monotonically nonincreasing in $y$, e.g., $s(x,y)=\widehat{\mu}(x)-y$. For each $i\in{\mathcal{I}}_{\textnormal{calib}}\cup{\mathcal{I}}_{\textnormal{test}}$, define $\widehat{S}_{i}=S(X_{i},c_{i})$, and for any $j\in[m]$, we define the conformal p-value as
$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}}{n+1}.$ (17)
We provide a proof for the validity of $p_{j}$ for testing $H_{j,0}$, as well as its comparison with Jin and Candès, 2023b in Appendix C.1 for completeness.
With these p-values, we can select the units whose conformal p-values are smaller than a threshold. In terms of the threshold choice, we consider two scenarios:
  1. 1.
    Fixed threshold: select test units whose conformal p-values (17) are below a fixed threshold $q\in(0,1)$, such as Bonferroni correction.
  2. 2.
    Data-driven threshold: select test units whose conformal p-values (17) are below a data-dependent threshold given by the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995).
4.2.1 Fixed threshold
We first consider $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:p_{j}\leq q\}$ with a threshold $q\in(0,1)$. We will see that there are only two potential forms of the reference set ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ when varying $y\in\mathbb{R}$. For $k\in\{0,1\}$, define
^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,0)\}\in\mathfrak{S}\Big{\}}$
$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (18)
where we recall that $\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}$ is the score in the ($n+\ell$)-th position after swapping $(i,j)$. For $k,\ell\in\{0,1\}$,
$T^{\textnormal{fixed}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}\leq q(n+1)\bigg{\}}.$ (19)
Proposition 6 presents the simplified expression for ^$\widehat{\mathcal{R}}_{n+j}(y)$ for any selected unit and any selection taxonomy, whose proof is deferred to Appendix B.7.
Proposition 6.
Suppose the selection set contains the units whose conformal p-values (17) are below a fixed value $q\in(0,1)$, i.e., $\widehat{{\mathcal{S}}}=\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]\colon p_{j}\leq q\}$. Then for any selection taxonomy $\mathfrak{S}$ and any $j\in[m]$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},0}_{n+j}$ when $y>c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},1}_{n+j}$ when $y\leq c_{n+j}$, for any $y\in\mathbb{R}$.
Finally, the JOMI prediction interval is therefore
$\widehat{{\mathcal{C}}}_{\alpha}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\big{\}}\cup\big{\{}y\in{\mathcal{Y}}:y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\big{\}},$
where $\widehat{q}_{k}={\textnormal{Quantile}}(\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed},k}_{n+j}}\cup\{\infty\};1-\alpha)$ for $k=0,1$. See Algorithm LABEL:alg:pval_fix for a summary.
The computation complexity of Algorithm LABEL:alg:pval_fix is polynomial in $n$ and $m$. Note that for each $j$, computing $T_{j}^{\textnormal{fixed}}(k,\ell)$ costs $O(m)$, computing ^$\widehat{\mathcal{R}}_{n+j}^{\textnormal{fixed},k}$ costs $O(mn)$, and computing $\widehat{C}_{\alpha,n+j}$ costs $O(n)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $T_{j}^{\textnormal{fixed}}(k,\ell)$ as in (19) for $(k,\ell)\in\{0,1\}$. Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},1}$ as (4.2.1). Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\text{fixed},k}}\cup\{\infty\})$, $k=0,1$. Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 2 JOMI for conformal p-values with fixed threshold
4.2.2 Data-driven p-value threshold
According to Jin and Candès, 2023b , applying the BH procedure (Benjamini and Hochberg,, 1995) to the p-values $\{p_{j}\}_{j=1}^{m}$ yields a selection set with controlled false discovery rate (FDR). Formally, let $p_{(1)}\leq p_{(2)}\leq\ldots\leq p_{(m)}$ be the order statistics of the conformal p-values. The BH selection set $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ is
$\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\colon p_{j}\leq k^{*}q/m\},\quad k^{*}=\max\Big{\{}k:p_{(k)}\leq kq/m\Big{\}},$
with the convention that $\max\varnothing=0$.
For the BH selection rule, the reference set can also be simplified to two cases. For $k=0,1$, define
^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,0)\}\in\mathfrak{S}\Big{\}}$
$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (20)
and for $k,\ell=0,1$, we define
$T^{\textnormal{BH}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}}{1+\sum_{j\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{(k,\ell)}\}}\cdot\frac{m}{n+1}\leq q\bigg{\}}.$ (21)
We prove the equivalence in the next proposition, deferring the proof to Appendix B.8.
Proposition 7.
For the BH selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},1}$ when $y\leq c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},0}$ when $y>c_{n+j}$ as defined in (4.2.2) for any $y\in\mathbb{R}$.
Therefore, the JOMI prediction set is of a similar form as the fixed-threshold case. We summarize the entire workflow in Algorithm 3. Similarly, the computation cost of $\widehat{C}_{\alpha,n+j}$ is $O(mn)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $T_{j}^{\textnormal{BH}}(k,\ell)$ as in (21) for $(k,\ell)\in\{0,1\}$. Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},1}$ as (4.2.2). Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\textnormal{BH},k}}\cup\{\infty\})$, $k=0,1$. Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 3 JOMI for conformal p-values with BH-threshold
4.3 Selection based on conformal prediction sets
Practitioners may also select test unit(s) based on the property of a preliminary prediction set. For example, they may select units whose prediction intervals are shorter than a threshold (Sokol et al.,, 2024), or whose prediction sets that entirely lie above a threshold. Note that the original prediction intervals are no longer valid conditional on being selected (Jin and Candès, 2023b, ). In this section, we instantiate our general framework to re-calibrate prediction sets for the units selected in such a way.
We thus consider two stages of prediction set construction: the one constructed in the first stage, which we call the preliminary prediction set, is used for determining the selection set $\widehat{\mathcal{S}}$, while that in the second stage, which we call the selective prediction set, is the one we are to build with JOMI.
Recall the SCP framework in Section 3.1. Let $S(x,y)$ and $V(x,y)$ be the nonconformity score for the two stages, respectively. The level $1-\beta$ preliminary prediction set for the $j$-th test unit is
$\widehat{C}^{\text{prelim}}_{\beta}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:S(X_{n+j},y)\leq\eta\big{\}},$ (22)
where $\eta$ is the $K\,:=\,\lceil(1-\beta)(n+1)\rceil$-th smallest element in $\{S_{i}\}_{i=1}^{n}$.
By (22), the preliminary prediction set is fully determined by $X_{n+j}$ and $\eta$. We can thus consider a general selection rule ${\mathcal{L}}:{\mathcal{X}}\times\mathbb{R}\mapsto\{0,1\}$, where ${\mathcal{L}}(X_{n+j},\eta)=1$ means selecting the unit. For instance, selecting prediction intervals shorter than a threshold $\lambda$ gives ${\mathcal{L}}(x,\eta)=\mathds{1}\{\widehat{q}_{U}(x)-\widehat{q}_{L}(x)+2\eta\leq\lambda\}$ if we use CQR for $\widehat{C}_{\beta}^{\text{prelim}}$ (Romano et al.,, 2019), i.e., setting $S(x,y)=\max\{\widehat{q}_{L}(x)-y,y-\widehat{q}_{U}(x)\}$, where $\widehat{q}_{L}(x)$ and $\widehat{q}_{U}(x)$ are estimates of some lower and upper conditional quantiles. As another example, for binary $Y$, we might want to select the units whose prediction set is a singleton, then the selection rule can be written as ${\mathcal{L}}(x,\eta)=\mathds{1}\big{\{}S(x,1)\leq\eta<S(x,0)\text{ or }S(x,0)\leq\eta<S(x,1)\big{\}}$.
The selection set for a selection rule ${\mathcal{L}}$ is
$\widehat{{\mathcal{S}}}_{\textnormal{ps}}=\big{\{}j\in[m]:{\mathcal{L}}(X_{n+j},\eta)=1\big{\}}.$
We are to derive a general form for a slightly conservative version of the JOMI prediction set, which nevertheless has tight coverage in all our numerical experiments (see Section 6 for examples). We only present the results for $\mathfrak{S}=2^{[m]}$ for the ease of exposition. Formally, we define
$\widehat{C}^{\textnormal{ps}}_{\alpha,n+j}:=\leavevmode\nobreak$$\{y:\eta^{-}\leq S(X_{n+1},y)\leq\eta^{+}\}$ (23)
$\cup\{y:V(X_{n+j},y)\leq q_{1},S(X_{n+1},y)<\eta^{-})\}$
$\cup\{y:V(X_{n+j},y)\leq q_{2},S(X_{n+1},y)>\eta^{+})\}$
where $\eta^{+}$ and $\eta^{-}$ are the $(K+1)$-th and $(K-1)$-th smallest element in $\{S_{i}\}_{i=1}^{n}$, respectively, and
$q_{1}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta^{-},\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta^{-})=1\big{\}}\Big{)};$
$q_{2}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta,{\mathcal{L}}(X_{i}.\eta^{+})=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta,\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta)=1\big{\}}\Big{)}.$ (24)
We prove the validity of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ in (23) below, whose proof is deferred to Appendix B.9.
Proposition 8.
For any selection rule ${\mathcal{L}}$, any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{ps}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{ps}}\in\mathfrak{S}$, the prediction set $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is a superset of the JOMI prediction set $\widehat{C}_{\alpha,n+j}$ defined in (8).
The procedure is summarized in Algorithm 4. For each $j$, the computation cost of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is $O(m)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; selection taxonomy $\mathfrak{S}$;
selective miscoverage level $\alpha$; preliminary micoverage level $\beta$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
Compute $q_{1}$ and $q_{2}$ as in (4.3). Compute $\eta^{+}$ as the $(K+1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$, where $K=\lceil(1-\beta)(n+1)\rceil$. Compute $\eta^{+}$ as the $(K-1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$ for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ as in (23).
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 4 JOMI for selection based on preliminary prediction sets
5 Application to drug discovery
In drug discovery, powerful prediction machines are increasingly used to guide the search of promising drug candidates. For such high-stakes decision-making, it is important to quantify the uncertainty in complex prediction machines (Svensson et al.,, 2017; Jin and Candès, 2023b, ; Laghuvarapu et al.,, 2024). Meanwhile, scientists will naturally focus on a subset of seemingly promising drugs.
In this section, we apply JOMI to several application scenarios in drug discovery with a selective nature. In some cases, JOMI yields shorter prediction intervals than vanilla conformal prediction when the latter is under-confident; in others, it makes the just right inflation of the prediction interval to provide exact selection-conditional coverage.
Application scenarios.
We study two problems in drug discovery:
  1. 1.
    Drug property prediction (DPP), a classification problem where the binary outcome indicates whether a drug candidate binds to a pre-specified disease target, and the covariates are the (encoded) chemical structure of the drug compound.
  2. 2.
    Drug-target-interaction prediction (DTI), a regression problem where each sample is a pair of drug and disease target. The outcome of interest is a real-valued variable indicating the binding affinity of that pair. The covariates are the (concatenated) encoded structure of the drug and the disease.
The datasets and prediction models are based on the DeepPurpose library (Huang et al.,, 2020), with details in subsequent subsections. In each problem, we consider several realistic selection rules.
Selection rules.
We consider three types of realistic selection rules ${\mathcal{S}}$:
  1. (1)
    Covariate-dependent top-K selection: selecting drugs with highest predicted binding affinities $\widehat{\mu}(X_{n+j})$.
    1. (i)
      Top-K selection for test data. When the scientist has a fixed budget of investigating $K$ drug candidates in the next phase, one may select $K$ test samples with the largest values of $\widehat{\mu}(X_{j})$.
    2. (ii)
      Top-K selection with mixed data. When the scientist is to investigate other properties for promisingly active drug candidates in the next phase, they may select $K$ units in ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$ with the highest predicted affinities.
    3. (iii)
      Calibration-referenced selection. The scientist may use the calibration data as reference and select test samples whose predicted activities are greater than the $K$-th highest in $\{\widehat{\mu}(X_{i})\}_{i\in{\mathcal{I}}_{\textnormal{calib}}}$.
  2. (2)
    Conformalized selection. The scientist might also obtain a subset of active drugs while controlling the FDR below some $q\in(0,1)$. In this case, $\widehat{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the set of test drugs picked by the conformalized selection method in Jin and Candès, 2023b at FDR level $q\in(0,1)$.
  3. (3)
    Selection with constraints. The cost of developing a drug is a function $C(x)$, and one wants to select as many drugs with highest predictions $\widehat{\mu}(X_{n+j})$ as possible while ensuring the total cost is below some constant $C$, or maximize the total reward while controlling the total budget.
For brevity, each selection rule is reported for one application, while we defer others to Appendix A.
Evaluation metrics.
We evaluate the selection-conditional (mis)coverage via the consistent estimator
$\widehat{\textrm{Miscov}}=\frac{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j})}{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}})},$
where $\widehat{P}$ is the empirical probability over all repeats. We also evaluate the size of prediction sets by averaging the cardinality of $\widehat{\mathcal{C}}_{j}$ for classification or length of $\widehat{\mathcal{C}}_{j}$ for regression over selected test units in all repeats. We also evaluate (but do not show in figures for brevity) the false coverage rate via $\widehat{\textrm{FCR}}=\widehat{E}\Big{[}\frac{\sum_{j=n+1}^{n+m}\mathds{1}\{j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j}\}}{1\vee|\widehat{\mathcal{S}}|}\Big{]},$ where $\widehat{E}$ is the empirical mean over all repeats.
5.1 Drug property prediction
We first consider the task of predicting the binary binding property of drug candidates. We use the HIV screening data in DeepPurpose with a total sample size of $n_{\textnormal{tot}}=41127$. The numerical features $X\in{\mathcal{X}}$ are encoded by Extended-Connectivity FingerPrints (Rogers and Hahn,, 2010, ECFP) which characterize topological properties of the drug candidates. A small-scale neural network $\widehat{\mu}\colon{\mathcal{X}}\to[0,1]$ is trained on randomly sampled $20\%$ of the entire dataset. Then, the remaining is randomly split into two equally-sized folds ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$. The exchangeability among data points is thus satisfied. Sampling of the training data and model training is independently repeated $10$ times; within each, we randomly split calibration/test data for $100$ times. This leads to $N=1000$ independent runs.
5.1.1 Top-K selection
We first consider the selection rule (i) top-K with test data. For a trained predictor $\widehat{\mu}$, we experiment with selecting (a) the highest $K$ values of $\widehat{\mu}(X_{j})$ and (b) the lowest $K$ values of $\widehat{\mu}(X_{j})$ among all test data. We observe similar patterns in (ii) top-K selection with both test and calibration data, as well as (iii) calibration-referenced selection. For brevity, we defer results for these selection rules to Appendix A.1.
For a confidence level $\alpha\in(0,1)$, we apply the vanilla conformal prediction and JOMI (both deterministic and randomized) to the selected units at level $\alpha$, with $\mathfrak{S}=2^{{\mathcal{I}}_{{\textnormal{test}}}}$. Since the selection size is fixed, valid selection-conditional coverage implies FCR control at level $\alpha$. We vary $\alpha\in\{0.1,0.2,\dots,0.9\}$ and $K\in\{20,100,1000,2000,5000,10000,15000\}$. We set $V$ as the APS score (Romano et al.,, 2020)
Figure 2 shows the empirical selection-conditional coverage $\widehat{\textnormal{Miscov}}$ on the left panel and the average prediction set size on the right panel, both at marginal miscoverage levels $\alpha\in\{0.1,0.8\}$. The solid (dashed) lines show the results when we select units with the highest (lowest) predicted affinities.
Figure 2: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for vanilla conformal prediction (Vanilla_CP), JOMI and randomized JOMI (JOMI_rand) for test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units. The reds dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$. The results are averaged over $N=1000$ runs.
The orange curves (Vanilla_CP) show that vanilla conformal prediction with APS score is over-confident for units with the lowest predicted affinities while under-confident for units with highest predicted affinities. In contrast, the purple (JOMI) and blue (JOMI_rand) curves both show valid coverage for our proposed methods. Also, using the APS score introduces visible gap between the actual coverage and $1-\alpha$ for JOMI due to distretization, which is made exact by its randomized version.
Interestingly, vanilla conformal prediction with the APS score yields zero-cardinality prediction sets for units with lowest predicted affinities with $\alpha=0.8$ (the orange dashed line in the right panel of Figure 2). This is because vanilla CP covers other test units with very high rate, and thus marginal coverage is guaranteed even with empty prediction sets for selected units. Of course, this is worrying if one cares more about these units at the bottom.
The selective behavior of these methods also depends on the choice of the nonconformity score $V$. Appendix A.2 shows the results for selection rules (i)-(iii) with the binary score. While our proposed methods consistently achieve valid coverage for selected units, with the binary score, vanilla CP is over-confident for units with highest predicted affinities yet under-confident for units with lowest predicted affinities. Also, the coverage gap of JOMI due to discretization is less visible with the binary score.
5.1.2 Conformalized selection
We then consider conformalized selection, where the focal test units are those believed to obey $Y=1$ with false discovery rate control at level $q\in(0,1)$. This problem was investigated in Bao et al., (2024) with no exact finite-sample coverage guarantees in theory (though their heuristic method performs reasonably in their empirical studies). We apply conformalized selection (Jin and Candès, 2023b, ) at FDR level $q\in\{0.2,\dots,0.9\}$ and $c_{i}\equiv 0.5$ to determine $\widehat{\mathcal{S}}$, and construct prediction intervals for units in $\widehat{\mathcal{S}}$ using either the APS score or the binary score. Experiments are repeated for $N=1000$ independent runs.
Figure 3 depicts the miscoverage $\widehat{\textnormal{Miscov}}$ and prediction interval lengths with nominal miscoverage level $\alpha\in\{0.1,0.8\}$ under various FDR levels $q$ and two choices of nonconformity score $V$.
Figure 3: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by conformalized selection (for finding $Y=1$ at target FDR levels $q\in\{0.1,0.2,\dots,0.9\}$) when applying vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), randomized JOMI (JOMI_rand), the method of Bao et al., (2024) (Heuristic and Heuristic_rand). Results with the APS score are in solid lines, while those with the binary score are in dashed lines. Details are otherwise as Figure 2.
Vanilla CP is still less calibrated for selected units: it is over-confident with both scores for $\alpha=0.1$, while being over-confident with binary score and under-confident with APS score at $\alpha=0.8$. In contrast, JOMI and JOMI_rand achieves valid selection-conditional coverage in all scenarios. There is some gap for JOMI with the APS score due to discretization, but not for JOMI_rand or the binary score. We observe an even lower empirical FCR than the conditional miscoverage for our methods (so they achieve valid FCR control); this is because the selection set can sometimes be empty for small values of FDR level $q$. Finally, compared with the heuristic methods of Bao et al., (2024), our method usually achieves smaller prediction set sizes whereas their method seems overly conservative.
5.1.3 Selection with constraints
We now consider selecting units with highest predicted binding affinities within a total budget of subsequent development. In this case, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})=\{j\in{\mathcal{D}}_{\textnormal{test}}\colon\widehat{\mu}(X_{j})\geq\widehat{\mu}_{[k]}\}$, where $\widehat{\mu}_{[k]}$ is the $k$-th order statistic of $\{\widehat{\mu}(X_{j})\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$, with $k=\max\{k\colon\sum_{\ell=k}^{m}L_{[k]}\leq C\}$, and $\{L_{j}\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$ are the costs.
We create semi-synthetic datasets since the original HIV data does not contain the cost information. Specifically, for each $i\in{\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i}))+2|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$ are i.i.d. random variables that capture other cost-related information. After leaving $20\%$ as the training set, we randomly sample the data without replacement so that $n=2500$ and $m=2500$.
The average miscoverage, prediction set size, and reference set size are reported in Figure 4. Interestingly, after adding cost constraints, vanilla conformal prediction is over-confident with the binary score and under-confident with the APS score. In contrast, our methods always yield near-exact coverage. From the right-most plot, we see that |^$|\widehat{\mathcal{R}}_{j}|$ is positively correlated with the number of selected test units.
Figure 4: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=500$ independent runs of Vanilla_CP, JOMI, and JOMI_rand. The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
5.2 Drug-target-interaction prediction
We now turn to a regression task, drug-target-interaction prediction. We use the DAVIS dataset in the DeepPurpose library (Huang et al.,, 2020). The outcome $Y$ is a continuously-valued variable indicating the binding affinity of a drug-target pair. The features $X$ combine the encodings of the drug compound and the target, where we encode the drug via a convolutional neural network (CNN) structure, and the target via a Transformer encoder. Our framework is agnostic to the choice of these encoding methods.
We train a point predictor $\widehat{\mu}(\cdot)\colon{\mathcal{X}}\to\mathbb{R}$ with $20\%$ randomly selected data in the entire dataset using a small neural network. The remaining $80\%$ data are then randomly split into two equally-sized folds as ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$. Similar to DPP, the training phase is repeated for $10$ independent runs; within each run, we independently run $100$ random split of calibration/test data and subsequent selection and construction of prediction sets. We use the nonconformity score $V(x,y)=|y-\widehat{\mu}(x)|$ throughout.
5.2.1 Conformalized selection for heterogeneous thresholds
We then study the behavior of vanilla CP and our proposed methods for selected units by running conformalized selection (Jin and Candès, 2023b, ) to find test pairs $j\in{\mathcal{I}}_{\textnormal{test}}$ whose outcomes are greater than data-dependent thresholds $c_{j}$. Here, we set $c_{j}$ as the $0.7$-th quantile of activities for training pairs with the same target sequence as the $j$-th data, while controlling the FDR below $q\in\{0.2,0.3,\dots,0.9\}$. The method of Bao et al., (2024) is not applicable to the heterogeneous thresholds here.
Figure 5 shows the miscoverage (the left panel) and average length of prediction intervals (the right panel) using JOMI and vanilla conformal prediction. For those most promising units picked by conformalized selection, vanilla CP (orange) tends to be over-confident with exceedingly high miscoverage. In contrast, our proposed methods achieve near-exact coverage across all values of $q$ and $\alpha$. Again, we observe that the FCR of our methods is even lower than $\widehat{\textnormal{Miscov}}$, empirically yielding valid FCR control.
Figure 5: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug-target-interaction for vanilla conformal prediction (Vanilla_CP) and JOMI (JOMI and JOMI_rand), at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for test units selected by conformalized selection at nominal FDR levels $q\in\{0.1,0.2,\dots,0.9\}$.
Our method can sometimes lead to two separate prediction intervals. We calculate the average number of intervals among selected test units in each experiment and report their boxplot in Figure 6. The frequency of producing two segments is very low across all cases.
Figure 6: Boxplot of the average number of intervals produced by JOMI and JOMI_rand among selected test units, over $N=1000$ independent runs, at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for various experiment configurations.
5.2.2 Selection under budget constraint
Finally, we consider a covariate-dependent selection rule, where the scientist aims to maximize the predicted activities of selected drugs subject to a fixed budget for subsequent development.
Formally, for each drug-target pair $i\in[m+n]$, we let $c_{i}$ be the $0.7$-th quantile of true activity scores in the training set with the same target, $\widehat{\mu}(X_{i})$ be its predicted binding affinity, and $L_{i}>0$ be its development costs. Then, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ aims to solve the following optimization problem:
$\mathop{\text{maximize}}_{S\subseteq\{n+1,\dots,m+m\}}$ $\quad\sum_{j\in S}\big{[}\widehat{\mu}(X_{j})-c_{j}\big{]}$ (25)
subject to $\quad\sum_{j\in S}L_{j}\leq\bar{L},$
where $\bar{L}=200$ is a budget limit. Again, as the dataset does not come with drug development costs, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
The optimization problem (25) is known as the Knapsack problem which is NP-hard. Nevertheless, there are efficient approximate solvers; in our experiments, we use the Python package mknapsack (mknapsack,, 2023). Existing methods such as Bao et al., (2024) cannot deal with such a complicated selection process. In contrast, our framework tackles this problem with a computation complexity that is polynomial in $m$, $n$, and the complexity of the subroutine ${\mathcal{S}}(\cdot,\cdot)$.
In Figure 7, we report the empirical miscoverage rate, average length of prediction intervals, and the sizes of ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ and the reference set ^$\widehat{\mathcal{R}}_{j}$ for all selected units, across $N=1000$ independent runs of our procedures. While vanilla conformal prediction is over-confident (with a much higher miscoverage rate), our methods yield exact coverage for the focal units. From the length of prediction sets, we see that longer prediction intervals are needed to cover promising drug-target pairs. While the number of selected test units are relatively stable, the reference set size may vary a bit.
Figure 7: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand. The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
6 Application to health risk prediction
Prediction machines are also widely used in healthcare for guiding clinical decision-making. These decisions may be determined by some complicated underlying processes such as clinical resource optimization (Ahmadi-Javid et al.,, 2017; Master et al.,, 2017) or preliminary uncertainty quantification (Olsson et al.,, 2022), and once made, would have substantial consequences.
In this section, we demonstrate the application of our framework to health risk prediction settings where units are picked by some complicated underlying processes. We will consider three selection rules:
  1. (1)
    Selection with constraints. Clinical decision makers may optimize a performance measure subject to certain constraints such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014), and send patients to different care categories. In Section 6.1, we consider a stylized example where we minimize the total predicted ICU stay subject to total cost budget.
  2. (2)
    Selecting small conformal prediction sets. Based on preliminary conformal prediction, one may suggest human intervention for units with large set sizes (Olsson et al.,, 2022; Sokol et al.,, 2024) while leaving those with small prediction sets unattended. In Section 6.2, we re-calibrate predictive inference for those whose preliminary prediction sets are shorter than a threshold.
  3. (3)
    Selecting lower prediction bounds. Practitioners may also focus on units whose preliminary prediction intervals lie above a threshold (Levitskaya,, 2023). In Section 6.3, we study selected test units whose preliminary prediction intervals have lower bounds above a threshold.
Among the above, (1) is covariate-dependent, and (2) and (3) depend on some preliminary uncertainty quantification criterion. All of them are efficiently tackled by the computation tricks in Section 4.
Our experiments use the ICU-stay data in the MIMIC-IV dataset (Johnson et al.,, 2023). Data (and features in $X$) are pre-processed using Gupta et al., (2022), with the outcome $Y$ being the length of ICU stay. We use random forests in the scikit-learn Python package to train a point prediction model $\widehat{\mu}(\cdot)$ and two quantile regression models $\widehat{q}_{1-\beta/2}(\cdot)$ and $\widehat{q}_{\beta/2}(\cdot)$ using a heldout training set. Calibration and test data are randomly split with $n=3000$ and $m=2000$.
6.1 Selection with constraints
We first study the case where test units are selected by minimizing the total predicted ICU stay time subject to a budget constraint. Given the point predictor $\widehat{\mu}(\cdot)$, we solve a Knapsack problem similar to (25) using the same approximate algorithm, where we maximize the summation of $\widehat{\mu}(X_{j})$ in the selected subset, and the cost $L_{j}$ is generated in the same way as Section 5.2.2 and then taken the ceiling.
Figure 8 shows the empirical miscoverage rate, length of prediction interval, and sizes of the selection set and reference sets. While vanilla conformal prediction is over-confident, our method achieves exact coverage for selected test units. We also see a slightly positive correlation of |^$|\widehat{\mathcal{R}}_{j}|$ and $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$.
Figure 8: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=500$ independent runs of Vanilla_CP, JOMI, and JOMI_rand. The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
6.2 Selecting small-sized prediction sets
We then consider the second selection rule, where we first build preliminary conformal prediction intervals $\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})$ via the score function $S(x,y)=|y-\widehat{\mu}(x)|/\widehat{\sigma}(x)$ (Lei et al.,, 2018), and both point prediction $\widehat{\mu}(\cdot)$ and conditional standard deviation estimation are trained via random forests. We then select those test units with $|\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})|\leq 5$, i.e., the upper and lower bounds of the preliminary prediction intervals are less than $5$ days apart. This mimics the ideas in (Ren et al.,, 2023; Sokol et al.,, 2024) where small-sized prediction sets are “certified” as confident.
After selection, we leverage the method in Section 4.3 to construct $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ for all selected test units $n+j$. Since $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is a superset of the exact output $\widehat{C}_{\alpha}(X_{n+j})$, we evaluate its empirical coverage to investigate whether it is over-conservative. Also, note from (23) that it is the union of three subsets; we also evaluate the number of disjoint segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$.
The miscoverage and length of prediction sets are reported in the first two plots in Figure 9. We observe that selectively certifying short prediction intervals can lead to under-coverage (orange curve), while JOMI achieves exact coverage (purple curve) by inflating the prediction sets, meaning that the superset $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is quite tight. The average number of disjoint segments in the right plot of Figure 9, which shows that $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is always one single interval.
Figure 9: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI when test units with short preliminary prediction sets are selected. The $x$-axis are nominal miscoverage levels $\alpha\in\{0.1,0.2,\dots,0.9\}$.
6.3 Selecting high prediction sets
Finally, we study selection rule (iii) which is also based on a preliminary prediction set constructed in the same away as Section 6.2. We imagine that practitioners select a test unit $n+j$ if the upper bound of $\widehat{C}^{\text{pre}}(X_{n+j})$ is below $6$, i.e., it appears the patient will stay in ICU less than 6 days.
The miscoverage rate, length of prediction sets, and number of disjoint segments averaged over $N=500$ independent runs of JOMI (using the method in Section 4.3) and vanilla conformal prediction are summarized in Figure 10. We see that preliminary prediction sets with low upper bounds tend to under-cover for small $\alpha$, while JOMI achieves exact coverage despite that we construct a superset of $\widehat{C}_{\alpha}(X_{n+j})$. However, JOMI may produce multiple segments for large values of $\alpha$.
Figure 10: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI. The nominal miscoverage levels on the $x$-axis are $\alpha\in\{0.1,0.2,\dots,0.9\}$.
Appendix A Additional experiment results
This section collects additional experimental results that are omitted in the main text. Appendix A.1 shows results for DPP with selection rules (ii) and (iii) with the APS score, omitted in Section 5.1. Appendix A.2 shows results for DPP with selection rules (i)-(iii) with the binary score. Appendix A.3 shows results for DTI with selection rules (ii) and (iii) that are omitted in Section 5.2.
A.1 Additional results for APS score in drug property prediction
In this part, we present results for DPP when test units are selected via (ii) top-K in both calibration and test data (Appendix A.1.1), and (iii) only calibration data as reference (Appendix A.1.2).
A.1.1 Top-K selection in mixed sample
We use the same values of $(\alpha,K)$ and the same scheme of repeated experiments as in Section 5.1.1, with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$. Figure 11 shows that vanilla conformal prediction incurs a large coverage gap for selected units, while JOMI and its randomized version both achieve valid selection-conditional coverage. The coverage gap due to discretization is eliminated again by randomization. Randomization also reduces the prediction set sizes on the right panel, but again, with $\alpha=0.8$, we see the concerning issue that it suffices for vanilla CP to construct empty prediction sets for selected units to achive marginal coverage.
Figure 11: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units whose $\widehat{\mu}(X_{n+j})$ are top-K (highest, solid line) or bottom-K (lowest, dashed line) among both calibration and test units. Details are otherwise the same as Figure 2.
A.1.2 Calibration-referenced selection
Now we select test units whose predicted affinities must be greater (or smaller) than the $K$-th largest (or smallest) predictions in the calibration set. We use the choices of $(\alpha,K)$ and the scheme of repeated experiments as in Section 5.1.1 with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$.
Figure 12: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by (iii) whose $\widehat{\mu}(X_{n+j})$ are greater than the $K$-th highest (solid line) or smaller than the $K$-th lowest (dashed line) in calibration units. Details are otherwise as Figure 2.
Figure 12 shows that our proposed methods achieve valid coverage (exact with JOMI_rand) while vanilla conformal prediction fails to. Again, vanilla CP is either over-confident or under-confident for units with lowest or highest predicted affinities.
A.2 Additional results for binary score in drug property prediction
In Figures 1314, and 15, we show the empirical selection-conditional coverage and prediction set sizes for selection rules (i)-(iii) when the prediction sets are constructed with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$. We observe that while vanilla CP is either over-confident or under-confident, our proposed methods achieve valid coverage across all configurations. The coverage gap of the non-randomized JOMI is smaller with this specific score function compared with the APS score.
Figure 13: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (i) top-K (or bottom-K) in test data. Details are otherwise as Figure 2.
Figure 14: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (ii) top-K (or bottom-K) in both calibration and test data. Details are otherwise as Figure 11.
Figure 15: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units that are (iii) greater than the $K$-th highest (or smaller than the $K$-th lowest) predicted affinites in the calibration data. Details are as Figure 12.
A.3 Additional results for drug-target-interaction prediction
In this section, we collect results for DTI with (1) Covariate dependent top-K selection. We present results under selection rule (1-i) in Appendix A.3.1, (1-ii) in Appendix A.3.2 and (1-iii) in Appendix A.3.3.
A.3.1 Top-K selection for fixed K
We consider the case where test units are selected via (i) top-K (or bottom-K) among predicted affinities in the test set. We apply the vanilla conformal prediction and our proposed methods with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$, so that selection-conditional coverage implies FCR control. Results under selection rules (ii) and (iii) demonstrate similar patterns, which we defer to Appendix A.3.
Figure 16 shows $\widehat{\textrm{Miscov}}$ (the left panel) and the average length of the prediction intervals (the right panel) for $K\in\{20,100,1000,2000,5000,10000\}$ and nominal miscoverage level $\alpha\in\{0.1,0.8\}$. We see that vanilla conformal prediction is over-confident for pairs with the highest predicted affinites, while being under-confident for the lowest. In contrast, our proposed methods achieve near-exact coverage under different nominal levels. By comparing the solid and dashed lines in the right panel, we observe that in this problem, the intrinstic uncertainty in higher-predicted-affinity pairs is larger, so a longer prediction interval is needed to achieve certain coverage.
Figure 16: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), and randomized JOMI applied to test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units. The red dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
A.3.2 Top-K selection in mixed sample
Figure 17 shows the miscoverage and average length of prediction intervals when (ii) the focal test units are those whose predicted affinities are among the $K$ highest or lowest in both calibration and test data. We see that vanilla CP is over-confident for highest-prediction units, while under-confident for lowest-prediction units. In contrast, our proposed methods always yield valid coverage.
Figure 17: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ among both calibration and test units. Details are otherwise the same as Figure 2.
A.3.3 Calibration-referenced selection
Figure 18 shows the results with (iii) calibration-referenced selection, i.e., the selected test units are those whose predicted affinities are greater than (or smaller than) the $K$-th highest (or lowest) in the calibration data. Similar to rules (i) and (ii), vanilla CP is not calibrated for these focal units, while our methods yield near-exact coverage across the whole spectrum of $K$.
Figure 18: Empirical FCR in drug-target interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ surpass the highest $K$ predicted affinities in the calibration data (or smaller than the lowest). Details are otherwise the same as Figure 16.
Appendix B Technical proofs
B.1 Proof of Proposition 1
[…]
$\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j}){\,|\,}j\in\widehat{\mathcal{S}})$ $=\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j})\,\big{|}\,j\in\widehat{\mathcal{S}},\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}\big{)}\cdot\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})$
$\geq(1-\alpha)\cdot\sum_{\ell\in{\mathcal{L}}}\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})=1-\alpha,$
where we use the fact that $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint and their union is the full set. Thus, it satisfies the weak selection-conditional coverage (2).
B.2 Proof of Proposition 2
Suppose $\widehat{C}_{\alpha}$ satisfies the strong selection-conditional coverage (3) for a class of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, it holds that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$. By definition, and since $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint, we have
FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}(\widehat{{\mathcal{S}}}),\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}$ (26)
$=\sum_{r=1}^{m}\sum_{\ell\in{\mathcal{L}}}\mathbb{E}\bigg{[}\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\cdot\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}.$ (27)
Noting that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ hence $|\widehat{{\mathcal{S}}}|=r(\ell)$ when $\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}$, we can rewrite the above as
$\sum_{\ell\in{\mathcal{L}}}\sum^{m}_{j=1}\frac{1}{r(\ell)}\mathbb{E}\big{[}\mathds{1}\big{\{}\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell},j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\big{\}}\bigg{]}$ (28)
[…]
where the last inequality follows from the strong selection-conditional coverage. Finally, the FCR control proof is completed by noting that
(28)$=\sum_{\ell\in{\mathcal{L}}}\frac{\alpha}{r(\ell)}\mathbb{E}\Big{[}|\widehat{{\mathcal{S}}}|\cdot\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\Big{]}=\alpha\cdot\mathbb{P}\big{(}\widehat{{\mathcal{S}}}\neq\varnothing\big{)}\leq\alpha.$
B.3 Proof of Proposition 3
Consider the following example with $m=2$ test units and ${\mathcal{Y}}=\mathbb{R}$. Suppose $\alpha\in(0,1)$. The selection rule returns
$\widehat{{\mathcal{S}}}=\begin{cases}\{1,2\}&\text{with probability }1-\frac{2\alpha}{1+\alpha},\\ \{1\}&\text{with probability }\frac{\alpha}{1+\alpha},\\ \{2\}&\text{with probability }\frac{\alpha}{1+\alpha}.\end{cases}$
Consider a prediction set $\widehat{C}_{\alpha}$ such that
$\widehat{C}_{\alpha}\big{(}X_{n+1};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+1};\{1\}\big{)}=\varnothing,\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{2\}\big{)}=\varnothing.$
Since the construction is symmetric in $1$ and $2$, we focus on $j=1$ without loss of generality.
[…]
For the FCR, we have
$\text{FCR}=\frac{\alpha}{1+\alpha}+\frac{\alpha}{1+\alpha}=\frac{2\alpha}{1+\alpha}>\alpha.$
Therefore, the constructed prediction set $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage but violates the FCR control.
B.4 Proof of Theorem 1
Proof of (a).
To prove (11), it suffices to show that
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{j}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S}\Big{)}\geq 1-\alpha.$ (32)
Recall that ${\mathcal{D}}_{j}={\mathcal{D}}_{{\textnormal{calib}}}\cup\{Z_{n+j}\}$, and ${\mathcal{D}}_{j}^{c}=\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}}$ is the set of test data without the $j$-th one; ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$ and ${\mathcal{I}}_{j}^{c}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$. Let $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ be the unordered set of ${\mathcal{D}}_{j}$, and we denote the realized value of the unordered set as $[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$, i.e., $z_{i}$’s are fixed realized values of $Z_{i}$’s, but we do not know which values in $[d_{j}]$ correspond to the calibration set and which to the test point. We also write $z_{i}=(x_{i},y_{i})$ and $v_{i}=V(x_{i},y_{i})$.
The exchangeability of $\{Z_{i}\}_{i\in{\mathcal{I}}_{j}}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$ implies that, given knowledge of the unordered realized data values in $[d_{j}]$, the probability of any assignment of these values to data points is equal. That is, for any permutation $\pi$ of ${\mathcal{I}}_{j}$,
$\mathbb{P}\big{(}Z_{1}=z_{\pi(1)},\dots,Z_{n}=z_{\pi(n)},Z_{n+j}=z_{\pi(n+j)}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{1}{(n+1)!}.$
Next, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we define the following subset of ${\mathcal{I}}_{j}$:
$\widetilde{R}=\big{\{}i\in{\mathcal{I}}_{j}\colon j\in{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i}),\,{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\in\mathfrak{S}\big{\}},$
where $[D_{\textnormal{calib}}^{i}]=[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ is an unordered set, and $D_{\textnormal{test}}^{i}=(X_{n+1},\dots,x_{i},\dots,X_{n+m})$. Note that the collection $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, so $\widetilde{R}$ is also fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
Recall that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})=\{i\in{\mathcal{I}}_{\textnormal{calib}}:j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\}$. We let ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})$ if $j\notin\widehat{\mathcal{S}}$ and ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}$ if $j\in\widehat{\mathcal{S}}$. We claim that given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}=\{v_{i}\colon i\in\widetilde{R}\},$ (33)
where repetition of elements is allowed on both sides. To see this, let $\pi$ denote the permutation on ${\mathcal{I}}_{j}$ such that $V_{i}=v_{\pi(i)}$ and $V_{n+j}=v_{\pi(n+j)}$. Then
$\widehat{\mathcal{S}}={\mathcal{S}}\big{(}\underbrace{[z_{\pi(1)},z_{\pi(2)},\dots,z_{\pi(n)}]}_{\textrm{calibration data}},\underbrace{X_{n+1},\dots,x_{\pi(n+j)},\dots,X_{n+m}}_{\textrm{test data}}\big{)}={\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(n+1)}],D_{\textnormal{test}}^{\pi(n+1)}\big{)},$ (34)
$\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}={\mathcal{S}}\big{(}\underbrace{[z_{\pi(1)},\dots,z_{\pi(i-1)},z_{\pi(n+j)},z_{\pi(i+1)},\dots,z_{\pi(n)}]}_{\textrm{calibration data after swap}},\underbrace{X_{n+1},\cdots,x_{\pi(i)},\dots,X_{n+m}}_{\textrm{test data after swap}}\big{)}$
$\qquad\qquad={\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}$
for all $i\in\{1,\dots,n\}$. Denoting $\widehat{\mathcal{S}}=\widehat{\mathcal{S}}^{{\textnormal{swap}(n+j,j)}}$ for notational simplicity, we have,
$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}$ =$\overset{\rm(a)}{=}\big{\{}V_{i}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\big{\}}$
=$\overset{\rm(b)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}\big{\}}$
=$\overset{\rm(c)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\pi(i)}],D_{{\textnormal{test}}}^{\pi(i)}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}\in\mathfrak{S}\big{\}}$
=$=\big{\{}v_{\ell}\colon\ell\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\ell}],D_{{\textnormal{test}}}^{\ell}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}\in\mathfrak{S}\big{\}}$
=$\overset{\rm(d)}{=}\{v_{i}\colon i\in\widetilde{R}\}.$
Above, step (a) follows from the definition of ^$\widehat{\mathcal{R}}_{n+j}^{+}$; steps (b) and (c) are due to (34), and step (d) follows from the definition of ${\widetilde{R}}$. We thus completed the proof of Eqn. (33).
Returning to the proof of (32), we have by Bayes’ rule that
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S},\leavevmode\nobreak\ [d_{j}],{\mathcal{D}}_{j}^{c}\Big{)}$ (35)
[…]
Again, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, the only randomness is the assignment of $[z_{1},\dots,z_{n},z_{n+j}]$ to $(Z_{1},\dots,Z_{n},Z_{n+j})$. By definition and the exchangeability of ${\mathcal{D}}_{j}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$, it is straightforward that
$\mathbb{P}\big{(}j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}$ $=\sum_{i\in{\mathcal{I}}_{j}}\mathbb{P}\big{(}Z_{n+j}=z_{i}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}\cdot\mathds{1}\{i\in{\widetilde{R}}\}$ (36)
$=\sum_{i\in{\mathcal{I}}_{j}}\frac{\mathds{1}\{i\in\widetilde{R}\}}{n+1}=\frac{|\widetilde{R}|}{n+1}.$ (37)
On the other hand, by (33),
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\Big{(}1-\alpha;\{V_{i}\}_{i\in{\mathcal{R}}_{n+j}^{+}(Y_{n+j})}\Big{)},j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\Big{)}$
[…]
Since ${\widetilde{R}}$ is determistic conditional on $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
($=\leavevmode\nobreak$ $\sum_{\ell\in{\widetilde{R}}}\mathbb{P}\big{(}V_{n+1}=v_{\ell}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}\cdot\mathds{1}\{v_{\ell}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{v_{i}\colon i\in\widetilde{R}\}\}$ (40)
[…]
where the last step is because $\mathbb{P}(Z_{n+1}=z_{\ell}{\,|\,}[D_{n+1}],{\mathcal{D}}_{n+1}^{c})=1/(n+1)$ for every $\ell$. By the definition of the quantile function,
(41)$\geq\frac{(1-\alpha)|\widetilde{R}|}{n+1}.$
Combining the above and (37), we have thus verified (35). Then by the tower property, we conclude (32), and thus (11).
When the elements in ${\widetilde{R}}$ are distinct, we additionally have the following determistic inequality following from the definition of the quantile function:
(41)$\leq\frac{(1-\alpha)|{\widetilde{R}}|+1}{n+1},$
which then leads to
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\Big{(}1-\alpha;\{V_{i}\}_{i\in{\mathcal{R}}_{n+j}^{+}(Y_{n+j})}\Big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S},[d_{j}],{\mathcal{D}}_{j}^{c}\Big{)}$
[…]
where the last step is because we are conditioning on the event $j\in\widehat{\mathcal{S}}$. Taking expectation with respect to $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$ on both sides, we obtain (12) and complete the proof of statement (a).
Proof of (b).
To conveniently denote repeated values in $[z_{1},\dots,z_{n},z_{n+j}]$ and $[v_{1},\dots,v_{n},v_{n+j}]$ we let $[1],[2],\dots,[n],[n+j]$ be a permutation of ${\mathcal{I}}_{j}$ so that $v_{[1]}\leq v_{[2]}\leq\cdots\leq v_{[n]}\leq v_{[n+j]}$. Then $[d_{j}]$ can be equivalently represented by $\{z_{[i]}\}_{i\in{\mathcal{I}}_{j}}$, and for any $i\in{\mathcal{I}}_{j}$,
$\mathbb{P}\big{(}Z_{n+j}=z_{[i]}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{1}{n+1}.$
Following the proof of part (a) up to Equation (33), we know that
$\mathbb{P}\big{(}j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{|\widetilde{R}|}{n+1}.$
On the other hand, due to (33) and the fact that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}=\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})$. when $j\in\widehat{\mathcal{S}}$, for any $t\in(0,1)$,
$\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{i}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{i}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
$=$ $\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{[i]}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{[i]}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
$=$ $\leavevmode\nobreak\ \sum_{\ell\in{\mathcal{I}}_{j}}\mathbb{P}\bigg{(}Z_{n+j}=z_{[\ell]},\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{[i]}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{[i]}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)},$
where the last equality is a decomposition of disjoint events. When $Z_{n+j}=z_{[\ell]}$, note that $j\in\widehat{\mathcal{S}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$ is equivalent to $[\ell]\in\widetilde{R}$. Thus, the last equation above equals to
$\sum_{\ell\in{\mathcal{I}}_{j}}\mathbb{P}\bigg{(}Z_{n+j}=z_{[\ell]},\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{v_{[\ell]}<v_{[i]}\}+U_{j}\cdot\sum_{i\in\widetilde{R}}\mathds{1}\{v_{[\ell]}=v_{[i]}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ [\ell]\in\widetilde{R}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
[…]
where we used the take-out property of conditional probability/expectation in the first step; the probability in the second line is only with respect to $U_{j}\sim\textnormal{Unif}[0,1]$, and the last equality is a standard result that is straightforward to check with exchangeability (see, e.g., Vovk et al., (2005, Proposition 2.4)). Putting them togther and letting $t=1-\alpha$, we have
$\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}\,\big{|}\,j\in\widehat{\mathcal{S}},[d_{j}],{\mathcal{D}}_{d}^{c}\big{)}=\frac{\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j},\leavevmode\nobreak\ j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}}{\mathbb{P}\big{(}j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}}=1-\alpha,$
which concludes the proof of statement (b).
B.5 Proof of Proposition 4
We start by defining an alternative stopping time that is determined solely by ${\mathcal{D}}_{j}^{c}$:
${\widetilde{T}}_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\geq m-K\Big{\}}.$
Abusing the notation a bit, we call $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$ and ${\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ the corresponding stopping times defined with $({\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}$, ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)})$. Since ${\widetilde{T}}_{\textnormal{topk}}$ is agnostic to the swapping of unit $i$ and $n+j$, ${\widetilde{T}}^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
We claim that when $S_{n+j}>T_{\textnormal{topk}}$, $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, To see this, note first that for any $t\in{\mathcal{T}}$,
$\sum_{i\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{S_{i}\leq t\}\geq\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ {\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}.$
On the other hand, one can check that
$\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}\overset{\rm(a)}{=}\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}+\mathds{1}\{S_{n+j}\leq T_{\textnormal{topk}}\}\overset{\rm(b)}{\geq}m-K.$ (42)
Above, step (a) is because $S_{n+j}>T_{\textnormal{topk}}$ and step (b) follows from the choice of $T_{\textnormal{topk}}$. With the definition of ${\widetilde{T}}_{\textnormal{topk}}$, we can then conclude from (42) that $T_{\textnormal{topk}}\geq{\widetilde{T}}_{\textnormal{topk}}$, and hence $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
Similarly, we have ${\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ by construction. When $S_{i}>T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$,
$\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}=\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}+\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}\geq m-K.$
By the definition of $T_{\textnormal{topk}}$, we then have $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\geq T_{\textnormal{topk}}$, and therefore $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}=T_{\textnormal{topk}}$.
Fix $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$, we proceed to show that ^$\widehat{{\mathcal{R}}}_{\textnormal{fixed}}=\widehat{{\mathcal{R}}}_{n+j}$. For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{topk}}$, there is $S_{i}>T_{\textnormal{topk}}$. We then have $S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$, since $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}\geq T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$. Therefore, $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, and
$\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ =$=\big{\{}\ell\in[m]:S_{n+\ell}^{{\textnormal{swap}(i,j)}}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\big{\}}=\big{\{}\ell\in[m]:S_{n+\ell}^{\textnormal{swap}(i,j)}>{\widetilde{T}}_{\textnormal{topk}}\big{\}}$
=$\overset{\rm(a)}{=}\{j\}\cup\big{\{}\ell\in[m]\backslash\{j\}:S_{n+\ell}>{\widetilde{T}}_{\textnormal{topk}}\big{\}}$
=$=\{j\}\cup\big{\{}\ell\in[m]\backslash\{j\}:S_{n+\ell}>T_{\textnormal{topk}}\big{\}}=\widehat{{\mathcal{S}}},$
wher step (a) follows from the fact that $S_{n+j}^{\textnormal{swap}(i,j)}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{topk}}$. As a result, $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$. On the other hand, for any $i\in\widehat{{\mathcal{R}}}_{n+j}$, we have that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$. Then there is
$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{topk}}\Rightarrow S_{i}>T_{\textnormal{topk}}.$ (43)
Combining all the above, we complete the proof.
B.6 Proof of Proposition 5
Fix $j\in\widehat{{\mathcal{S}}}$. Define the proxy stopping time that is agnostic to the swapping of units $i$ and $n+j$:
${\widetilde{T}}_{\textnormal{cq}}=\inf\bigg{\{}t\in{\mathcal{T}}:\sum_{i\in{\mathcal{I}}_{j}}\mathds{1}\{S_{i}\leq t\}\geq qn\bigg{\}}.$
By construction, ${\widetilde{T}}_{\textnormal{cq}}\leq T_{\textnormal{cq}}$. We claim that when $S_{n+j}>{\widetilde{T}}_{\textnormal{cq}}$, $T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$. To see this, note that
$\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$ (44)
where the last inequality follows from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. By the definition of $T_{\textnormal{cq}}$, (44) implies that ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}$ and therefore $T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
Similarly, we have ${\widetilde{T}}_{\textnormal{cq}}\leq T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}$ by construction. When $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$,
$\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$
with the last step following from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. We then have ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}^{\textnormal{swap}(i,j)}$ and therefore $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
Next, we proceed to show that ^$\widehat{{\mathcal{R}}}_{n+j}=\widehat{{\mathcal{R}}}_{\textnormal{cq}}$. For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{cq}}$, there is $S_{i}>T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$. Then, $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$, and
$\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ =$=\{\ell\in[m]:S_{n+\ell}^{\textnormal{swap}(i,j)}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\}{=}\{\ell\in[m]:S_{n+\ell}^{\textnormal{swap}(i,j)}>{\widetilde{T}}_{\textnormal{cq}}\}$
=$\overset{\rm(a)}{=}\{j\}\cup\{\ell\in[m]\backslash\{j\}:S_{n+\ell}>{\widetilde{T}}_{\textnormal{cq}}\}=\widehat{{\mathcal{S}}}_{\textnormal{cq}},$
where step (a) follows from the fact that $S^{\textnormal{swap}(i,j)}_{n+j}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$. The above implies that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and and we have $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$. Reversely, for $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$,
$S_{n+j}^{\textnormal{swap}(i,j)}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{cq}}\Rightarrow S_{i}>T_{\textnormal{cq}},$
completing the proof.
B.7 Proof of Proposition 6
To start, we claim that the $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ is equivalent to
${\widetilde{S}}_{\textnormal{fixed}}=\{j\in[m]:\widehat{{\mathcal{S}}}_{n+j}\geq T\},\text{ where }T=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}\geq q(n+1)\bigg{\}}.$
To see this, suppose $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$, then $p_{j}\leq q$. We can check
$\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}=p_{j}\leq q.$
By the definition of $T$, $\widehat{S}_{n+j}\geq T$. Conversely, if $\widehat{S}_{n+j}\geq T$, then
$p_{j}=\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}\leq\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T,Y_{i}\leq c_{i}\}\Big{)}\leq q\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}.$
We then have $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}={\widetilde{S}}_{\textnormal{fixed}}$. We shall work with ${\widetilde{S}}_{\textnormal{fixed}}$ in the following.
Fix $j\in{\widetilde{S}}_{\textnormal{fixed}}$ and a selection taxonamy $\mathfrak{S}$ such that ${\widetilde{S}}_{\textnormal{fixed}}\in\mathfrak{S}$. We prove the proposition for $y\leq c_{n+j}$ and $y>c_{n+j}$ separately. For notational simplicity, we define the swapped stopping time for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ and $y\in{\mathcal{Y}}$ as
$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$
Case 1: $y\leq c_{n+j}$.
The goal is to show that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$. We prove this for $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
  • $Y_{i}\leq c_{i}$.
    Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$, i.e., $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
    ${1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i},y\leq c_{n+j}\}}\leq q(n+1).$ (45)
    Since $y\leq c_{n+j}$, the above inequality is equivalent to
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}\leq q(n+1).$ (46)
    Since $\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{i},Y_{i}\leq c_{i}\}=1$, we further have
    ${\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}}\leq q(n+1).$
    The above and the definition of $T_{j}^{\textnormal{fixed}}(1,0)$ implies that $\widehat{{\mathcal{S}}}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)$. Next, note that
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{{\mathcal{S}}}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$
[…]
    where the first step is because $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$; the second step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$. Eqn. (B.7) implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{{\textnormal{swap}(i,j)}}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{{\textnormal{swap}(i,j)}}(y),Y_{i}\leq c_{i}\}=1$, we have
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}$
[…]
    following the definition of $T^{\textnormal{swap}(i,j)}(y)$. We then conclude that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{fixed}}_{j}(1,0)$ and therefore $T_{j}^{(1,0)}=T^{{\textnormal{swap}(i,j)}}(y)$. As a result,
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}.$
    We have therefore verified that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
    Next, given $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, our goal now is to show that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. By definition, there is
    $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)\text{ and }\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
    Since $y\leq c_{n+j}$ and $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,0),Y_{i}\leq c_{i}\}=1$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{{\mathcal{S}}}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$
[…]
    where the last step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$. The above implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$, and therefore $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$, and $y\leq c_{n+j}$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}$
[…]
    following the definition of $T^{\textnormal{swap}(i,j)}(y)$. As a result, $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(1,0)\Rightarrow T_{j}^{\textnormal{fixed}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$. The swapped selection set
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)$ =$=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
    It follows that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
  • $Y_{i}>c_{i}$.
    Note that for any $t\in{\mathcal{T}}$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t\}$
[…]
    becuase $y\leq c_{n+j}$ and $Y_{i}>c_{i}$. Inspecting the definition of $T_{j}^{\textnormal{fixed}}(1,1)$ and $T^{\textnormal{swap}(i,j)}(y)$, we can conclude that $T_{j}^{\textnormal{fixed}}(1,1)=T^{\textnormal{swap}(i,j)}(y)$. As a result, we have $\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
    $\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,1)\}=\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}.$
    Therefore $i\in\widehat{{\mathcal{R}}}_{n+y}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed}}(y)$.
Combing the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$ for $y\leq c_{n+j}$.
Case 2: $y>c_{n+j}$.
We again consider the case of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$ separately.
  • $Y_{i}\leq c_{i}$.
    Fix $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. By definition, there is $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $y>c_{n+j}$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$
[…]
    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$. We thus have $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
[…]
    by the definition of $T_{j}^{\textnormal{fixed}}(0,0)$. As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)=T^{\textnormal{swap}(i,j)}(y)$, and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(0,0)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    It then follows that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$ Note that since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
[…]
    following the definition of $T_{j}^{\textnormal{fixed}}(0,0)$. As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$
[…]
    by the choice of $T^{\textnormal{swap}(i,j)}(y)$. The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)\Rightarrow T^{\textnormal{swap}(i,j)}=T_{j}^{\textnormal{fixed}}(0,0)$. Consequently, we have
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$
    We have thus shown that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
  • $Y_{i}>c_{i}$.
    Since $y>c_{n+j}$ and $Y_{i}>c_{i}$, for any $t\in{\mathcal{T}}$ there is
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}=1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}.$
    implying that $T_{j}^{\textnormal{fixed}}(0,1)=T^{\textnormal{swap}(i,j)}(y)$. Consequently, $\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    Then there is $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
Combining the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}=\widehat{{\mathcal{R}}}_{n+j}(y)$, and therefore ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$, for any $y>c_{n+j}$. The proof is now complete.
B.8 Proof of Proposition 7
We start from making the observation that $\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\in[m]:\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$, where
$T^{\textnormal{BH}}=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq t\})\vee 1}\frac{m}{1+n}\leq q\bigg{\}}.$
To see this, let $j^{*}=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$ and $\widehat{S}_{n+(1)},\widehat{S}_{n+(2)},\ldots,\widehat{S}_{n+(m)}$ denote the order statistics of $\{\widehat{S}_{n+j}\}_{j\in[m]}$ in an decending order. It then suffices to show that $j^{*}=k^{*}$. When $k^{*}\geq 1$,
$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq\widehat{S}_{n+(k^{*})},Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq\widehat{S}_{n+(k^{*})}\})\vee 1}\frac{m}{1+n}=p_{(k^{*})}\frac{m}{k^{*}}\leq q,$ (48)
where the last inequality follows from the definition of $k^{*}$. From the definition of $T^{\textnormal{BH}}$, we can conclude that $\widehat{S}_{n+(k^{*})}\geq T^{\textnormal{BH}}$; since $\widehat{S}_{n+(j^{*})}$ is the smallest element in $\{\widehat{S}_{n+j}\}_{j\in[m]}$ that is no less than $T^{\textnormal{BH}}$, we also have that $\widehat{S}_{n+(k^{*})}\geq\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$. Conversely,
$p_{(j^{*})}\frac{m}{j^{*}}$ =$\overset{\rm(a)}{=}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+(j^{*})},Y_{i}\leq c_{i}\}}{n+1}\cdot\frac{m}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}},Y_{i}\leq c_{i}\}}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}\cdot\frac{m}{n+1}\overset{\rm(c)}{\leq}q.$
Above, step (a) uses the definition of $j^{*}$, step (b) is because $\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$ and step (c) is by the definition of $T^{\textnormal{BH}}$. From the form of the BH procedure, we have that $j^{*}\geq k^{*}$, and therefore $j^{*}=k^{*}$.
When $k^{*}=0$, for any $t\in{\mathcal{T}}$, we let $j(t)=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\}$. If $j(t)>0$, then
$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{n+1}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{j(t)}\frac{m}{n+1}$
[…]
where the last step is because $k^{*}=0$. The above implies that $T^{\textnormal{BH}}>\widehat{S}_{n+(1)}\Rightarrow j^{*}=0$. The equivalence is therefore established.
As before, we define for any $i\in[n]$, $j\in[m]$, and $y\in{\mathcal{Y}}$ the swapped stopping time:
$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$.
We are now ready to prove the proposition. We consider the two cases of $y\leq c_{n+j}$ and $y>c_{n+j}$ separately.
Case 1: $y\leq c_{n+j}$.
We discuss the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
  • $Y_{i}\leq c_{i}$. Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),T_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{1+n}$
[…]
    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$. By the choice of $T^{\textnormal{BH}}_{j}(1,0)$, there is $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(1,0)$. As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$. Conversely,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
[…]
    following from the definition of $T^{\textnormal{BH}}_{j}(1,0)$. We therefore have $T^{\textnormal{BH}}_{j}(1,1)\geq T^{\textnormal{swap}(i,j)}(y)$, and then $T^{\textnormal{BH}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$. The above also implies that
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    By definition, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
    On the other hand, suppose $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH}}(y)$, we have $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}$. We can check that
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
[…]
    by the definition of $T^{\textnormal{BH}}_{j}(1,0)$; the definition of $T^{\textnormal{swap}(i,j)}(y)$ implies that $T^{\textnormal{BH}}_{j}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$, there is
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
[…]
    and therefore $T^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\Rightarrow T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(1,0)$. We then have that $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
    ^$\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}.$
    That is, $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. We have thus shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}\leq c_{i}$.
  • If $Y_{i}>c_{i}$, then for any $t\in{\mathcal{T}}$,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    By their definition, we have $T^{\textnormal{BH}}_{j}(1,1)=T^{\textnormal{swap}(i,j)}$. As a consequence,
    $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\iff j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,1)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    As a result, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}>c_{i}$.
Concluding the above two cases, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)$ when $y\leq c_{n+j}$.
Case 2: $y>c_{n+j}$.
We again prove ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$ for $y>c_{n+j}$ in two cases.
  • $Y_{i}\leq c_{i}$. Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
[…]
    by the construction of $T^{\textnormal{swap}(i,j)}(y)$; therefore, $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)$. Meanwhile, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{i}\leq c_{i}\}=1$,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}}\frac{m}{n+1}$
[…]
    following from the definition of $T^{\textnormal{BH}}_{j}(0,0)$. Therefore $T^{\textnormal{BH}}_{j}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$. As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,0)$ and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0)$ and therefore
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    which is a result of the definition of $T^{\textnormal{BH}}(0,0)$. The above also implies that $T^{\textnormal{BH}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Next,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}\}}\frac{m}{n+1}$
[…]
    where the last step is a result of $T^{\textnormal{swap}(i,j)}(y)$’s definition. The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$. We then have
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}\in\mathfrak{S}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Combining the above, we have shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)$ for any $i$ such that $Y_{i}\leq c_{i}$.
  • $Y_{i}>c_{i}$. In this case, we have for any $t\in{\mathcal{T}}$ that
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    By their definitions, there is $T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(0,1)$, and therefore
    $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$
    and
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,1)\}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i$ such that $Y_{i}>c_{i}$.
Combining the two cases, we have shown that ^$\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $y>c_{n+j}$. The proof is now complete.
B.9 Proof of Proposition 8
Recall that
$\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(y)}\cup\{\infty\},1-\alpha\big{)}\Big{\}},$
where ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ ia the reference set. We consider the following three cases: $S(X_{n+j},y)<\eta^{-}$, $S(X_{n+j},y)>\eta^{+}$ and $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$. The case where $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$ is obvious, so we discuess the remaining two cases in the following.
As before, we define $\eta^{\textnormal{swap}(i,j)}(y)$ to be the $K$-th smallest elements in $\{\widehat{S}^{\textnormal{swap}(i,j)}_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}$.
Case 1: $S(X_{n+j},y)<\eta^{-}$.
Consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)<\eta^{-}$ and $y\in\widehat{C}_{\alpha,n+j}$. We claim that in this case
^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta^{-},{\mathcal{L}}(X_{i},\eta^{-})=1\}.$
To see this, note that when $S_{i}\leq\eta^{-}$, we have $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$. and therefore ${\mathcal{L}}(X_{i},\eta)={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$; when $S_{i}>\eta^{-}$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{-}$, and therefore ${\mathcal{L}}(X_{i},\eta^{-})={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$. As a result, we have that ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}$ and therefore
$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(1)}}\cup\{\infty\},1-\alpha\big{)}$
Case 2: $S(X_{n+j},y)>\eta^{+}$.
We now consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)>\eta^{+}$ and $y\in\widehat{C}_{\alpha,n+j}$. In this case, we instead have
^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(2)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta,{\mathcal{L}}(X_{i},\eta^{+})=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta,{\mathcal{L}}(X_{i},\eta)=1\}.$
This is because when $S_{i}\leq\eta$, $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{+}$ and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta^{+})$; when $S_{i}>\eta$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$, and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta)$. Consequently,
$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(2)}}\cup\{\infty\},1-\alpha\big{)}.$
Combining the three cases, we finish the proof.
Appendix C Auxiliary results
C.1 The conformal p-values
The conformal p-value and its random counterpart defined in Jin and Candès, 2023b are
$\bar{p}_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}\text{ and }$
$\bar{p}^{\textnormal{rand}}_{j}=\frac{U_{n+j}(1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}=\widehat{S}_{n+j}\})+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}>\widehat{S}_{n+j}\}}{n+1}.$
Lemma 1.
Under the condition of Theorem 1, for any $j\in[m]$, the conformal p-values $p_{j}$ and $p_{j}^{\textnormal{rand}}$ defined in (17) satisfy that
$\mathbb{P}\big{(}p_{j}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\text{ and }\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,$
for any $t\in[0,1]$.
Furthermore, there is $p_{j}\leq\bar{p}_{j}$ and $p_{j}^{\textnormal{rand}}\leq\bar{p}_{j}^{\textnormal{rand}}$ deterministically, where the equalities hold for the nonformity score $S(x,y)=\widehat{\mu}(X)-\big{(}\sup_{x}2|\widehat{\mu}(x)|\big{)}\cdot y$.
Proof.
For any $j\in[m]$ and $t\in[0,1]$, it can be checked that when $Y_{n+j}\leq c_{n+j}$,
$p_{j}\leq t\iff\widehat{{\mathcal{S}}}_{n+j}\text{ is among the }k^{*}\text{-th largest element in }{\mathcal{N}}_{j}:=\{\widehat{S}_{i}:i\in{\mathcal{I}}_{j},Y_{i}\leq c_{i}\},$
where $k^{*}=\min(|{\mathcal{N}}_{j}|,\lfloor t(n+1)\rfloor)$. Let $T_{k^{*}}$ denote the value of the $k^{*}$-th largest element in ${\mathcal{N}}_{j}$. Then apparently, both $k^{*}$ and $T_{k^{*}}$ are invariant to the permutation on ${\mathcal{I}}_{j}$. Consequently,
$\mathbb{P}(p_{j}\leq t,Y_{n+j}\leq c_{n+j})$ $=\mathbb{E}\big{[}\mathds{1}\{Y_{n+j}\leq c_{n+j},\widehat{S}_{n+j}\geq T_{k^{*}}\}\big{]}\overset{\rm(a)}{=}\frac{1}{n+1}\sum_{i\in{\mathcal{I}}_{j}}\mathbb{E}\Big{[}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{k^{*}}\}\Big{]}$
$=\frac{1}{n+1}\mathbb{E}\Big{[}\sum_{i\in{\mathcal{N}}_{j}}\mathds{1}\{\widehat{S}_{i}\geq T_{k^{*}}\}\Big{]}\overset{\rm(b)}{\leq}\frac{k^{*}}{n+1}\leq t,$
where step (a) is due to the exchangeability of ${\mathcal{D}}_{j}$ and step (b) follows from the fact that the number of elements greater than or equal to $T_{k^{*}}$ in ${\mathcal{N}}_{j}$ is at most $k^{*}$.
Next, we consider the randomized version. For any $i\in{\mathcal{I}}_{j}$, let $E_{i}$, $G_{i}$ and $G^{+}_{i}$ denote the number of elements in ${\mathcal{N}}_{j}$ equal to, no less than, and greater than $\widehat{S}_{i}$, respectively, i.e.,
$E_{i}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}=\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\},\leavevmode\nobreak\ G_{i}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}\leq\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\},\leavevmode\nobreak\ G_{i}^{+}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}<\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\}.$
Using these notations,
$\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}=\mathbb{P}\big{(}U_{n+j}E_{n+j}+G_{n+j}^{+}\leq(n+1)t,Y_{n+j}\leq c_{n+j}\big{)}$ (49)
[…]
where the second step follows from exchangeability. Let $i^{*}$ denote the index of the largest element in $\{\widehat{S}_{i}:i\in{\mathcal{N}}_{j},G_{i}^{+}<(n+1)t\}$ (with ties broken arbitrarily). Then
($=\leavevmode\nobreak$ $\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}|{\mathcal{N}}_{j}|+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot\Big{(}G_{i^{*}}^{+}+E_{i^{*}}\cdot\mathds{1}\Big{\{}U_{n+j}\leq\frac{(n+1)t-G_{i^{*}}^{+}}{E_{i^{*}}}\Big{\}}\Big{)}\Big{]}$
$=\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}(n+1)t+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot(n+1)t\Big{]}$
$\leq t.$
By definition,
$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1}$ $\overset{\rm(a)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},S_{i}\geq\widehat{S}_{n+j}\}}{n+1}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}=\bar{p}_{j}.$
Above, the step (a) is because when $Y_{i}\leq c_{i}$, $S_{i}\geq\widehat{S}_{i}$; when $S(x,y)=\widehat{\mu}(x)-(\sup_{x}2|\widehat{\mu}(x)|)\cdot\mathds{1}\{y>c\}$, the $\leq$” is replaced by “$=$” in step (a) and (b). ∎

### v3

@@
 Abstract
-Conformal prediction builds marginally valid prediction intervals which cover the unknown outcome of a randomly drawn new test point with a prescribed probability.
-In practice, a common scenario is that, after seeing the test unit(s), practitioners decide which test unit(s) to focus on in a data-driven manner, and wish to quantify the uncertainty for the focal unit(s).
-In such cases, marginally valid prediction intervals for these focal units can be misleading due to selection bias.
-This paper presents a general framework for constructing a prediction set with finite-sample exact coverage conditional on the unit being selected.
-Its general form works for arbitrary selection rules, and generalizes Mondrian Conformal Prediction to multiple test units and non-equivariant classifiers.
-We then work out computationally efficient implementation of our framework for a number of realistic selection rules, including top-K selection, optimization-based selection, selection based on conformal p-values, and selection based on properties of preliminary conformal prediction sets.
+Conformal prediction builds marginally valid prediction intervals that cover the unknown outcome of a randomly drawn test point with a prescribed probability.
+However, in practice, data-driven methods are often used to identify specific test unit(s) of interest, requiring uncertainty quantification tailored to these focal units.
+In such cases, marginally valid conformal prediction intervals may fail to provide valid coverage for the focal unit(s) due to selection bias.
+This paper presents a general framework for constructing a prediction set with finite-sample exact coverage, conditional on the unit being selected by a given procedure.
+The general form of our method accommodates arbitrary selection rules that are invariant to the permutation of the calibration units, and generalizes Mondrian Conformal Prediction to multiple test units and non-equivariant classifiers.
+We also work out computationally efficient implementation of our framework for a number of realistic selection rules, including top-K selection, optimization-based selection, selection based on conformal p-values, and selection based on properties of preliminary conformal prediction sets.
 The performance of our methods is demonstrated via applications in drug discovery and health risk prediction.
 1 Introduction
-Conformal prediction is a versatile framework for quantifying the uncertainty of any black-box prediction model, by issuing a prediction set that covers an unknown outcome with a prescribed probability.
+Conformal prediction is a versatile framework for quantifying the uncertainty of any black-box prediction model, by issuing a prediction set that covers the unknown outcome with a prescribed probability.
 Formally, suppose the task is to predict an outcome $Y\in{\mathcal{Y}}$ based on features $X\in{\mathcal{X}}$.
@@
 The probability in (1) is over both the calibration data and the test point (Vovk et al.,, 2005; Lei et al.,, 2018).
-With this finite-sample, distribution-free guarantee, the conformal prediction set $\widehat{C}_{\alpha,n+1}$ describes a range of plausible values the unknown outcome $Y_{n+1}$ can take, thereby expected to inform downstream decision-making based on any black-box prediction model.
-With such a promise, methods for issuing marginally-valid – in the sense of (1) – prediction sets have been developed for various problems; see e.g., Angelopoulos and Bates, (2021) for a review.
+With this finite-sample, distribution-free guarantee, the conformal prediction set $\widehat{C}_{\alpha,n+1}$ describes a range of plausible values the unknown outcome $Y_{n+1}$ may take, thereby expected to inform downstream decision-making based on the black-box prediction model.
+With such a promise, methods for constructing marginally valid – in the sense of (1) – prediction sets have been developed for various problems; see e.g., Angelopoulos and Bates, (2021) for a review.
 In many downstream applications, however, people are often only interested in a selective subset of units.
-For example, practitioners may act upon a unit if it exhibits an interesting property (Olsson et al.,, 2022; Sokol et al.,, 2024; Levitskaya,, 2023); in some cases, practitioners may want to focus on a subset suggested by some complicated deta-dependent decision process such as resource optimization (Svensson et al.,, 2018; Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014).
-It would be misleading if the prediction sets fail to deliver the promised coverage guarantee for the selected units.
+For example, practitioners may only act upon a unit if it exhibits an interesting property (Olsson et al.,, 2022; Sokol et al.,, 2024; Levitskaya,, 2023), or focus only on a subset of test units picked by a complicated data-dependent process such as resource optimization (Svensson et al.,, 2018; Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014).
+It would be misleading for the practitioners if the prediction sets fail to deliver the promised coverage guarantee for the selected unit(s).
 Let us discuss a few applications where such cases may arise.
-  • In drug discovery, an important task is to predict the binding affinity of a drug candidate to a target to inform drug prioritization (Laghuvarapu et al.,, 2024).
-Scientists may only focus on drug candidates with highest predicted affinites, or selected by an false discovery rate (FDR)-controlling procedure (Jin and Candès, 2023b, ), or whose prediction sets only cover high values (Svensson et al.,, 2017), or that optimize resource usage (Svensson et al.,, 2018).
-It will be misleading if more than an $\alpha$-fraction of prediction sets for the selected drugs fail to cover the actual binding affinities.
+  • In drug discovery, an important task is to predict the binding affinity of a drug candidate to a disease target, which informs subsequent drug prioritization (Laghuvarapu et al.,, 2024).
+Among many drug candidates, scientists may only focus on those with highest predicted affinities, or those selected by a false discovery rate (FDR)-controlling procedure (Jin and Candès, 2023b, ), or whose prediction sets only cover high values (Svensson et al.,, 2017), or those optimizing resource usage (Svensson et al.,, 2018).
+It may lead to a waste of resources if the prediction sets for the selected drugs fail to cover the actual binding affinities with an exceedingly high chance.
   • In business decision-making, companies may take different inventory decisions based on whether a conformal prediction set suggests a strong demand or a weak demand (Levitskaya,, 2023).
-Similarly, it will be problematic if the average coverage among strong-demand predictions is below $1-\alpha$.
+Similarly, it will be problematic if a strong-demand prediction cannot cover with at least $1-\alpha$ of the time.
   • In disease diagnosisOlsson et al., (2022) suggest human intervention if the prediction set for a disease status is too large (this implicitly declares confidence in small prediction sets; see similar ideas in Ren et al., (2023); Sokol et al., (2024)).
-However, it would be concerning if more than an $\alpha$-fraction of small-sized prediction sets (“approved” as confident) miss the true disease status.
-  • In healthcare management, clinical decision makers may send patients to different healthcare categories after optimizing a performance measure (such as waiting time) subject to certain constraints, such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014).
-It is thus natural to only include a data-dependent subset of patients in downstream uncertainty quantification.
-In all these cases, it is highly desirable that a prediction set should cover the unknown outcome for a unit of interest with a prescribed probability.
+However, it would be concerning if, with more than a chance $\alpha$, the small-sized prediction sets – “approved” as confident – miss the true disease status.
+  • In healthcare management, patients may be sent to different healthcare categories based a program that optimizes some performance measure (such as waiting time) subject to certain constraints, such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014).
+The subset of patients in each category is therefore data-dependent.
+In all these examples, it is highly desirable that a prediction set should cover the unknown outcome for a unit of interest with a prescribed probability.
 This motivates a stronger, selection-conditional guarantee.
-Supposing there are $m\geq 1$ test units $\{X_{n+j}\}_{j=1}^{m}$, we shall aim for
+Supposing there are $m\geq 1$ test units ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$, we aim for
 $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\big{)}\geq 1-\alpha,$ (2)
-where ${\mathcal{D}}_{\textnormal{calib}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is the calibration data, ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ is the test data, and ${\mathcal{S}}(\cdot,\cdot)\subseteq[m]:=\{1,\dots,m\}$ is a data-driven process to decide which units are of interest.
-In particular, we focus on the case where $\widehat{C}_{\alpha,n+j}$ is constructed after the selection set ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is determined.
-This is similar to post-selection inference (Lee et al.,, 2016; Markovic et al.,, 2017; Tibshirani et al.,, 2018), yet we study a predictive inference problem; see Section 1.2 for more discussion.
-A prediction set with marginal validity (1) does not necessarily cover an unknown outcome conditional on being of interest as (2).
-Such selection issue has caught recent attention in the literature of predictive inference.
-Through analysis of a real drug discovery dataset, Jin and Candès, 2023b (, Section 1) demonstrate that more than $30\%$ of seemingly promising prediction sets (in the sense that $\widehat{C}_{\alpha}(X_{n+1})=\{1\}$) miss the actual outcomes for $\alpha=0.01$, i.e., with a marginal coverage of $99\%$.
-We shall see more examples of this issue in our numerical experiments under various selection processes.
-Focused on downstream decisions, Jin and Candès, 2023b ; Jin and Candès, 2023a develop methods that select units with large outcomes while guranteeing control of FDR.
-For uncertainty quantification, Bao et al., (2024) propose methods to construct prediction sets for selected test units with (approximately) false coverage rate (FCR) control.
-Their methods only achieve exact guarantees for a limited class of selection rules such as top-$K$ selection, while we are to present a general framework that works for arbitrary selection rules.
-In addition, FCR control and selection-conditional coverage (2) are related but distinct error notions, which we discuss in Section 2.2.
+where ${\mathcal{D}}_{\textnormal{calib}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is the calibration data, and ${\mathcal{S}}(\cdot,\cdot)$ is a data-driven process to decide the units of interest, which maps the calibration and test data to a subset of $[m]:=\{1,\dots,m\}$.
+Our target is similar in spirit to post-selection inference (Lee et al.,, 2016; Markovic et al.,, 2017; Tibshirani et al.,, 2018), but we consider predictive inference settings and develop distinct techniques; see Supplementary Section A.1 for more discussion.
+Throughout, we focus on settings where the prediction sets are constructed after ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is determined.
+By separating the selection process from the (post hoc) uncertainty quantification step, we leave the freedom of defining selection rules to the practitioners.
+A prediction set with marginal validity (1) does not necessarily cover an unknown outcome conditional on being of interest as in (2).
+Such a selection issue has recently been raised in the literature of predictive inference: through analysis of a real drug discovery dataset, Jin and Candès, 2023b (, Section 1) demonstrate that more than $30\%$ of seemingly promising prediction sets (in the sense that $\widehat{C}_{\alpha,n+j}=\{\text{active}\}$) miss the actual outcomes when the nominal marginal miscoverage rate is $\alpha=0.01$.
+We shall see more examples of this issue in our numerical experiments with various selection processes.
+New techniques are therefore needed for constructing prediction sets achieving (2).
+1.1 Exchangeability via reference sets
+Accounting for selection in conformal prediction is a delicate task since it breaks exchangeability.
+The core of conformal prediction is to leverage the exchangeability among the calibration and test data, such that their “prediction residuals”, referred to as nonconformity scores, are comparable in distribution (see Section 3.1 for more details).
+The calibration scores therefore inform the magnitude of uncertainty in a new test point (Vovk et al.,, 2005).
+However, given a selection event, the calibration data are no longer exchangeable with the test point, leading to violation of the coverage guarantee mentioned above.
+In general, the selection-conditional distributions of these scores are complex since the selection event can be highly data-dependent.
+Such a challenge motivates our new framework, named JOint Mondrian Conformal Inference (JOMI), which builds prediction sets that achieve selection-conditional coverage, with (2) as a special case.
+As visualized in Figure 1, our key idea is to find a “reference set” — a data-dependent subset of calibration data that remain exchangeable with respect to the new test point conditional on the selection event.
+This reference set thus provides calibrated quantification of uncertainty for a selected unit.
+The mechanism we devise accommodates arbitrary selection rules ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that are invariant to permutations of data in ${\mathcal{D}}_{\textnormal{calib}}$, and can be computed efficiently for a wide range of commonly used selection rules.
+Figure 1: Visualization of the intuition behind the reference set.
+(a) Marginally, the calibration data (black) are exchangeable with respect to the test point (blue).
+(b) The calibration data are not exchangeable with respect to the test point (shaded blue) given a selection event.
+(c) We find calibration data which, when posited as a “test point”, would lead to the same selection event.
+(d) The reference set consists of calibration data that are exchangeable with respect to the test point given selection, and we use them to construct JOMI prediction sets.
 We also note that the selection-conditional coverage (2) may be implied by other stronger notions such as conditional coverage: $\mathbb{P}(Y_{n+1}\in\widehat{C}_{\alpha,n+1}{\,|\,}X_{n+1}=x)\geq 1-\alpha$ for $\mathbb{P}$-almost all $x\in{\mathcal{X}}$.
-Conditional coverage implies (2) if $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are mutually independent and the selection rule only depends on the features.
-However, it is not achievable by finite-length prediction intervals in the absence of distributional assumptions (Foygel Barber et al.,, 2021), and the underlying selection rule may be more complicated.
-In contrast, we may view (2) as a “relaxed” version of conditional coverage that is both achievable and relevant for practical use.
-1.1 Overview of contributions
-In this paper, we propose a unified framework, named JOint Mondrian Conformal Inference (JOMI), that builds valid prediction sets for focal units picked by any data-dependent selection process.
-To this end, we make the following contributions.
-In Section 3, we introduce the general formulation of JOMI, whose crux is to find a data-dependent “reference” set of calibration data that are exchangeable conditional on the selection event.
-In this sense, our method can be viewed as a generalization of Mondrian Conformal Prediction (Vovk et al.,, 2005).
-We prove that for exchangeable data, the prediction set $\widehat{C}_{\alpha,n+j}$ produced by JOMI covers $Y_{n+j}$ with probability at least $1-\alpha$ conditional on a selection event.
-The selection event can be “test unit $j$ is selected”, which leads to (2).
-It can also be more granular, such as “test unit $j$ is selected, and there are $k$ selected test units”.
-Our framework works for any selection rule ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that is permutation-invariant to ${\mathcal{D}}_{\textnormal{calib}}$, without any modeling assumptions on the data generating process.
-In Section 4, we work out efficient implementation of JOMI for a number of selection processes that may be of practical interest, including:
+Conditional coverage implies (2) if $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are mutually independent and the selection rule only depends on test features.
+However, it is not achievable by finite-length prediction intervals without distributional assumptions (Vovk et al.,, 2005; Barber et al.,, 2021), and practical selection rules may also depend on other information.
+Given these considerations, we may view (2) as a “relaxed” version of conditional coverage that is both achievable and relevant for practical use.
+1.2 Preview of contributions
+In Section 3, we introduce the general formulation of JOMI, including the construction of reference set and its use in deriving a selection-conditionally valid prediction set.
+We prove that under exchangeability, the prediction set $\widehat{C}_{\alpha,n+j}$ produced by JOMI covers $Y_{n+j}$ with probability at least $1-\alpha$ conditional on a selection event.
+The selection event can be “test unit $j$ is selected”, which leads to (2); it can also be more granular, such as “test unit $j$ is selected, and there are $k$ selected test units”.
+This framework is valid for any selection rule ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that is permutation-invariant to ${\mathcal{D}}_{\textnormal{calib}}$, without any modeling assumptions on the data generating process.
+In Section 4, we study the computational aspect of JOMI.
+We show that when $|{\mathcal{Y}}|<\infty$, our method can be computed with a worst-case complexity of $O(|{\mathcal{Y}}|mn)$ times that of the selection rule.
+Moreover, for general continuous outcome space ${\mathcal{Y}}$, we work out efficient implementation of JOMI for a number of selection processes that may be of practical interest, including:
   • Covariate-dependent selection.
-When the selection rule do not involve $\{Y_{i}\}_{i=1}^{n}$, the computation complexity of our method is (at most) $O(mn)$ times that of the selection rule.
-This imples efficient implementation for a wide range of problems, including various forms of top-K selection and selecting based on complicated constrained optimization programs.
-As a special case, we recover the method of Bao et al., (2024) for top-K selection among test data, and yield valid solutions to other ranking-based selection methods attempted in their work.
+When the selection rule does not involve $\{Y_{i}\}_{i=1}^{n}$, the computation complexity of our generic method is (at most) $O(mn)$ times that of the selection rule.
+This implies efficient implementation for a wide range of problems, including various forms of top-K selection (for which the computation can be further reduced to $O(m+n)$) and selection based on complicated constrained optimization programs.
+As a special case, we recover the method of Bao et al., (2024) for top-K selection among test data, and provide valid solutions to other ranking-based selection methods attempted in their work.
   • Conformal p-value-based selection.
-We also derive efficient implementation for selection rules based on conformal p-values with both fixed and data-dependent thresholds.
-For instance, our prediction set is finite-sample exactly valid for units selected by Conformalized Selection (Jin and Candès, 2023b, ), which is studied in Bao et al., (2024) with approximate guarantee.
-  • Selection based on preliminary conformal prediction set.
-In addition, we derive a general, efficient instantiation when unit(s) whose marginal prediction sets demonstrate an interesting property are selected, such as being of a short length, or having a lower bound above a threshold.
-Our methods can be useful for re-calibrating the uncertainty quantification of such seemingly promising units.
-Finally, we demonstrate the applications of our methods via a few realistic selection rules that may occur in drug discovery (Section 5) and health risk prediction (Section 6).
-Our results show that marginally valid prediction sets may undercover or overcover, while our methods lead to valid coverage.
-1.2 Related work
-Our methods extend conformal prediction, which is a general framework for building marginally valid prediction intervals (Vovk et al.,, 2005; Lei et al.,, 2018; Romano et al.,, 2019, 2020).
-Compared with vanilla conformal prediction, we aim at a new goal, which is to achieve selection-conditional coverage for a unit of interest, instead of marginally for an “average” unit.
-Particularly related to our framework is Mondrian Conformal Prediction (Vovk et al.,, 2003, MCP), which uses a subset of calibration data to build a prediction interval for coverage conditional on a pre-determined or permutation-invariant class membership.
-However, achieving selection-conditional coverage is a new goal with distinct motivations, and our techniques and scopes of application are significantly different (see Remark 1).
-The selection-conditional guarantee we seek for is closely related to post-selection inference (POSI) literature (Berk et al.,, 2013; Tibshirani et al.,, 2016; Chernozhukov et al.,, 2015) and in particular conditional selective inference (Lee et al.,, 2016; Markovic et al.,, 2017; Reid et al.,, 2017; Tibshirani et al.,, 2018; Kivaranovic and Leeb,, 2020; McCloskey,, 2024).
-In POSI, one important task is to build confidence intervals for a selected subset of model parameters conditional on the selection event, which is close to our goal of selection-conditional coverage guarantee.
-Methods for this goal usually leverage specific problem structure such as linearity and distribution of the estimators to derive the confidence intervals (Lee et al.,, 2016; Tian and Taylor,, 2017; Andrews et al.,, 2024; Liu,, 2023).
-As we focus on predictive inference (we have units instead of parameters), we leverage the exchangeability across units to achieve valid selection-conditional coverage, hence our methods are very different.
-Another related error notion in the selective inference literature is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005); we compare them in more details in Section 2.2.
-Methods for achieving FCR control have been proposed for various settings (Benjamini and Yekutieli,, 2005; Weinstein et al.,, 2013; Weinstein and Ramdas,, 2020; Zhao and Cui,, 2020).
-For conformal prediction, Bao et al., (2024) propose prediction sets for selected units while achieving (approximate) FCR control.
-However, their methods only work for multiple test units.
-In addition, they achieve exact guarantee only for selection rules that are invariant to permutations of calibration and test units and top-$K$ selection (where $K$ test units with the highest predictions are selected); we will show that when applying JOMI to these problems, we yield the same prediction sets as theirs.
-Meanwhile, our framework leads to exact guarantee for some other problems studied in Bao et al., (2024) with approximate FCR control.
-Our framework also tackles many more settings with complicated selection rules that cannot be handled by theirs.
-Selective inference issues in conformal prediction or predictive inference have caught recent attention.
-Most of existing works study multiple testing, i.e., selecting individuals that obey some properties with false discovery rate (FDR) control, including selecting outliers (Bates et al.,, 2023; Marandon et al.,, 2022; Liang et al.,, 2022) and selecting high-quality data (Jin and Candès, 2023b, ; Jin and Candès, 2023a, ).
-The construction of conformal prediction sets is an orthogonal direction, and our methods also work for selection based on such procedures (we develop an efficient method in Section 4.2.2).
-Under a similar name, Sarkar and Kuchibhotla, (2023) proposed “post-selection inference for conformal prediction”, but they mainly focus on selecting the confidence level $\alpha\in(0,1)$ instead of selecting the units.
-In addition, they fall within the “simultaneous inference” strand in POSI so that their inference is valid simultaneously for all confidence levels, other than conditional on a selection event.
-For these reasons, our methods and guarantees are quite different.
+We derive efficient implementation for a class of selection rules based on thresholding conformal p-values with “stopping time-type” cutoffs.
+For instance, our prediction set is finite-sample exactly valid for units selected by the Conformal Selection method (Jin and Candès, 2023b, ), which is studied in Bao et al., (2024) with approximate FCR guarantee.
+  • Selection based on preliminary conformal prediction sets.
+In addition, we derive a general, efficient instantiation when selecting units whose marginal prediction sets demonstrate certain interesting properties, such as being of a short length, or having a lower bound above some threshold.
+Our method can be useful for re-calibrating the uncertainty quantification of such seemingly promising units.
+Finally, we demonstrate the application of our methods via several realistic selection rules that may occur in drug discovery (Section 5) and health risk prediction (Section 6).
+Our results show that marginal prediction sets may undercover or overcover the selected units, while our methods always achieve the promised coverage guarantees.
+Due to space constraints, we delegate a comprehensive literature review to Supplementary Section A.1. We end this section with some useful notations.
+Notations.
+For a positive integer $n\in\mathbb{N}^{+}$, write $[n]:=\{1,\dots,n\}$.
+We write the data pair as $Z=(X,Y)$, so that the calibration data is ${\mathcal{D}}_{\textnormal{calib}}=\{Z_{i}\}_{i=1}^{n}$.
+For any $j\in[m]$, we define the augmented calibration set as ${\mathcal{D}}_{j}={\mathcal{D}}_{\textnormal{calib}}\cup\{Z_{n+j}\}$ and the remaining test set as ${\mathcal{D}}^{c}_{j}={\mathcal{D}}_{\textnormal{test}}\backslash\{X_{n+j}\}$.
+The unordered set of ${\mathcal{D}}_{j}$ is denoted $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$, which provides the order statistics.
 2 Problem setup
-We are interested in inferring the unknown outcome $Y$ of certain “focal” units selected in a data-driven manner.
-Following the split conformal prediction framework (Vovk et al.,, 2005; Lei et al.,, 2018), our prediction set is based on a fitted prediction model obtained from a training set ${\mathcal{D}}_{\textnormal{train}}$.
-Recall the calibration data ${\mathcal{D}}_{\textnormal{calib}}=\{Z_{i}\}_{i=1}^{n}$, $Z_{i}=(X_{i},Y_{i})$, and test points $\{Z_{n+j}\}_{j=1}^{m}$, $Z_{n+j}=(X_{n+j},Y_{n+j})$.
-The outcomes $\{Y_{n+j}\}_{j=1}^{m}$ are unobserved and of interest.
-We write ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ for brevity.
-We assume ${\mathcal{D}}_{\textnormal{train}}$ is independent of ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, and we condition on ${\mathcal{D}}_{\textnormal{train}}$, treating the fitted models as fixed.
+Following the split conformal prediction framework (Vovk et al.,, 2005; Lei et al.,, 2018), we build our prediction sets based on a prediction model fitted on a training set ${\mathcal{D}}_{\textnormal{train}}$, assuming ${\mathcal{D}}_{\textnormal{train}}$ is independent of the calibration and test data.
+In what follows, we always condition on ${\mathcal{D}}_{\textnormal{train}}$, thereby treating the fitted models as fixed.
+In this section, we formally introduce the selection-conditional coverage guarantees and compare them to other related notions in the literature.
 2.1 Selection-conditional coverage
-We call $Z_{n+j}$ a focal unit if $j\in\widehat{\mathcal{S}}$, where $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\subseteq[m]$ is obtained from a selection rule ${\mathcal{S}}$ that may depend on the observe data.
+Recall that $j\in[m]$ is a focal unit if $j\in\widehat{\mathcal{S}}$, where $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\subseteq[m]$ is obtained from a selection rule ${\mathcal{S}}$ that depends on the observed data.
 Its potential dependence on ${\mathcal{D}}_{\textnormal{train}}$ is clear since we treat ${\mathcal{D}}_{\textnormal{train}}$ as fixed.
-Without loss of generality, we posit that ${\mathcal{S}}$ is a determinististic function; when the selection rule is randomized, one can condition on its randomness and follow our framework.
+Without loss of generality, we posit that ${\mathcal{S}}$ is a deterministic function; when the selection rule is randomized, one can condition on its randomness and follow our framework.
 For each test unit $j$, we wish to construct a prediction set $\widehat{C}_{\alpha,n+j}\subseteq{\mathcal{Y}}$ such that
 $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\geq 1-\alpha,$ (3)
-where $\alpha\in(0,1)$ is the confidence level, and $\mathfrak{S}\subseteq 2^{[m]}$ is some pre-specified collection of subsets of $[m]$. We call $\mathfrak{S}$ the selection taxonomy in what follows.
-Equation (3) is the most general form of the selection-conditional coverage we aim for in this paper, where difference choices of $\mathfrak{S}$ leads to a spectrum of granularity in the conditioning event.
-For instance, taking $\mathfrak{S}=\{S_{0}\}$ for some $S_{0}\subseteq[m]$ achieves coverage conditional on selecting a specific set $S_{0}$; this is similar to the coverage guarantee in (Lee et al.,, 2016) for high-dimensional inference.
-Taking $\mathfrak{S}=\{S\subseteq[m]\colon|S|=r\}$ for some $r\leq m$ achieves coverage conditional on getting a specific number of selected units.
-Finally, taking $\mathfrak{S}=2^{[m]}$ puts no restrictions on the selection set, giving rise to the guarantee (2) introduced in the beginning, coverage conditional only on a unit being selected.
+where $\alpha\in(0,1)$ is the confidence level, and $\mathfrak{S}\subseteq 2^{[m]}$ is some pre-specified collection of subsets of $[m]$.
+We call $\mathfrak{S}$ the selection taxonomy in what follows.
+Different choices of $\mathfrak{S}$ lead to a spectrum of granularity in the conditioning event.
+For instance, taking $\mathfrak{S}=2^{[m]}$ puts no restrictions on the selection set, giving rise to the guarantee (2) introduced in the beginning (coverage conditional on a unit being selected).
+Taking $\mathfrak{S}=\{S\subseteq[m]\colon|S|=r\}$ for some $r\leq m$ achieves coverage conditional on selecting a specific number of units.
+Finally, taking $\mathfrak{S}=\{S_{0}\}$ for some $S_{0}\subseteq[m]$ achieves coverage conditional on selecting a specific set; this is similar to the coverage guarantee conditional on a selected model in Lee et al., (2016) for high-dimensional parameter inference.
+In words, the selection-conditional coverage guarantee (3) can be interpreted as follows: imagine there are infinitely many independent realizations of ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$; among those realizations where the selection event of interest happens, the prediction set $\widehat{C}_{\alpha,n+j}$ will cover the true outcome for at least $1-\alpha$ fraction of times.
+Here, $\mathfrak{S}$ is introduced to provide practitioners with the language to specify the granularity of the conditioning event, allowing them to tailor the guarantee to their specific needs.
+Moreover, it allows us to properly describe the relationship between selection-conditional coverage and FCR control, as we will see below.
 2.2 Relations among notions of selective coverage
-Before introducing our methods, we take a moment here to compare different notions of selective coverage. Readers who are more interested in our methodology may skip the remaining of this section.
-Our first observation is that (3) implies (2) under appropriate conditions. The proof of the next Proposition is in Appendix B.1.
+Before introducing our methods, we take a moment to compare different notions of selective coverage. Readers who are more interested in our methodology may skip the remaining of this section.
+Our first observation is that (3) implies (2) under appropriate conditions. The proof of the next proposition is in Supplementary Section D.1.
 Proposition 1.
-If $\widehat{C}_{\alpha,n+j}$ satisfies (3) for a set of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, then it satisfies (2).
-To distinguish the selection-conditional coverage in (3) and that in (2), we refer to the former as strong selection-conditional coverage and the latter as weak selection-conditional coverage.
-Another widely used post-selection-type guarantee is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), defined as the expected proportion of selected units missed by the prediction set:
+Suppose a family of prediction sets $\{\widehat{C}_{\alpha,n+j}^{(\ell)}\}_{\ell\in{\mathcal{L}}}$ satisfy $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$.
+Define the prediction set $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ when $j\in\widehat{\mathcal{S}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$.
+Then $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}})\geq 1-\alpha$.
+To distinguish the selection-conditional coverage in (3) and that in (2), we refer to (3) as strong selection-conditional coverage and (2) as weak selection-conditional coverage.
+Another widely used post-selection guarantee is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), defined as the expected proportion of selected units missed by the prediction set:
 $\text{FCR}\,:=\,\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]},$ (4)
 where $a\vee b=\max\{a,b\}$ for any $a,b\in\mathbb{R}$.
-As introduced earlier, Bao et al., (2024) mainly focuses on constructing prediction sets with FCR control.
+Previous works (Weinstein and Ramdas,, 2020; Bao et al.,, 2024; Gazin et al.,, 2024) mainly focus on constructing prediction sets with FCR control.
 However, with $m=1$, it holds that $\text{FCR}=\mathbb{P}(j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j}))\leq\alpha$ for any marginally valid prediction set.
-Thus, the FCR is not very meaningful when we consider one test point.
-We claim that the strong selection-conditional coverage implies FCR control with a proper choice of the selection taxonomy. The proof of the following proposition can be found in Appendix B.2.
+Thus, FCR does not always address the selection issue.
+The following proposition shows that strong selection-conditional coverage implies FCR control with a proper choice of selection taxonomy, whose proof is in Supplementary Section D.2.
 Proposition 2.
-[…]
-As a side note, we find that the weak selection-conditional coverage does not necessarily imply FCR control, although in some special cases both can be true (see, e.g., Bao et al., (2024) for such examples).
-We put this as a proposition below, with a counterexample given in Appendix B.3.
+Suppose a family of prediction sets $\{\widehat{C}_{\alpha,n+j}^{(\ell)}\}_{\ell\in{\mathcal{L}}}$ satisfy $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of disjoint selection taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$.
+Define the prediction set $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ when $j\in\widehat{{\mathcal{S}}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$.
+Then its FCR (4) is upper bounded by $\alpha\cdot\mathbb{P}(\widehat{{\mathcal{S}}}\neq\varnothing)\leq\alpha$.
+The selection taxonomies in Proposition 2 require the selection set to be of a specific size, without other conditions on its form.
+This can be automatically satisfied by some selection rules such as top-K selection.
+Extending Proposition 2 to non-exact scenarios yields useful insights in practice.
+For instance, consider the family of taxonomies with $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon r(\ell)-\delta\leq|S|\leq\delta+r(\ell)\}$ for some small $\delta\in\mathbb{N}^{+}$.
+When the selection size is stable, weak selection-conditional coverage implies $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\approx 1-\alpha$, which leads to $\textrm{FCR}\approx\alpha\cdot\mathbb{P}(\widehat{{\mathcal{S}}}\neq\varnothing)$.
+This approximate result helps explain several phenomena in our numerical experiments: first, the FCR is usually controlled empirically by our method; second, we sometimes observe significantly lower FCR than the selection-conditional coverage when $\mathbb{P}(\widehat{{\mathcal{S}}}=\varnothing)$ is moderately large.
+As we shall see later, a version of our proposed method achieves coverage guarantees conditional on being selected and the selection size, which requires checking a slightly more complex condition in the reference set construction when the selection size varies.
+The computation of such prediction sets can be done efficiently for all the instances provided in the paper, with the corresponding worst-case computational complexity explicitly stated in each section.
+One potential concern, though, is that the additional condition on the selection size may reduce the number of calibration data points in the reference set, especially when the selection step is highly variable.
+This could potentially lead to wider and/or unstable prediction intervals.
+As a side note, the weak selection-conditional coverage does not necessarily imply FCR control, although in some special cases both can be true (see, e.g., Bao et al., (2024) for such examples).
+We put this as a proposition below, with a counterexample given in Supplementary Section D.3.
 Proposition 3.
-There exists an instance and prediction sets $\widehat{C}_{\alpha}$ such that $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage at level $\alpha$ but violates the FCR control at level $\alpha$.
-We close this part by introducing several notations that will be used throughout the paper.
-Notations.
-We denote the index sets ${\mathcal{I}}_{{\textnormal{calib}}}=\{1,\dots,n\}$ and ${\mathcal{I}}_{\textnormal{test}}=\{n+1,\dots,n+m\}$.
-Let $S:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ be the score function for selection, and $V:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ the nonconformity score function for constructing the prediction set.
-For example, $S(x,y)=y$ and $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$.
-For any $j\in{\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, define $S_{i}=S(X_{i},Y_{i})$ and $V_{j}=V(X_{j},Y_{j})$.
-Let ${\mathcal{D}}_{j}={\mathcal{D}}_{\textnormal{calib}}\cup\{Z_{n+j}\}$, and ${\mathcal{I}}_{j}={\mathcal{I}}_{\textnormal{calib}}\cup\{n+j\}$ be the corresponding indices; define also ${\mathcal{D}}^{c}_{j}={\mathcal{D}}_{\textnormal{test}}\backslash\{Z_{n+j}\}$ and ${\mathcal{I}}^{c}_{j}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$.
-For any $j\in[m]$, we define $Z_{n+j}(y)=(X_{n+j},y)$, the value of $Z_{n+j}$ if $Y_{n+j}$ were to take a hypothesized value $y\in{\mathcal{Y}}$.
+There exists an instance and prediction sets $\{\widehat{C}_{\alpha,n+j}:j\in\widehat{{\mathcal{S}}}\}$ that satisfy the weak selection-conditional coverage at level $\alpha$ but violate the FCR control at level $\alpha$.
+We end this section with a remark on the interpretation of selection-conditional coverage and FCR control.
+Remark 1 (Interpretation of selection-conditional coverage and FCR control).
+As shown by Proposition 2, the guarantee in (3) is stronger than FCR control (for a proper choice of the selection taxonomy).
+In fact, the former has often been used as a device to derive the latter in the literature (e.g., Weinstein et al.,, 2013; Bao et al.,, 2024).
+In terms of interpretation, there are two main differences between selection-conditional coverage and FCR.
+First, the guarantee of FCR is averaged over all the selection events, including the case of empty selection set where the false coverage proportion is by definition zero.
+Therefore, even with FCR control, one could still suffer from a high false coverage proportion when the selection set is non-empty as long as this is compensated by the cases where the selection set is empty.
+On the other hand, selection-conditional coverage delivers guarantees conditioning on selection events of interest, which prevents the aforementioned undesired situation.
+Second, FCR control provides a guarantee that is averaged over all the selected units: it could be possible that for some selected units, the coverage is much lower than the nominal level, and for others the coverage is much higher, so that the average coverage over all the selected units is controlled at the nominal level.
+In contrast, selection-conditional coverage provides guarantees specific to each selected units.
+Finally, we note that the distinction between the two concepts can be asymptotically negligible in special cases. We refer to Lemma 1 in the supplementary material for such an instance.
 3 JOMI: a unified framework
 3.1 Warm-up: split conformal prediction
-To warm up, we briefly summarize the split conformal prediction (SCP) method (Vovk et al.,, 2005; Lei et al.,, 2018) and the ideas for its finite-sample coverage due to exchangeability.
-We begin with a nonconformity score function $V\colon{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ determined by the training set, so that $V(x,y)$ informs how well a hypothetical value $y\in{\mathcal{Y}}$ conforms to a machine prediction.
+To warm up, we briefly summarize the split conformal prediction (SCP) method (Vovk et al.,, 2005; Lei et al.,, 2018) and how it achieves finite-sample coverage under exchangeability.
+SCP starts with a nonconformity score function $V\colon{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ determined by ${\mathcal{D}}_{\textnormal{train}}$, so that $V(x,y)$ informs how well a hypothetical value $y\in{\mathcal{Y}}$ conforms to a machine prediction.
 For instance, one may set $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$.
-Other popular choices include CQR (Romano et al.,, 2019) for regression and APS (Romano et al.,, 2020) for classification.
-We compute $V_{i}=V(X_{i},Y_{i})$ for $i\in[n]$.
-The split conformal prediction set for a test point $j\in[m]$ is then
+Other popular choices include conformalized quantile regression (CQR, Romano et al.,, 2019) for regression and adaptive prediction sets (APS, Romano et al.,, 2020) for classification.
+Compute $V_{i}=V(X_{i},Y_{i})$ for $i\in[n]$. The split conformal prediction set for test unit $j\in[m]$ is
 $\widehat{C}_{\alpha,n+j}^{\text{SCP}}=\Big{\{}y\colon V(X_{n+j},y)\leq\text{Quantile}\big{(}1-\alpha;\{V_{i}\}_{i=1}^{n}\cup\{+\infty\}\big{)}\Big{\}},$ (5)
-where $\text{Quantile}(1-\alpha;\cdot)$ is the empirical $(1-\alpha)$-th quantile of the set in the second argument.
-When $(Z_{1},\dots,Z_{n},Z_{n+j})$ are exchangeable, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ achieves valid marginal coverage (1(Vovk et al.,, 2005).
-It helps motivate our approach to first introduce the ideas behind the validity of $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$. Note that
+where $\text{Quantile}(1-\alpha;\cdot)$ is the $(1-\alpha)$-th empirical quantile of the set in the second argument.
+When $(Z_{1},\dots,Z_{n},Z_{n+j})$ are exchangeable, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ achieves (1(Vovk et al.,, 2005).
+It helps motivate our approach to see the ideas behind the validity of $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$.
+In words, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ finds hypothesized values of $y$ that make $V(X_{n+j},y)$ look similar to calibration scores.
+Note that
 $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\text{SCP}}\big{)}=\mathbb{P}\Big{(}V_{n+j}\leq\text{Quantile}\big{(}1-\alpha;\{V_{i}\}_{i=1}^{n}\cup\{V_{n+j}\}\big{)}\Big{)},$ (6)
 where $V_{n+j}=V(X_{n+j},Y_{n+j})$.
-Define the unordered set $[Z]_{j}=[Z_{1},\dots,Z_{n},Z_{n+j}]$.
-Due to exchangeability, conditional on the event $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, the probability of $Z_{n+j}$ taking on each value in $z_{1},\dots,z_{n+j}$ is equal.
-The right-handed side of (6) is thus lower bounded by $1-\alpha$.
-The key challenge of achieving selection-conditional coverage like (3) is that, conditional on the selection event, the data points $\{Z_{1},\dots,Z_{n},Z_{n+j}\}$ are not exchangeable any more.
-More specifically, given selection, and further conditional on $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, it is unclear how $Z_{n+j}$ is distributed over $z_{1},\dots,z_{n+j}$.
-As such, a natural idea is to find a subset of calibration data that are “exchangeable” with the new test point even given the selection event.
-We introduce our methods in the next.
-3.2 Conformal inference via locally exchangeable subgroups
+Recall the unordered set $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ where $Z_{i}=(X_{i},Y_{i})$.
+Conditional on the event $\{[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]\}$, the only randomness is in the ordering of $(Z_{1},\dots,Z_{n},Z_{n+j})$; due to exchangeability, the probability of $Z_{n+j}$ taking on each value in $z_{1},\dots,z_{n},z_{n+j}$ is equal.
+Therefore, conditional on $[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]$, the chance of $V_{n+j}$ being no greater than the $(1-\alpha)$-th quantile of $[v_{1},\dots,v_{n},v_{n+j}]$, where $v_{i}=v(z_{i})$, is at least $1-\alpha$, i.e.,
+$\mathbb{P}\Big{(}V_{n+j}\leq\text{Quantile}\big{(}1-\alpha;\{v_{i}\}_{i=1}^{n}\cup\{v_{n+j}\}\big{)}\,\Big{|}\,[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]\Big{)}\geq 1-\alpha.$ (7)
+This leads to (6) by the tower property. Therefore, inverting the criterion on the right-hand side of (6) gives a valid prediction set for $Y_{n+j}$.
+The reason why vanilla SCP may fail to achieve selection-conditional coverage like (3) is that, conditional on the selection event, the data points $\{Z_{1},\dots,Z_{n},Z_{n+j}\}$ are no longer exchangeable.
+In other words, in (7), it is unclear how $V_{n+j}$ is distributed over $v_{1},\dots,v_{n+j}$ if we additionally condition on the selection event.
+As such, a natural remedy is to find a subset of calibration data that are “exchangeable” with respect to the test point conditioning on the selection event, and leverage their scores to calibrate the prediction of the test unit.
+We introduce our methods below.
+3.2 Conformal inference via a reference set
 Fix a test unit $j\in[m]$.
-Recall that $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the focal set, and $Z_{n+j}(y)=(X_{n+j},y)$ is the imputed data point with a hypothesized response $y\in{\mathcal{Y}}$.
-For any calibration data $i\in{\mathcal{I}}_{\textnormal{calib}}$, we define the “swapped” calibration data ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)$ and test data ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}(y)$ to be
+Recall that $Z_{n+j}(y)=(X_{n+j},y)$ is the imputed test point with a hypothesized response $y\in{\mathcal{Y}}$.
+The core of our method is to find calibration points that are exchangeable with respect to the test point conditional on the selection set.
+At a high level, they are calibration units $i\in[n]$ that are “indistinguishable” with $j\in[m]$, in the sense that treating $Z_{n+j}(y)$ as a calibration point and $Z_{i}$ as a test point results in the same selection event.
+We formalize this idea via a “swap” operation.
+For any calibration unit $i\in[n]$, we define the “swapped” calibration data ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)$ and the swapped test data ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}$ as follows:
 ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)=(Z_{1}^{{\textnormal{swap}(i,j)}}(y),Z_{2}^{{\textnormal{swap}(i,j)}}(y),\dots,Z_{n}^{{\textnormal{swap}(i,j)}}(y)),$
 ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}=(X_{n+1}^{{\textnormal{swap}(i,j)}},X_{n+2}^{{\textnormal{swap}(i,j)}},\dots,X_{n+m}^{{\textnormal{swap}(i,j)}}),$
-where
-$Z_{\ell}^{{\textnormal{swap}(i,j)}}(y)=\begin{cases}Z_{n+j}(y)&\ell=i,\\ Z_{\ell}&\ell\neq i.\end{cases}\quad X_{n+\ell}^{{\textnormal{swap}(i,j)}}=\begin{cases}X_{i}&\ell=j,\\ X_{n+\ell}&\ell\neq j.\end{cases}$
+where for $k\in[n]$ and $\ell\in[m]$,
+$Z_{k}^{{\textnormal{swap}(i,j)}}(y)=\begin{cases}Z_{n+j}(y)&k=i,\\ Z_{k}&k\neq i.\end{cases}\quad X_{n+\ell}^{{\textnormal{swap}(i,j)}}=\begin{cases}X_{i}&\ell=j,\\ X_{n+\ell}&\ell\neq j.\end{cases}$
+That is, ${\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}(y)$ and ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)}$ are the calibration and test data if we treat $Z_{i}$ as the $j$-th test point, and $Z_{n+j}(y)$ as the $i$-th calibration point.
+Figure 2 is an illustration of the swap operation.
 […]
 Figure 2: Graphical illustration of ${\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}(y)$ and ${\mathcal{D}}_{\textnormal{test}}^{{\textnormal{swap}(i,j)}}$.
-That is, they are the calibration and test data if we treat $Z_{i}$ as the $j$-th test point, and $Z_{n+j}(y)$ as the $i$-th calibration point.
-Figure 1 is an illustration of the swapping operation.
-Applying the same selection rule, the selection set after swapping and with the hypothesized $y$ is then
+Applying the same selection rule ${\mathcal{S}}$ to the swapped data, we define the swapped selection set with the hypothesized $y$ as
 $\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y)={\mathcal{S}}\big{(}{\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y),{\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}\big{)}.$
 Then, we define the “reference set” for achieving (3) with taxonomy $\mathfrak{S}$ as
-^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}\colon j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y),\text{ and }\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}\big{\}}.$ (7)
-In words, the reference set is the collection of calibration points $i$ such that, after swapping unit $i$ and $n+j$, the test point $j$ remains in the focal set and the focal set remains in $\mathfrak{S}$.
-We use the notation ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ to emphasize that ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is a data-dependent mapping from ${\mathcal{Y}}$ to subsets of ${\mathcal{I}}_{{\textnormal{calib}}}$.
-With all the preparation, we define our prediction set via
+^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in[n]\colon j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y),\text{ and }\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}\big{\}}.$ (8)
+In words, the reference set is the collection of calibration points $i\in[n]$ such that, after swapping unit $i$ and $n+j$, the (posited) test point $j$ (which is the original unit $i$) remains in the focal set and the focal set remains in $\mathfrak{S}$.
+We use the notation ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ to emphasize that ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is a data-dependent mapping from ${\mathcal{Y}}$ to the power set of $[n]$.
+With all the preparation, we define our prediction set for $Y_{n+j}$ as
 $\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\cup\{+\infty\}\big{)}\Big{\}}.$ (9)
 As we shall show shortly, our prediction set (9) achieves near-exact coverage when the nonconformity score is continuous and the reference set is of moderate size.
-In fact, it is possible to achieve exact coverage by introducing extra randomness:
+In addition, we can achieve exact coverage by introducing extra randomness:
 $\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}$
-=$\quad=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}(y)|}\leq 1-\alpha\bigg{\}},$ (9)
-where $U_{1},\ldots,U_{m}$ are i.i.d. random variables drawn from $\textrm{Unif}[0,1]$ which are independent of the data.
-Remark 1 (Connection to MCP).
-Our framework is closely related to Mondrian Conformal Prediction (MCP(Vovk et al.,, 2003) which applies to one test unit ($m=1$).
-Given $K$ categories $[K]$, they assume access to a classifier ${\mathcal{K}}$ that gives the class label for all units: $(\kappa_{1},\kappa_{2},\dots,\kappa_{n+1})={\mathcal{K}}(Z_{1},Z_{2},\dots,Z_{n+1}),$ where $\kappa_{i}\in[K]$ is the class label for unit $i$.
-The classifier ${\mathcal{K}}$ is required to be invariant to any permutation $\pi$ of $[n+1]$, i.e., $(\kappa_{\pi(1)},\kappa_{\pi(2)},\dots,\kappa_{\pi(n+1)})={\mathcal{K}}(Z_{\pi(1)},Z_{\pi(2)},\dots,Z_{\pi(n+1)}).$ The goal of MCP is to construct a prediction set for the test unit with cateogory-wise coverage, i.e., $\mathbb{P}(Y_{n+1}\in\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}{\,|\,}\kappa_{n+1}=k)$ for each $k\in[K]$.
-This is achieved by finding a subset of calibration data for each hypothesized $y\in{\mathcal{Y}}$,
-~$\widetilde{{\mathcal{R}}}(y)=\{i\in{\mathcal{I}}_{\textnormal{calib}}:{\mathcal{K}}({\mathcal{D}}_{{\textnormal{calib}}},(X_{n+1},y))_{[i]}=k\},$ (10)
-and then constructing the prediction set as
-$\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+1},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widetilde{{\mathcal{R}}}(y)}\cup\{\infty\}\big{)}\Big{\}}.$
-While the setups seem different, JOMI effectively reduces to MCP when ${\mathcal{K}}$ is a binary classifier and ${\mathcal{S}}$ is equivariant under permutations of $[n+1]$.
-To see this, one can view the selection rule ${\mathcal{S}}$ as categorizing test points into two classes (selected and not selected), and selection-conditional coverage (3) as coverage given the “selected” category.
-In this way, one can check that our reference set ^$\widehat{{\mathcal{R}}}_{n+1}(y)$ coincides with ~$\widetilde{{\mathcal{R}}}(y)$ defined in (10), and thus the prediction sets also coincide.
-With a distinct goal of selection-conditional coverage, we additionally address the challenge that realistic selection rules are often asymmetric to calibration and (multiple) test points.
-To this end, JOMI generalizes MCP via a more delicate but still general framework for finding exchangeable subgroups.
-As we jointly consider the calibration and test points when defining the reference set, we call our method JOint Mondrain Conformal Inference (JOMI).
-3.3 Theoretical guarantee
-Theorem 1 confirms the validity of our prediction sets (8) and (3.2).
+$=$ $\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}(y)|}\leq 1-\alpha\bigg{\}},$ (10)
+where $U_{1},\ldots,U_{m}$ are i.i.d.
+random variables drawn from $\textrm{Unif}[0,1]$ independent of the data.
+In Supplementary Sections A.2 and A.3, we discuss in detail the connection of our method to Mondrian conformal prediction (MCP) and comparison with the BY procedure (Benjamini and Yekutieli,, 2005).
+While the BY procedure is a natural and heuristic solution to post-selection inference, it suffers from over-conservativeness and non-adaptivity to the selection event.
+3.3 Theoretical guarantees
+Theorem 1 confirms the conditional validity of our prediction sets $\widehat{C}_{\alpha,n+j}$ in (9) and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ in (3.2).
 Theorem 1.
-Suppose ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is invariant to permutations of ${\mathcal{D}}_{\textnormal{calib}}$.
-Suppose $\{Z_{i}\}_{i=1}^{n}\cup\{Z_{n+j}\}$ are exchangeable conditional on $\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$ for any $j\in[m]$.
-Then, for any selection taxonomy $\mathfrak{S}$,
+Suppose ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is invariant to permutations of ${\mathcal{D}}_{\textnormal{calib}}$, and that $\{Z_{i}\}_{i=1}^{n}\cup\{Z_{n+j}\}$ are exchangeable conditional on $\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$ for any $j\in[m]$.
+Then, for any selection taxonomy $\mathfrak{S}$, the following statements hold.
   1. (a)
@@
     $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\geq 1-\alpha.$ (11)
-    Furthermore, if ties between $V_{1},\ldots,V_{n},V_{n+j}$ occur with probability zero, then
+    Furthermore, if ties among $V_{1},\ldots,V_{n},V_{n+j}$ occur with probability zero, then
     $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}\,\big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}\big{)}\leq 1-\alpha+\mathbb{E}\bigg{[}\frac{1}{1+|\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})|}\bigg{]}.$ (12)
@@
     $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S})=1-\alpha.$
-We defer the detailed proof of Theorem 1 to Appendix B.4.
-Let us provide some intuitions for the validity results here.
-Similar to the ideas of SCP in Section 3.1, the key fact we rely on is that, plugging in the true value $y=Y_{n+j}$, data in the reference set and the new test point are still exchangeable conditional on the event that $j$ is selected.
-To be specific, imagine that we know the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]=[z_{1},\dots,z_{n},z_{n+j}]$ and the values of the remaining test points $\{X_{n+\ell}\}_{\ell\neq j}$.
-The only randomness is thus in the ordering of $Z_{1},\dots,Z_{n+j}$.
-Meanwhile, our reference set ^$\widehat{\mathcal{R}}_{j}(\cdot)$ is constructed in a delicate way such that the (unordered) scores in the reference set $[V_{i}\colon i\in\widehat{\mathcal{R}}_{j}(Y_{n+j})]$ is fully determined by the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]$ if ${\mathcal{S}}(\cdot,\cdot)$ is invariant to permutations of the “calibration” data in the first argument.
-Then, by the exchangeability of $Z_{1},\dots,Z_{n},Z_{n+j}$, the probability of $Y_{n+j}$ taking on any value in $\widehat{R}$ is equal given the unordered set, leading to the validity in Theorem 1.
+We defer the detailed proof of Theorem 1 to Supplementary Section D.4 and provide some intuition here.
+Similar to the ideas of SCP in Section 3.1, the key fact we rely on is that, plugging in the true value $y=Y_{n+j}$, data in the reference set and the new test point are still exchangeable conditional on the selection event.
+To be specific, to prove (11), we are to show a stronger result:
+$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{V_{n+j}\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S},\,[{\mathcal{D}}_{j}],\,{\mathcal{D}}_{j}^{c}\Big{)}\geq 1-\alpha,$ (13)
+where we recall $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ and ${\mathcal{D}}_{j}^{c}={\mathcal{D}}_{\textnormal{test}}\backslash\{X_{n+j}\}$.
+For any fixed values $z_{1},\dots,z_{n},z_{n+j}$, once given the unordered values $[{\mathcal{D}}_{j}]=[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$ and the values of other test points $\{X_{n+\ell}\}_{\ell\neq j}$, the only randomness is in the ordering of $Z_{1},\dots,Z_{n+j}$ among $[z_{1},\dots,z_{n},z_{n+j}]$.
+Meanwhile, we show that our reference set ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is constructed in a delicate way such that the (unordered set of) scores $\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{V_{n+j}\}$ is fully determined by $[z_{1},\dots,z_{n},z_{n+j}]$.
+That is, $\widehat{R}_{n+j}^{+}:=[V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}]$ is fully determined given $[d_{j}]$.
+Then, (13) reduces to
+$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\widehat{R}_{n+j}^{+}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S},\,[{\mathcal{D}}_{j}],\,{\mathcal{D}}_{j}^{c}\Big{)}\geq 1-\alpha,$ (14)
+Finally, by the exchangeability of $Z_{1},\dots,Z_{n},Z_{n+j}$, the probability of $V_{n+j}$ taking on any value in $\widehat{R}_{n+j}^{+}$ is equal given $[{\mathcal{D}}_{j}]$ and ${\mathcal{D}}_{j}^{c}$, leading to the validity in Theorem 1 via Bayes’ rule.
 4 Computationally tractable instances
-Section 3 provides a general framework for constructing valid prediction sets conditional on selection events, but computing the prediction sets according to their definitions requires looping over all possible values of $y$, often a computationally intractable task.
-In this section, we instantiate the generic procedure through concrete examples, in which special structures allow us to efficiently compute the prediction sets (8) and (3.2).
-We focus on three classes of selection rules: the selection rules using solely the covariates, those based on conformal p-values, and those based on conformal prediction sets.
+So far, we have presented a general framework for constructing valid prediction sets conditional on general selection events.
+However, computing the prediction sets according to their definition requires looping over all possible values of $y\in{\mathcal{Y}}$, which can be computationally intractable.
+When $|{\mathcal{Y}}|$ is finite (and relatively small), our proposed method can be efficiently implemented according to its definition; the corresponding computational complexity is at most $O(|{\mathcal{Y}}|mn)$ times the complexity of the selection rule.
+In this section, we instantiate our general procedure beyond the small $|{\mathcal{Y}}|$ setting with concrete examples where special structures enable efficient computation.
+We focus on three classes of selection rules that can be of practical interest: selection using only the covariates, selection based on conformal p-values, and selection based on conformal prediction sets.
+When practitioners are willing to slightly modify the selection rule to improve computation efficiency, they may refer to Supplementary Section E.4 where we discuss extensions to simplify the construction of prediction sets by further splitting the calibration data.
 4.1 Covariate-dependent selection rules
-When the selection rule depends solely on the covariates, the reference sets can be efficiently computed by definition, since ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ no longer depends on $y$ (we shall also suppress the dependence of ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ and $S(x,y)$ on $y$ throughout Section 4.1).
-The complete procedure for constructing $\widehat{C}_{\alpha,n+j}$’s and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ with an arbitrary covariate-dependent selection rule is summarized in Algorithm 1, and the computational complexity the procedure is $O(mn\cdot C_{\mathcal{S}})$, with $C_{\mathcal{S}}$ being the complexity of the selection process.
-Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$;
-selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$;
+We first consider covariate-dependent selection rules, i.e., when ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is only a function of $\{X_{i}\}_{i=1}^{n+m}$.
+Under such rules, the reference set no longer depends on $y$; we shall suppress the dependence on $y$ and write ^$\widehat{{\mathcal{R}}}_{n+j}(y)\equiv\widehat{{\mathcal{R}}}_{n+j}$ throughout this subsection.
+Here, ^$\widehat{{\mathcal{R}}}_{n+j}$ can be efficiently computed by looping over $i\in[n]$.
+The complete procedure for constructing $\widehat{C}_{\alpha,n+j}$ and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ with an arbitrary covariate-dependent selection rule is summarized in Algorithm 1.
+Its overall computational complexity is $O(mn\cdot C_{\mathcal{S}})$, with $C_{\mathcal{S}}$ being the complexity of the selection process.
+Algorithm 1 JOMI for arbitrary covariate-dependent selection rules
 […]
-for $j=1,\ldots,m$ do
-       /* Constructing reference sets */
-       Initialize ^$\widehat{{\mathcal{R}}}_{n+j}=\emptyset$;
-       for $i=1,\ldots,n$ do
-             ^$\widehat{{\mathcal{R}}}_{n+j}\leftarrow\widehat{{\mathcal{R}}}_{n+j}\cup\{i\}$   if $j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}$ and $\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}$;
-       end for
-       /* Constructing prediction sets */
-       if form = dtm then
-             $\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}}\cup\{+\infty\}\big{)}\Big{\}}.$
-       end if
-      if form = rand then
-             Sample $U_{j}\sim\textnormal{Unif}[0,1]$;
-             $\widehat{C}_{\alpha,n+j}=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}|}\leq 1-\alpha\bigg{\}}.$
-       end if
-end for
-Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
-Algorithm 1 JOMI for arbitrary covariate-dependent selection rules
-By Theorem 1, the output of Algorithm 1 is valid as long as the selection rule is invariant to the permutation of the calibration points. This includes many commonly-used selection rules:
+[…]
+[…]
+[…]
+[…]
+Sample $U_{j}\sim\textnormal{Unif}[0,1]$.
+[…]
+[…]
+By Theorem 1, the output of Algorithm 1 is valid as long as the selection rule does not rely on the ordering of the calibration points. This includes many commonly-used selection rules:
   1. (1)
-    Top-K selection: the $K$ test units with the highest scores $S(X_{i})$ are selected.
+    Top-K selection.
+The $K$ test units with the highest scores $S(X_{i})$ are selected, where ${\mathcal{S}}\colon{\mathcal{X}}\to\mathbb{R}$ is a pre-trained score function.
+For example, $S(X_{i})$ may be the predicted binding affinity for a drug with chemical structure $X_{i}$ or the predicted health risk of a patient with features $X_{i}$, and the drug discovery process or clinical system admits a fixed number of new units.
   2. (2)
-    Selection based on joint quantiles: unit $i$ is selected if its score $S(X_{i})$ is surpasses the $q$-th quantile of both the calibration and test scores.
+    Selection based on joint quantiles.
+A unit $j$ is selected if its score $S(X_{n+j})$ surpasses the $q$-th quantile of both the calibration and test scores $\{S(X_{i})\}_{i=1}^{n+m}$.
+For instance, a scientist may be interested in toxicities of drug candidates in ${\mathcal{D}}_{\textnormal{test}}$ where those in ${\mathcal{D}}_{\textnormal{calib}}$ have been tested, but they only focus on drugs with highest predicted activities $S(X_{i})$ in the entire library.
   3. (3)
-    Selection based on calibration quantiles: a test unit $i$ is selected if its score $S(X_{i})$ surpasses the $q$-th quantile of calibration scores.
+    Selection based on calibration quantiles.
+A test unit $j$ is selected if its score $S(X_{n+j})$ surpasses the $q$-th quantile of calibration scores $\{S(X_{i})\}_{i=1}^{n}$.
+This may happen when a doctor uses the predicted health risks of existing patients to determine a normal range, and picks test units anticipated to have a relatively extreme health risk.
   4. (4)
-    Selection with constraints: each unit $i$ has a score $S(X_{i})$ and a cost $C(X_{i})$.
-The selection rule amounts to finding a subset $\widehat{{\mathcal{S}}}$ of test units that maximizes $\sum_{j\in\widehat{{\mathcal{S}}}}S(X_{n+j})$ subject to $\sum_{j\in\widehat{{\mathcal{S}}}}C(X_{n+j})\leq C$, where $C$ reflects the total budget.
-As we shall show shortly, it turns out that for rules (1)-(3) we can further improve the computational efficiency by explicitly writing down the reference sets.
-We also note that rules (1)-(2) were considered in Bao et al., (2024), in which the authors proposed methods achieving the weak selection-conditional coverage in (2) and FCR.
-The prediction sets proposed therein coincide with ${\mathcal{C}}_{\alpha,n+j}$, and we shall show that they in fact achieve the strong selection-conditional coverage in (3) for free.
-We provide specially tailored solutions for selection rules (1)-(3) below.
-Throughout, we let ${\mathcal{T}}=\{S_{i}\}_{i\in{\mathcal{I}}_{\textnormal{calib}}\cup{\mathcal{I}}_{\textnormal{test}}}$.
-4.1.1 Top-K selection
-In this example, we assume that $S_{n+j}$’s are distinct for all $j\in[m]$ almost surely (otherwise it is unclear how to define the “top $K$” elements).
-Under the top-K selection rule, the $K$ test units with the highest scores are selected.
-We can write the top-K selection set as
-$\widehat{{\mathcal{S}}}_{\textnormal{topk}}=\{j\in[m]:S_{n+j}>T_{\textnormal{topk}}\},\quad\text{where}\quad T_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{j\in[m]}}\mathds{1}\{S_{n+j}\leq t\}\geq m-K\Big{\}}.$
-For this selection rule, the reference set of each $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ takes a unified form:
-^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{topk}}\}.$ (13)
-Intuitively, $T_{\textnormal{topk}}$ refers to the $(m-K)$-th smallest score among the test units; when $S_{n+j}$ is among the top $K$, replacing $S_{n+j}$ with another score that is greater than $T_{\textnormal{topk}}$ does not change the selection outcome.
-This intuition is formalized in the following proposition, with the proof deferred to Appendix B.5.
+    Selection by black-box optimization procedures.
+A test unit $j$ is selected by running an arbitrary (even black-box) optimization program that does not involve $\{Y_{i}\}_{i\in{\mathcal{I}}_{\textnormal{calib}}}$.
+One such example is optimization under constraints: apart from the score $S(X_{n+j})$, each unit $j$ is associated with a cost $C_{n+j}$, and $\widehat{{\mathcal{S}}}$ is the subset of test units that maximizes $\sum_{j\in\widehat{{\mathcal{S}}}}S(X_{n+j})$ subject to $\sum_{j\in\widehat{{\mathcal{S}}}}C_{n+j}\leq c$, where $c$ reflects the total budget.
+This may happen in healthcare management systems that optimize resources subject to constraints and send patients to different care categories, or in candidate screening for job interviews subject to budget and diversity constraints.
+More generally, in drug discovery, scientists may run a complex Bayesian optimization algorithm to select the next batch of drugs to evaluate (Pyzer-Knapp,, 2018), which may be viewed as a black box process.
+No matter how complicated these optimization programs are, as long as they do not involve calibration labels, JOMI supports efficient uncertainty quantification afterwards.
+We will see some stylized examples in our numerical experiments.
+We additionally show in Supplementary Section B that we can further improve the computational efficiency of JOMI for rules (1)-(3) by deriving exact forms of the reference sets, where in each case the computation complexity is $O(\max\{m,n\})$.
+We also note that rules (1)-(2) were considered in Bao et al., (2024), where they propose methods that achieve (2) and FCR control.
+The prediction sets proposed therein coincide with ours, and our results imply that they in fact achieve the strong selection-conditional coverage in (3) for free.
+4.2 Selection based on conformal p-values
+The second class of selection rules we study concern selecting units whose outcomes satisfy certain conditions while controlling some type-I error.
+To this end, the test units are selected by thresholding a class of conformal p-values, where each p-value is computed via contrasting a test point with the calibration data.
+As such, the selection rule can be complicated and asymmetric.
+We will follow the framework of Jin and Candès, 2023b to define the p-values and selection rules, who study the problem of discovering test units with large outcomes.
+Examples include selecting drugs with sufficiently high binding affinities, finding highly competent job candidates, and identifying patients who benefit from a treatment, etc.
+In these problems, predictions from machine learning models serve as proxies for the true outcomes of interest that are too expensive or impossible to evaluate, and the selection procedure leverages the power of predictions to select units with large outcomes while ensuring error control.
+Given test points $\{X_{n+j}\}_{j\in[m]}$ and (potentially random) thresholds $c_{n+j}\in\mathbb{R}$, the goal is to select those $Y_{n+j}>c_{n+j}$ while controlling the number/fraction of false positives.
+The statistical evidence for detecting a large outcome is quantified by the so-called “conformal p-values”.
+Suppose the calibration data are $\{(X_{i},Y_{i},c_{i})\}_{i=1}^{n}$, such that the tuples $\{X_{i},Y_{i},c_{i}\}_{i=1}^{n+m}$ are exchangeable.
+Assume access to a score function $S\colon{\mathcal{X}}\times\mathbb{R}\to\mathbb{R}$ such that $S(x,y)$ is non-increasing in $y$ for any $x\in{\mathcal{X}}$.
+An example is $S(x,y)=\widehat{\mu}(x)-y$ where $\widehat{\mu}(x)$ is a point predictor trained on ${\mathcal{D}}_{\textnormal{train}}$.
+We then compute $\widehat{S}_{i}=S(X_{i},c_{i})$ for $i\in[n+m]$, and define the conformal p-values111 Our p-value slightly modify the definition in Jin and Candès, 2023b for the ease of describing our prediction sets.
+Similar to the original ones, our p-values control the type-I error in detecting one large outcome.
+In addition, using our p-values in their original procedures maintains error control with improved power; we discuss these results in Supplementary Section E.2 for completeness.
+$p_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}}{n+1},\quad j\in[m].$ (15)
+Hereafter, we call $S$ the selection score function.
+We can show that $p_{j}$ is valid in the sense that $\mathbb{P}(p_{j}\leq t,\leavevmode\nobreak\ Y_{n+j}\leq c_{n+j})\leq t,$ $\forall t\in[0,1]$.
+That is, testing with $p_{j}$ controls the type-I error in finding one large outcome, accounting for the randomness in the outcomes as well.
+We then select test units whose conformal p-values are below a threshold, the choice of which determines the type of error control guarantee.
+Some examples are given below.
+  1. 1.
+    Fixed threshold.
+We select test units whose conformal p-values (15) are below a fixed threshold $q\in(0,1)$.
+This could happen when testing a single hypothesis, or testing multiple hypotheses with Bonferroni correction.
+The latter controls the family-wise error rate in finding large outcomes, which is useful in highly risk-sensitive settings such as disease diagnosis.
+  2. 2.
+    BH threshold.
+We select test units whose conformal p-values (15) are below a data-dependent threshold given by the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995).
+This selection procedure is shown in Jin and Candès, 2023b to control the FDR in detecting large outcomes, which is useful in exploratory screening such as drug discovery for ensuring efficient resource use in follow-up investigations.
+For generality, we consider the selection rule in the form of $\widehat{{\mathcal{S}}}_{\textnormal{cp}}=\{j\in[m]:\widehat{S}_{n+j}\geq\tau\}$, where $\tau$ is a stopping time adapted to the filtration $\{\sigma(\{A_{s}\}_{s\leq t},\{B_{s}\}_{s\leq t})\}_{t\in\mathbb{R}}$:
+$A_{s}=1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq s,Y_{i}\leq c_{i}\},\leavevmode\nobreak\ B_{s}=\sum_{j\in[m]}\mathds{1}\{\widehat{S}_{n+j}\geq s\}.$ (16)
+Equivalently, for any $t\in\mathbb{R}$, we can write $\mathds{1}\{\tau\leq t\}=f_{t}(\{A_{s}\}_{s\leq t},\{B_{s}\}_{s\leq t})$ for some function $f_{t}$.
+It can be shown that the two examples above are special cases of this general rule (these results can be found in Section E.3 in the supplementary material).
+We now provide a general solution for constructing selection-conditional prediction sets corresponding to such rules.
+The challenge here is that all the p-values depend on each other, as they leverage the same set of calibration data, and/or the data-driven threshold determined by all p-values would further add to the intricacy.
+We note that the BH-based rule is studied in Bao et al., (2024) with approximate FCR control, while we are to provide an efficient solution with exact coverage guarantee.
+The following proposition lays out the form of the reference set and its validity, with its proof delegated to Section D.5 in the supplementary material.
 Proposition 4.
-For the top-K selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}\in 2^{[m]}$, such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{topk}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ for all $y\in{\mathcal{Y}}$.
-Computing ^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ as a universal reference set reduces the computation complexity of the reference construction step in Algorithm 1 to $O(\max(m,n))$.
-It is also worth mentioning that the definition of reference set here is agnostic to the choice of $\mathfrak{S}$.
-4.1.2 Selection based on joint quantiles
-In some cases, practitioners may want to select a test unit $n+j$ if its score $S_{n+j}$ greater than the $q$-th quantile of the calibration and test scores.
-Equivalently, we could write the selection set as
-$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\{j\in[m]:S_{n+j}>T_{\textnormal{jq}}\},\quad\text{where}\quad T_{\textnormal{jq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{m+n}}\mathds{1}\{S_{i}\leq t\}\geq q(n+m)\Big{\}}$
-By definition, the selection threshold $T_{\textnormal{jq}}$ is invariant to the permutation of the calibration and test scores, and it is straightforward to check that for all $j\in\widehat{{\mathcal{S}}}$, the reference set is in the form of
-^$\widehat{{\mathcal{R}}}_{\textnormal{jq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{jq}}\}.$
-Using ^$\widehat{{\mathcal{R}}}_{jq}$ reduces the computation complexity of the reference set in Algorithm 1 to $O(\max(m,n))$.
-4.1.3 Selection based on calibration quantiles
-Consider now a selection rule, where test unit $n+j$ is selected if its score $S_{n+j}$ is greater than the $q$-th quantile of of the calibration scores.
-Equivalently, we could write the selection set as
-$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\big{\{}j\in[m]:S_{n+j}>T_{\textnormal{cq}}\big{\}},\text{ where }T_{\textnormal{cq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{n}}\mathds{1}\{S_{i}\leq t\}\geq qn\Big{\}}.$ (14)
-For each $j\in\widehat{{\mathcal{S}}}$, the refernence set consists of all calibration units in the top $q$-th quantile, i.e.,
-^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{cq}}\}.$ (15)
-The validity of ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ as the referenece set is established below, whose proof is deferred to appendix B.6.
+Suppose the selection set is $\widehat{{\mathcal{S}}}_{\textnormal{cp}}=\{j\in[m]:\widehat{S}_{n+j}\geq\tau\}$, where $\tau$ is determined by $\mathds{1}\{\tau\leq t\}=f_{t}(\{A_{s}\}_{s\leq t},\{B_{s}\}_{s\leq t})$, for some $f_{t}$ and $(A_{s},B_{s})$ defined in (16), $\forall t\in\mathbb{R}$.
+For any $\mathfrak{S}\in 2^{[m]}$ and $j\in[m]$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{cp}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{cp}}\in\mathfrak{S}$, the reference set can be simplified as
+^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)=\mathds{1}\{y\leq c_{n+j}\}\cdot\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp},1}+\mathds{1}\{y>c_{n+j}\}\cdot\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp},0},\quad\text{where}$ (17)
+^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp},k}=$$\{i\in[n]:Y_{i}\leq c_{i},\widehat{S}_{i}\geq\tau^{(j)}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(k,0)\}\in\mathfrak{S}\}$ (18)
+$\cup\{i\in[n]:Y_{i}>c_{i},\widehat{S}_{i}\geq\tau^{(j)}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(k,1)\}\in\mathfrak{S}\}.$ (19)
+For $k,\ell\in\{0,1\}$, the adjusted threshold $\tau^{(j)}(k,\ell)$ is given by
+$\mathds{1}\{\tau^{(j)}(k,\ell)\leq t\}=f_{t}(\{A_{s}^{(j)}(k,\ell)\}_{s\leq t},\{B_{s}^{(j)}\}_{s\leq t}),\text{where}$ (20)
+$A_{s}^{(j)}(k,\ell)=\ell+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq s,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq s\},\leavevmode\nobreak\ B^{(j)}_{s}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}.$
+Based on Proposition 4, the JOMI prediction set is given by
+$\widehat{C}_{\alpha,n+j}=\big{\{}y\in{\mathcal{Y}}:y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\big{\}}\cup\big{\{}y\in{\mathcal{Y}}:y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\big{\}},$
+where $\widehat{q}_{k}={\textnormal{Quantile}}(1-\alpha;\{V_{i}:i\in\widehat{{\mathcal{R}}}^{\textnormal{cp},k}_{n+j}\}\cup\{\infty\})$ for $k=0,1$.
+See Algorithm 2 for a summary of the complete procedure, where we only present the deterministic version for simplicity.
+The overall computation complexity is at most $O(m(m+n)|\widehat{\mathcal{S}}|)$.
+Algorithm 2 JOMI for selection based on conformal p-values
+Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$; selection rule ${\mathcal{S}}$; nonconformity score function $V(\cdot,\cdot)$.
+[…]
+Compute $\tau^{(j)}(k,\ell)$ as in (20), for $(k,\ell)\in\{0,1\}$.
+Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{cp},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{cp},1}$ as (18).
+Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}:i\in\widehat{\mathcal{R}}_{n+j}^{\text{cp},k}\}\cup\{\infty\})$, for $k=0,1$.
+[…]
+[…]
+4.3 Selection based on conformal prediction sets
+The final class of selection rules we study are based on the properties of (preliminary) prediction sets, usually constructed by running the vanilla SCP.
+Such use cases have appeared implicitly in many heuristic applications of conformal prediction.
+For example, practitioners may select units whose prediction intervals are shorter/longer than a threshold, which roughly indicates enough confidence (Sokol et al.,, 2024).
+People may also select units whose prediction sets entirely lie above a threshold, which roughly indicates a desired outcome (Svensson et al.,, 2017).
+Note that the original prediction intervals are no longer valid conditional on being selected (Jin and Candès, 2023b, ), and thus using them for interpreting downstream uncertainty can be misleading.
+In this section, we apply our general framework to re-calibrate prediction sets for the units selected in such a way.
+Formally, we consider two stages of prediction set construction.
+The one constructed in the first stage, called the preliminary prediction set, is used for determining the selection set $\widehat{\mathcal{S}}$.
+The one in the second stage, which we call the selective prediction set, is the one we are to build with JOMI.
+Following SCP in Section 3.1, we let $S(x,y)$ and $V(x,y)$ be the nonconformity score functions for the two stages, respectively.
+The ($1-\beta$)-level preliminary prediction set for the $j$-th test unit is
+$\widehat{C}^{\text{prelim}}_{\beta,n+j}=\big{\{}y\in{\mathcal{Y}}:S(X_{n+j},y)\leq\eta\big{\}},$ (21)
+where $\eta$ is the $K\,:=\,\lceil(1-\beta)(n+1)\rceil$-th smallest element in $\{S(X_{i},Y_{i})\}_{i=1}^{n}$.
+We consider any selection rule based on the preliminary prediction set $\widehat{C}_{\beta,n+j}^{\text{prelim}}$.
+Note that by (21), given the first-stage score function $S(\cdot,\cdot)$, the form of $\widehat{C}_{\beta,n+j}^{\text{prelim}}$ is fully determined by $X_{n+j}$ and $\eta$.
+We can thus express any selection rule through ${\mathcal{L}}:{\mathcal{X}}\times\mathbb{R}\mapsto\{0,1\}$, where ${\mathcal{L}}(X_{n+j},\eta)=1$ means selecting the unit and ${\mathcal{L}}(X_{n+j},\eta)=0$ otherwise.
+An example is selecting based on prediction interval lengths: following Sokol et al., (2024), suppose that we use CQR (Romano et al.,, 2019) in the first stage, i.e., $S(x,y)=\max\{\widehat{q}_{L}(x)-y,y-\widehat{q}_{U}(x)\}$, where $\widehat{q}_{L}(x)$ and $\widehat{q}_{U}(x)$ are estimates of some lower and upper conditional quantiles.
+Selecting prediction intervals shorter than a threshold $\lambda$ gives ${\mathcal{L}}(x,\eta)=\mathds{1}\{\widehat{q}_{U}(x)-\widehat{q}_{L}(x)+2\eta\leq\lambda\}$.
+As another example, for a binary outcome $Y$, we might want to select units whose prediction set is a singleton, leading to ${\mathcal{L}}(x,\eta)=\mathds{1}\big{\{}S(x,1)\leq\eta<S(x,0)\text{ or }S(x,0)\leq\eta<S(x,1)\big{\}}$.
+Having determined the selection rule ${\mathcal{L}}$, the selection set is thus ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})=\widehat{{\mathcal{S}}}_{\textnormal{ps}}=\big{\{}j\in[m]:{\mathcal{L}}(X_{n+j},\eta)=1\big{\}}.$ We are to derive a computationally efficient but slightly conservative version of the JOMI prediction set, which nevertheless has tight coverage in all our numerical experiments (see Section 6).
+Define
+$\widehat{C}^{\textnormal{ps}}_{\alpha,n+j}:=\leavevmode\nobreak$$\{y:\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}\}\cup\{y:V(X_{n+j},y)\leq q_{j,1}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ S(X_{n+j},y)<\eta^{-}\}$ (22)
+$\cup\{y:V(X_{n+j},y)\leq q_{j,2}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ S(X_{n+j},y)>\eta^{+}\}$
+where $\eta^{+}$ and $\eta^{-}$ are the $(K+1)$-th and $(K-1)$-th smallest element in $\{S_{i}\}_{i=1}^{n}$, respectively, and
+$q_{j,1}:={\textnormal{Quantile}}\Big{(}1-\alpha;\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta)=1\}\in\mathfrak{S}\big{\}}$
+$\qquad\qquad\qquad\qquad\qquad\cup\big{\{}V_{i}:S_{i}>\eta^{-},\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta^{-})=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{-})=1\}\in\mathfrak{S}\big{\}}\Big{)};$
+$q_{j,2}:={\textnormal{Quantile}}\Big{(}1-\alpha;\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta,{\mathcal{L}}(X_{i},\eta^{+})=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{+})=1\}\in\mathfrak{S}\big{\}}$
+$\qquad\qquad\qquad\qquad\qquad\cup\big{\{}V_{i}:S_{i}>\eta,\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta)=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta)=1\}\in\mathfrak{S}\big{\}}\Big{)}.$ (23)
+We prove the validity of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ in (22) below, whose proof is in Supplementary Section D.6.
 Proposition 5.
-For the selection rule (14), for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$, such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\widehat{{\mathcal{R}}}_{n+j}(y)$ defined in (7) for all $y\in{\mathcal{Y}}$.
-Similar to the previous two cases, directly computing ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ reduces the computation complexity of the reference set to $O(\max(m,n))$.
-As a side note, Bao et al., (2024) also considered the selection rule based on calibration quantiles (as an instance of “calibration-assisted selection”), for which they provide asymptoically FCR control guarantees under additional assumptions.
-4.2 Selection based on conformal p-values
-Suppose we wish to select the test units whose outcomes satisfy certain conditions.
-For example, Jin and Candès, 2023b studies the problem of selecting test units with large outcomes, where they formulate the task as hypothesis testing problems.
-Following their formulation, we consider for each $j\in[m]$ the following null and alternative hypotheses:
-$H_{j,0}:Y_{n+j}\leq c_{n+j}\text{ v.s. }H_{j,1}:Y_{n+j}>c_{n+j},$ (16)
-where $c_{n+j}$ is a (potentially random) threshold.
-When $H_{j,0}$ is rejected, we select the $j$-th test unit.
-For testing $H_{j,0}$in and Candès, 2023b devise the conformal p-value $p_{j}$ such that
-$\mathbb{P}(p_{j}\leq t,Y_{n+j}\leq c_{j})\leq t,\text{ for any }t\in[0,1].$
-In this paper, we follow their strategy, but work with a slightly different form of the conformal p-value that is uniformly more powerful than the original version in Jin and Candès, 2023b .
-Throughout Section 4.2, we assume the score function $S(x,y)$ to be monotonically nonincreasing in $y$, e.g., $s(x,y)=\widehat{\mu}(x)-y$.
-For each $i\in{\mathcal{I}}_{\textnormal{calib}}\cup{\mathcal{I}}_{\textnormal{test}}$, define $\widehat{S}_{i}=S(X_{i},c_{i})$, and for any $j\in[m]$, we define the conformal p-value as
-$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}}{n+1}.$ (17)
-We provide a proof for the validity of $p_{j}$ for testing $H_{j,0}$, as well as its comparison with Jin and Candès, 2023b in Appendix C.1 for completeness.
-With these p-values, we can select the units whose conformal p-values are smaller than a threshold. In terms of the threshold choice, we consider two scenarios:
-  1. 1.
-    Fixed threshold: select test units whose conformal p-values (17) are below a fixed threshold $q\in(0,1)$, such as Bonferroni correction.
-  2. 2.
-    Data-driven threshold: select test units whose conformal p-values (17) are below a data-dependent threshold given by the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995).
-4.2.1 Fixed threshold
-We first consider $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:p_{j}\leq q\}$ with a threshold $q\in(0,1)$.
-We will see that there are only two potential forms of the reference set ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ when varying $y\in\mathbb{R}$.
-For $k\in\{0,1\}$, define
-^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,0)\}\in\mathfrak{S}\Big{\}}$
-$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (18)
-where we recall that $\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}$ is the score in the ($n+\ell$)-th position after swapping $(i,j)$. For $k,\ell\in\{0,1\}$,
-$T^{\textnormal{fixed}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}\leq q(n+1)\bigg{\}}.$ (19)
-Proposition 6 presents the simplified expression for ^$\widehat{\mathcal{R}}_{n+j}(y)$ for any selected unit and any selection taxonomy, whose proof is deferred to Appendix B.7.
-Proposition 6.
-Suppose the selection set contains the units whose conformal p-values (17) are below a fixed value $q\in(0,1)$, i.e., $\widehat{{\mathcal{S}}}=\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]\colon p_{j}\leq q\}$.
-Then for any selection taxonomy $\mathfrak{S}$ and any $j\in[m]$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},0}_{n+j}$ when $y>c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},1}_{n+j}$ when $y\leq c_{n+j}$, for any $y\in\mathbb{R}$.
-Finally, the JOMI prediction interval is therefore
-$\widehat{{\mathcal{C}}}_{\alpha}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\big{\}}\cup\big{\{}y\in{\mathcal{Y}}:y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\big{\}},$
-where $\widehat{q}_{k}={\textnormal{Quantile}}(\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed},k}_{n+j}}\cup\{\infty\};1-\alpha)$ for $k=0,1$.
-See Algorithm LABEL:alg:pval_fix for a summary.
-The computation complexity of Algorithm LABEL:alg:pval_fix is polynomial in $n$ and $m$.
-Note that for each $j$, computing $T_{j}^{\textnormal{fixed}}(k,\ell)$ costs $O(m)$, computing ^$\widehat{\mathcal{R}}_{n+j}^{\textnormal{fixed},k}$ costs $O(mn)$, and computing $\widehat{C}_{\alpha,n+j}$ costs $O(n)$.
-Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
-selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
-for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
-       Compute $T_{j}^{\textnormal{fixed}}(k,\ell)$ as in (19) for $(k,\ell)\in\{0,1\}$.
-Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},1}$ as (4.2.1).
-Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\text{fixed},k}}\cup\{\infty\})$, $k=0,1$.
-Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
-end for
-Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
-Algorithm 2 JOMI for conformal p-values with fixed threshold
-4.2.2 Data-driven p-value threshold
-According to Jin and Candès, 2023b , applying the BH procedure (Benjamini and Hochberg,, 1995) to the p-values $\{p_{j}\}_{j=1}^{m}$ yields a selection set with controlled false discovery rate (FDR).
-Formally, let $p_{(1)}\leq p_{(2)}\leq\ldots\leq p_{(m)}$ be the order statistics of the conformal p-values.
-The BH selection set $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ is
-$\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\colon p_{j}\leq k^{*}q/m\},\quad k^{*}=\max\Big{\{}k:p_{(k)}\leq kq/m\Big{\}},$
-with the convention that $\max\varnothing=0$.
-For the BH selection rule, the reference set can also be simplified to two cases. For $k=0,1$, define
-^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,0)\}\in\mathfrak{S}\Big{\}}$
-$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (20)
-and for $k,\ell=0,1$, we define
-$T^{\textnormal{BH}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}}{1+\sum_{j\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{(k,\ell)}\}}\cdot\frac{m}{n+1}\leq q\bigg{\}}.$ (21)
-We prove the equivalence in the next proposition, deferring the proof to Appendix B.8.
-Proposition 7.
-For the BH selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},1}$ when $y\leq c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},0}$ when $y>c_{n+j}$ as defined in (4.2.2) for any $y\in\mathbb{R}$.
-Therefore, the JOMI prediction set is of a similar form as the fixed-threshold case.
-We summarize the entire workflow in Algorithm 3.
-Similarly, the computation cost of $\widehat{C}_{\alpha,n+j}$ is $O(mn)$.
-Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
-selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
-for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
-       Compute $T_{j}^{\textnormal{BH}}(k,\ell)$ as in (21) for $(k,\ell)\in\{0,1\}$.
-Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},1}$ as (4.2.2).
-Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\textnormal{BH},k}}\cup\{\infty\})$, $k=0,1$.
-Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
-end for
-Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
-Algorithm 3 JOMI for conformal p-values with BH-threshold
-4.3 Selection based on conformal prediction sets
-Practitioners may also select test unit(s) based on the property of a preliminary prediction set.
-For example, they may select units whose prediction intervals are shorter than a threshold (Sokol et al.,, 2024), or whose prediction sets that entirely lie above a threshold.
-Note that the original prediction intervals are no longer valid conditional on being selected (Jin and Candès, 2023b, ).
-In this section, we instantiate our general framework to re-calibrate prediction sets for the units selected in such a way.
-We thus consider two stages of prediction set construction: the one constructed in the first stage, which we call the preliminary prediction set, is used for determining the selection set $\widehat{\mathcal{S}}$, while that in the second stage, which we call the selective prediction set, is the one we are to build with JOMI.
-Recall the SCP framework in Section 3.1.
-Let $S(x,y)$ and $V(x,y)$ be the nonconformity score for the two stages, respectively.
-The level $1-\beta$ preliminary prediction set for the $j$-th test unit is
-$\widehat{C}^{\text{prelim}}_{\beta}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:S(X_{n+j},y)\leq\eta\big{\}},$ (22)
-where $\eta$ is the $K\,:=\,\lceil(1-\beta)(n+1)\rceil$-th smallest element in $\{S_{i}\}_{i=1}^{n}$.
-By (22), the preliminary prediction set is fully determined by $X_{n+j}$ and $\eta$.
-We can thus consider a general selection rule ${\mathcal{L}}:{\mathcal{X}}\times\mathbb{R}\mapsto\{0,1\}$, where ${\mathcal{L}}(X_{n+j},\eta)=1$ means selecting the unit.
-For instance, selecting prediction intervals shorter than a threshold $\lambda$ gives ${\mathcal{L}}(x,\eta)=\mathds{1}\{\widehat{q}_{U}(x)-\widehat{q}_{L}(x)+2\eta\leq\lambda\}$ if we use CQR for $\widehat{C}_{\beta}^{\text{prelim}}$ (Romano et al.,, 2019), i.e., setting $S(x,y)=\max\{\widehat{q}_{L}(x)-y,y-\widehat{q}_{U}(x)\}$, where $\widehat{q}_{L}(x)$ and $\widehat{q}_{U}(x)$ are estimates of some lower and upper conditional quantiles.
-As another example, for binary $Y$, we might want to select the units whose prediction set is a singleton, then the selection rule can be written as ${\mathcal{L}}(x,\eta)=\mathds{1}\big{\{}S(x,1)\leq\eta<S(x,0)\text{ or }S(x,0)\leq\eta<S(x,1)\big{\}}$.
-The selection set for a selection rule ${\mathcal{L}}$ is
-$\widehat{{\mathcal{S}}}_{\textnormal{ps}}=\big{\{}j\in[m]:{\mathcal{L}}(X_{n+j},\eta)=1\big{\}}.$
-We are to derive a general form for a slightly conservative version of the JOMI prediction set, which nevertheless has tight coverage in all our numerical experiments (see Section 6 for examples).
-We only present the results for $\mathfrak{S}=2^{[m]}$ for the ease of exposition.
-Formally, we define
-$\widehat{C}^{\textnormal{ps}}_{\alpha,n+j}:=\leavevmode\nobreak$$\{y:\eta^{-}\leq S(X_{n+1},y)\leq\eta^{+}\}$ (23)
-$\cup\{y:V(X_{n+j},y)\leq q_{1},S(X_{n+1},y)<\eta^{-})\}$
-$\cup\{y:V(X_{n+j},y)\leq q_{2},S(X_{n+1},y)>\eta^{+})\}$
-where $\eta^{+}$ and $\eta^{-}$ are the $(K+1)$-th and $(K-1)$-th smallest element in $\{S_{i}\}_{i=1}^{n}$, respectively, and
-$q_{1}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta^{-},\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta^{-})=1\big{\}}\Big{)};$
-$q_{2}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta,{\mathcal{L}}(X_{i}.\eta^{+})=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta,\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta)=1\big{\}}\Big{)}.$ (24)
-We prove the validity of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ in (23) below, whose proof is deferred to Appendix B.9.
-Proposition 8.
-For any selection rule ${\mathcal{L}}$, any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{ps}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{ps}}\in\mathfrak{S}$, the prediction set $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is a superset of the JOMI prediction set $\widehat{C}_{\alpha,n+j}$ defined in (8).
-The procedure is summarized in Algorithm 4. For each $j$, the computation cost of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is $O(m)$.
-Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; selection taxonomy $\mathfrak{S}$;
-selective miscoverage level $\alpha$; preliminary micoverage level $\beta$;
-selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
-Compute $q_{1}$ and $q_{2}$ as in (4.3).
+For any selection rule ${\mathcal{L}}$, any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{ps}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{ps}}\in\mathfrak{S}$, $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is a superset of the JOMI prediction set $\widehat{C}_{\alpha,n+j}$ defined in (9), and
+$\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}\backslash\widehat{C}_{\alpha,n+j}\subseteq\big{\{}y\in{\mathcal{Y}}\colon\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}\big{\}}.$
+By Proposition 5, the conservativeness of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is quite limited, as $\eta^{-}$ and $\eta^{+}$ are usually very close to each other.
+We also verify its tight empirical coverage in Section 6.
+The procedure is summarized in Algorithm 3. For each $j$, the computation cost of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is $O(m+n)$, and therefore the overall computation cost is $O(m(m+n))$.
+Algorithm 3 JOMI for selection based on preliminary prediction sets
+Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; selection taxonomy $\mathfrak{S}$; selective miscoverage level $\alpha$; first-stage miscoverage level $\beta$; selection rule ${\mathcal{S}}$; first-stage score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
+$\widehat{{\mathcal{S}}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$.
 Compute $\eta^{+}$ as the $(K+1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$, where $K=\lceil(1-\beta)(n+1)\rceil$.
-Compute $\eta^{+}$ as the $(K-1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$ for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
-       Compute $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ as in (23).
-end for
-Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
-Algorithm 4 JOMI for selection based on preliminary prediction sets
+[…]
+Compute $q_{j,1}$ and $q_{j,2}$ as in (4.3).
+[…]
+[…]
+Remark 2.
+We note that it is possible that the prediction interval produced by Algorithm 3 does not satisfy the constraints on the preliminary prediction sets.
+This is, however, a desired feature in our setting where the selection rule is given and the inference step is decoupled from the selection step; the selection-conditional prediction sets should be used as tools for informing further decisions/analysis.
+Otherwise, we may need to take an orthogonal strategy: change the selection algorithm, for which ideas from Jin and Candès, 2023b ; Gazin et al., (2024) may be useful.
 5 Application to drug discovery
 In drug discovery, powerful prediction machines are increasingly used to guide the search of promising drug candidates.
-For such high-stakes decision-making, it is important to quantify the uncertainty in complex prediction machines (Svensson et al.,, 2017; Jin and Candès, 2023b, ; Laghuvarapu et al.,, 2024).
-Meanwhile, scientists will naturally focus on a subset of seemingly promising drugs.
+For such high-stakes decisions, it is important to quantify the uncertainty in the predictions (Svensson et al.,, 2017; Jin and Candès, 2023b, ; Laghuvarapu et al.,, 2024).
+Meanwhile, selection issues naturally arise as scientists may only focus on seemingly promising drugs.
 In this section, we apply JOMI to several application scenarios in drug discovery with a selective nature.
@@
 Application scenarios.
-We study two problems in drug discovery:
-  1. 1.
-    Drug property prediction (DPP), a classification problem where the binary outcome indicates whether a drug candidate binds to a pre-specified disease target, and the covariates are the (encoded) chemical structure of the drug compound.
-  2. 2.
-    Drug-target-interaction prediction (DTI), a regression problem where each sample is a pair of drug and disease target.
+In the main text, we focus on drug property prediction (DPP), a classification problem where the binary outcome indicates whether a drug candidate binds to a pre-specified disease target, and the covariates are the (encoded) chemical structure of the drug compound.
+Due to limited space, we defer the results for several selection scenarios in drug-target-interaction prediction (DTI) to Supplementary Section C.4.
+DTI is a regression problem where each sample is a pair of drug and disease target.
 The outcome of interest is a real-valued variable indicating the binding affinity of that pair.
-The covariates are the (concatenated) encoded structure of the drug and the disease.
-The datasets and prediction models are based on the DeepPurpose library (Huang et al.,, 2020), with details in subsequent subsections. In each problem, we consider several realistic selection rules.
-Selection rules.
-We consider three types of realistic selection rules ${\mathcal{S}}$:
+The covariates are the (concatenated) encoded structure of both.
+Selection rules. We consider three types of realistic selection rules ${\mathcal{S}}$:
   1. (1)
-    Covariate-dependent top-K selection: selecting drugs with highest predicted binding affinities $\widehat{\mu}(X_{n+j})$.
+    Covariate-dependent top-K selection: selecting drugs with highest predicted binding affinities.
     1. (i)
-      Top-K selection for test data.
-When the scientist has a fixed budget of investigating $K$ drug candidates in the next phase, one may select $K$ test samples with the largest values of $\widehat{\mu}(X_{j})$.
+      Top-K among test data. When the scientist has a fixed budget of investigating $K$ drug candidates in the next phase, one may select $K$ test samples with the largest $\widehat{\mu}(X_{n+j})$.
     2. (ii)
-      Top-K selection with mixed data.
+      Top-K among mixed data.
 When the scientist is to investigate other properties for promisingly active drug candidates in the next phase, they may select $K$ units in ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$ with the highest predicted affinities.
@@
       Calibration-referenced selection.
-The scientist may use the calibration data as reference and select test samples whose predicted activities are greater than the $K$-th highest in $\{\widehat{\mu}(X_{i})\}_{i\in{\mathcal{I}}_{\textnormal{calib}}}$.
+The scientist may use the ${\mathcal{D}}_{\textnormal{calib}}$ as reference and select test samples whose predicted activities are greater than the $K$-th highest in $\{\widehat{\mu}(X_{i})\}_{i\in{\mathcal{I}}_{\textnormal{calib}}}$.
   2. (2)
-    Conformalized selection.
+    Conformal selection.
 The scientist might also obtain a subset of active drugs while controlling the FDR below some $q\in(0,1)$.
-In this case, $\widehat{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the set of test drugs picked by the conformalized selection method in Jin and Candès, 2023b at FDR level $q\in(0,1)$.
+In this case, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the set of test drugs picked by Conformal Selection in Jin and Candès, 2023b at FDR level $q\in(0,1)$.
   3. (3)
     Selection with constraints.
-The cost of developing a drug is a function $C(x)$, and one wants to select as many drugs with highest predictions $\widehat{\mu}(X_{n+j})$ as possible while ensuring the total cost is below some constant $C$, or maximize the total reward while controlling the total budget.
-For brevity, each selection rule is reported for one application, while we defer others to Appendix A.
+The cost of developing a drug is a variable $C_{n+j}$, and one wants to select as many drugs with the highest predictions $\widehat{\mu}(X_{n+j})$ as possible while ensuring the total cost is below some constant $C$, or maximize the total reward while controlling the costs.
 Evaluation metrics.
-We evaluate the selection-conditional (mis)coverage via the consistent estimator
-$\widehat{\textrm{Miscov}}=\frac{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j})}{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}})},$
-where $\widehat{P}$ is the empirical probability over all repeats.
-We also evaluate the size of prediction sets by averaging the cardinality of $\widehat{\mathcal{C}}_{j}$ for classification or length of $\widehat{\mathcal{C}}_{j}$ for regression over selected test units in all repeats.
-We also evaluate (but do not show in figures for brevity) the false coverage rate via $\widehat{\textrm{FCR}}=\widehat{E}\Big{[}\frac{\sum_{j=n+1}^{n+m}\mathds{1}\{j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j}\}}{1\vee|\widehat{\mathcal{S}}|}\Big{]},$ where $\widehat{E}$ is the empirical mean over all repeats.
-5.1 Drug property prediction
-We first consider the task of predicting the binary binding property of drug candidates.
-We use the HIV screening data in DeepPurpose with a total sample size of $n_{\textnormal{tot}}=41127$.
+We evaluate the selection-conditional (mis)coverage via the consistent estimator $\widehat{\textrm{Miscov}}=\frac{\sum_{j=1}^{m}\widehat{P}(j\in\widehat{\mathcal{S}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j})}{\sum_{j=1}^{m}\widehat{P}(j\in\widehat{\mathcal{S}})},$ where $\widehat{P}$ is the empirical probability over all repeats.
+The size of prediction sets is evaluated by averaging the cardinality of $\widehat{C}_{\alpha,n+j}$ for classification or length of $\widehat{C}_{\alpha,n+j}$ for regression over selected test units in all repeats.
+We also evaluate (but do not show in figures for brevity) the false coverage rate via $\widehat{\textrm{FCR}}=\widehat{E}\Big{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}\}}{1\vee|\widehat{\mathcal{S}}|}\Big{]},$ where $\widehat{E}$ is the empirical mean over all repeats.
+Data and prediction models.
+We use the HIV screening data in the DeepPurpose library (Huang et al.,, 2020) with a total sample size of $n_{\textnormal{tot}}=41127$.
 The numerical features $X\in{\mathcal{X}}$ are encoded by Extended-Connectivity FingerPrints (Rogers and Hahn,, 2010, ECFP) which characterize topological properties of the drug candidates.
@@
 Then, the remaining is randomly split into two equally-sized folds ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$.
-The exchangeability among data points is thus satisfied.
+The exchangeability among ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$ is thus satisfied.
 Sampling of the training data and model training is independently repeated $10$ times; within each, we randomly split calibration/test data for $100$ times.
-This leads to $N=1000$ independent runs.
-5.1.1 Top-K selection
+This leads to $N=1000$ runs in total.
+5.1 Top-K selection
 We first consider the selection rule (i) top-K with test data.
-For a trained predictor $\widehat{\mu}$, we experiment with selecting (a) the highest $K$ values of $\widehat{\mu}(X_{j})$ and (b) the lowest $K$ values of $\widehat{\mu}(X_{j})$ among all test data.
-We observe similar patterns in (ii) top-K selection with both test and calibration data, as well as (iii) calibration-referenced selection.
-For brevity, we defer results for these selection rules to Appendix A.1.
-For a confidence level $\alpha\in(0,1)$, we apply the vanilla conformal prediction and JOMI (both deterministic and randomized) to the selected units at level $\alpha$, with $\mathfrak{S}=2^{{\mathcal{I}}_{{\textnormal{test}}}}$.
-Since the selection size is fixed, valid selection-conditional coverage implies FCR control at level $\alpha$.
-We vary $\alpha\in\{0.1,0.2,\dots,0.9\}$ and $K\in\{20,100,1000,2000,5000,10000,15000\}$.
-We set $V$ as the APS score (Romano et al.,, 2020)
-Figure 2 shows the empirical selection-conditional coverage $\widehat{\textnormal{Miscov}}$ on the left panel and the average prediction set size on the right panel, both at marginal miscoverage levels $\alpha\in\{0.1,0.8\}$.
-The solid (dashed) lines show the results when we select units with the highest (lowest) predicted affinities.
-Figure 2: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for vanilla conformal prediction (Vanilla_CP), JOMI and randomized JOMI (JOMI_rand) for test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units.
+For a trained predictor $\widehat{\mu}$, we select (a) the highest $K$ values of $\widehat{\mu}(X_{j})$ and (b) the lowest $K$ values of $\widehat{\mu}(X_{j})$ among all test data.
+The results for (ii) top-K selection with both test and calibration data, as well as (iii) calibration-referenced selection are deferred to Supplementary Section C.1, showing similar patterns.
+For a confidence level $\alpha\in(0,1)$, we apply the vanilla conformal prediction and JOMI (both deterministic and randomized) to the selected units at level $\alpha$, with $\mathfrak{S}=2^{[m]}$.
+Since $|\widehat{\mathcal{S}}|$ is fixed, selection-conditional coverage implies FCR control at level $\alpha$ due to Proposition 2.
+In addition, we include the BY procedure (Benjamini and Yekutieli,, 2005) as a heuristic baseline.
+We vary $\alpha\in\{0.1,0.2,\dots,0.9\}$ and $K\in\{20,100,1000,2000,5000,10000,15000\}$, and set $V$ as the APS score (Romano et al.,, 2020).
+The results with another binary score is in Supplementary Section C.2.
+Figure 3 shows the empirical selection-conditional miscoverage in the left panel and the average prediction set size in the right panel, both at nominal levels $\alpha\in\{0.1,0.8\}$.
+The solid (resp.
+dashed) lines show the results when we select units with the highest (resp.
+lowest) predicted affinities.
+Figure 3: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for vanilla conformal prediction (Vanilla_CP), BY correction (BY), JOMI and randomized JOMI (JOMI_rand) for test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units.
 The reds dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
 The results are averaged over $N=1000$ runs.
-The orange curves (Vanilla_CP) show that vanilla conformal prediction with APS score is over-confident for units with the lowest predicted affinities while under-confident for units with highest predicted affinities.
+The orange curves (Vanilla_CP) show that vanilla conformal prediction with APS scores is over-confident for units with the lowest predicted affinities while under-confident for units with highest predicted affinities.
 In contrast, the purple (JOMI) and blue (JOMI_rand) curves both show valid coverage for our proposed methods.
-Also, using the APS score introduces visible gap between the actual coverage and $1-\alpha$ for JOMI due to distretization, which is made exact by its randomized version.
-Interestingly, vanilla conformal prediction with the APS score yields zero-cardinality prediction sets for units with lowest predicted affinities with $\alpha=0.8$ (the orange dashed line in the right panel of Figure 2).
-This is because vanilla CP covers other test units with very high rate, and thus marginal coverage is guaranteed even with empty prediction sets for selected units.
+Also, using the APS score introduces visible gap between the actual coverage and $1-\alpha$ for JOMI due to discretization, which is made exact by its randomized version.
+Finally, while BY has lower-than-nominal miscoverage which shows its validity as a heuristic method, it is overly conservative, leading to much larger prediction sets.
+Interestingly, vanilla conformal prediction with the APS score yields zero-cardinality prediction sets for units with the lowest predicted affinities with $\alpha=0.8$ (the orange dashed line in the right panel of Figure 3).
+This is because vanilla CP covers other test units with very high rate, and thus marginal coverage is guaranteed even with empty prediction sets for the selected units.
 Of course, this is worrying if one cares more about these units at the bottom.
-The selective behavior of these methods also depends on the choice of the nonconformity score $V$.
-Appendix A.2 shows the results for selection rules (i)-(iii) with the binary score.
-While our proposed methods consistently achieve valid coverage for selected units, with the binary score, vanilla CP is over-confident for units with highest predicted affinities yet under-confident for units with lowest predicted affinities.
-Also, the coverage gap of JOMI due to discretization is less visible with the binary score.
-5.1.2 Conformalized selection
-We then consider conformalized selection, where the focal test units are those believed to obey $Y=1$ with false discovery rate control at level $q\in(0,1)$.
+The behavior of these methods also depends on the choice of the nonconformity score $V$.
+Supplementary Section C.2 shows the results for selection rules (i)-(iii) with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$.
+With the binary score, JOMI consistently achieves valid coverage for selected units, and the coverage gap of JOMI due to discretization is less visible.
+For vanilla CP, opposite to the situations here, it is over-confident for units with the highest predicted affinities yet under-confident for units with the lowest predicted affinities.
+5.2 Conformal selection
+We then consider conformal selection, where the focal test units are those believed to obey $Y=1$ with false discovery rate control at level $q\in(0,1)$.
 This problem was investigated in Bao et al., (2024) with no exact finite-sample coverage guarantees in theory (though their heuristic method performs reasonably in their empirical studies).
-We apply conformalized selection (Jin and Candès, 2023b, ) at FDR level $q\in\{0.2,\dots,0.9\}$ and $c_{i}\equiv 0.5$ to determine $\widehat{\mathcal{S}}$, and construct prediction intervals for units in $\widehat{\mathcal{S}}$ using either the APS score or the binary score.
+We apply conformal selection (Jin and Candès, 2023b, ) at FDR level $q\in\{0.2,\dots,0.9\}$ and $c_{n+j}\equiv 0.5$ to determine $\widehat{\mathcal{S}}$, and construct prediction intervals for units in $\widehat{\mathcal{S}}$ using both the APS score and the binary score.
 Experiments are repeated for $N=1000$ independent runs.
-Figure 3 depicts the miscoverage $\widehat{\textnormal{Miscov}}$ and prediction interval lengths with nominal miscoverage level $\alpha\in\{0.1,0.8\}$ under various FDR levels $q$ and two choices of nonconformity score $V$.
-Figure 3: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by conformalized selection (for finding $Y=1$ at target FDR levels $q\in\{0.1,0.2,\dots,0.9\}$) when applying vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), randomized JOMI (JOMI_rand), the method of Bao et al., (2024) (Heuristic and Heuristic_rand).
+Figure 4 depicts $\widehat{\textnormal{Miscov}}$ and prediction interval lengths with nominal miscoverage level $\alpha\in\{0.1,0.8\}$ under various FDR levels $q$ and two choices of nonconformity score $V$.
+Figure 4: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by conformal selection at target FDR levels $q\in\{0.2,\dots,0.9\}$ when applying vanilla conformal prediction (Vanilla_CP), BY (BY), JOMI (JOMI), randomized JOMI (JOMI_rand), and the method of Bao et al., (2024) (Heuristic).
 Results with the APS score are in solid lines, while those with the binary score are in dashed lines.
-Details are otherwise as Figure 2.
-Vanilla CP is still less calibrated for selected units: it is over-confident with both scores for $\alpha=0.1$, while being over-confident with binary score and under-confident with APS score at $\alpha=0.8$.
+Details are otherwise the same as Figure 3.
+Vanilla CP is not calibrated for selected units: it is over-confident with both scores for $\alpha=0.1$, while being over-confident with binary score and under-confident with APS score at $\alpha=0.8$.
 In contrast, JOMI and JOMI_rand achieves valid selection-conditional coverage in all scenarios.
 There is some gap for JOMI with the APS score due to discretization, but not for JOMI_rand or the binary score.
-We observe an even lower empirical FCR than the conditional miscoverage for our methods (so they achieve valid FCR control); this is because the selection set can sometimes be empty for small values of FDR level $q$.
-Finally, compared with the heuristic methods of Bao et al., (2024), our method usually achieves smaller prediction set sizes whereas their method seems overly conservative.
-5.1.3 Selection with constraints
-We now consider selecting units with highest predicted binding affinities within a total budget of subsequent development.
-In this case, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})=\{j\in{\mathcal{D}}_{\textnormal{test}}\colon\widehat{\mu}(X_{j})\geq\widehat{\mu}_{[k]}\}$, where $\widehat{\mu}_{[k]}$ is the $k$-th order statistic of $\{\widehat{\mu}(X_{j})\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$, with $k=\max\{k\colon\sum_{\ell=k}^{m}L_{[k]}\leq C\}$, and $\{L_{j}\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$ are the costs.
+We observe an even lower empirical FCR than the conditional miscoverage for our methods (so they achieve valid FCR control); this is because the selection set can sometimes be empty for small values of FDR level $q$ (recall Proposition 3).
+Compared with the heuristic methods of Bao et al., (2024), our method usually achieves smaller prediction set sizes whereas their method seems overly conservative.
+We conjecture that this is due to a more delicate choice of the reference set.
+Finally, BY is also overly conservative despite valid empirical coverage.
+5.3 Selection with constraints
+We now consider selecting units with the highest predicted binding affinities within a total budget of subsequent development.
+In this case, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})=\{j\in[m]\colon\widehat{\mu}(X_{n+j})\geq\bar{\mu}\}$, where $\bar{\mu}=\max\{\mu\colon\sum_{j=1}^{m}L_{n+j}\mathds{1}\{\widehat{\mu}(X_{n+j})\geq\mu\}\leq C\}$, and $\{L_{n+j}\}_{j\in[m]}$ are the costs.
 We create semi-synthetic datasets since the original HIV data does not contain the cost information.
-Specifically, for each $i\in{\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i}))+2|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$ are i.i.d.
+Specifically, for each $i\in[n+m]$, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i}))+2|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$ are i.i.d.
 random variables that capture other cost-related information.
-After leaving $20\%$ as the training set, we randomly sample the data without replacement so that $n=2500$ and $m=2500$.
+Setting $20\%$ of the data aside as the training set, we randomly sample the data without replacement so that $n=2500$ and $m=2500$.
 The average miscoverage, prediction set size, and reference set size are reported in Figure 5.
@@
 From the right-most plot, we see that |^$|\widehat{\mathcal{R}}_{j}|$ is positively correlated with the number of selected test units.
-Figure 4: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=500$ independent runs of Vanilla_CP, JOMI, and JOMI_rand.
+As usual, BY is overly conservative.
+Figure 5: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for averaged reference set size {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in\widehat{{\mathcal{S}}}}$ versus $|\widehat{{\mathcal{S}}}|$ (right), across $N=500$ independent runs of Vanilla_CP, BY, JOMI, and JOMI_rand.
 The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
-5.2 Drug-target-interaction prediction
-We now turn to a regression task, drug-target-interaction prediction.
+Finally, we report the empirical FCR in the three tasks in Supplementary Section C.3.
+Consistent with our theory in Section 2.2, selection-conditional coverage implies FCR control in top-K selection, and the empirical FCR is also close to the nominal coverage level under conformal selection where the selection size is stable.
+However, in the constrained optimization task, we find that the empirical FCR of vanilla CP is much lower than the selection-conditional coverage.
+Similar to the ideas in Proposition 1, this is because the FCR is defined as zero when the selection set is empty.
+This shows the limitation of FCR as an error metric: even if it is controlled, on the “interesting” event that the selection set is non-empty, the actual coverage can be lower than anticipated.
+6 Application to health risk prediction
+Prediction machines are also widely used in healthcare for guiding clinical decision-making.
+These decisions may come from complicated underlying processes such as clinical resource optimization (Ahmadi-Javid et al.,, 2017; Master et al.,, 2017) or preliminary uncertainty quantification (Olsson et al.,, 2022).
+Accounting for selections from complicated decision processes is necessary for reliable and informative uncertainty quantification.
+In this section, we demonstrate the application of our framework to health risk prediction settings.
+We will consider three selection rules:
+  1. (1)
+    Selection with constraints.
+Clinical decision makers may optimize a performance measure subject to certain constraints such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014), and send patients to different care categories.
+In Section 6.1, we consider a stylized example where we minimize the total predicted ICU stay subject to total cost budget, where the selected units can be viewed as patients sent to a certain category.
+  2. (2)
+    Selecting small conformal prediction sets.
+Based on preliminary conformal prediction, one may suggest human intervention for units with large set sizes (Olsson et al.,, 2022; Sokol et al.,, 2024) while leaving those with small prediction sets unattended.
+In Section 6.2, we re-calibrate predictive inference for those whose preliminary prediction sets are shorter than a threshold.
+  3. (3)
+    Selection based on upper prediction bounds.
+Practitioners may also focus on units whose preliminary prediction intervals lie below a threshold (Levitskaya,, 2023).
+In Section 6.3, we study selected test units whose preliminary prediction intervals have upper bounds below a threshold.
+Among the above, rule (1) is covariate-dependent, while rules (2) and (3) can depend on the outcomes. All of them are efficiently tackled by the computation tricks in Section 4.
+Our experiments use the ICU-stay data in the MIMIC-IV dataset (Johnson et al.,, 2023).
+Data (and features in $X$) are pre-processed using the pipeline provided by Gupta et al., (2022), with the outcome $Y$ being the length of ICU stay.
+We use random forests in the scikit-learn Python package to train a point prediction model $\widehat{\mu}(\cdot)$ and two quantile regression models $\widehat{q}_{1-\beta/2}(\cdot)$ and $\widehat{q}_{\beta/2}(\cdot)$ using a holdout training set.
+Calibration and test data are randomly split with $n=3000$ and $m=2000$.
+As the discretization issue is minimal in this regression problem, we only present results for JOMI without randomization.
+6.1 Selection with constraints
+We first study the case where test units are selected by minimizing the total predicted ICU stay time subject to a budget constraint.
+Formally, for each patient $i\in[m+n]$, we let $\widehat{\mu}(X_{i})$ be its predicted ICU stay length, and $L_{i}>0$ be the budget needed for them.
+Then, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ aims to solve the following optimization problem:
+$\mathop{\text{maximize}}_{S\subseteq[m]}$ $\quad\sum_{j\in S}\widehat{\mu}(X_{n+j})$ (24)
+subject to $\quad\sum_{j\in S}L_{n+j}\leq\bar{L},$
+where $\bar{L}=200$ is a budget limit.
+Again, as the dataset does not come with drug development costs, we generate $L_{i}=\lceil\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}\rceil$, where $\bar{\mu}=\max_{i\in{\mathcal{D}}_{\textnormal{train}}}|\widehat{\mu}(X_{i})|$, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
+The optimization problem (24) is known as the Knapsack problem which is NP-hard.
+Nevertheless, there are efficient approximate solvers and we note that the validity of JOMI does not rely on exactness of the results; in our experiments, we use the Python package mknapsack (mknapsack,, 2023).
+Existing methods such as Bao et al., (2024) cannot deal with such a complicated selection process.
+In contrast, our framework tackles this problem with a computation complexity that is polynomial in $m$, $n$, and the complexity of the subroutine ${\mathcal{S}}(\cdot,\cdot)$.
+Figure 6 shows the empirical miscoverage, length of prediction interval, and sizes of the selection set and reference sets.
+While vanilla conformal prediction is over-confident and BY is overly conservative, our method achieves exact coverage for selected test units despite the complexity of the selection process.
+We also see a slightly positive correlation of |^$|\widehat{\mathcal{R}}_{n+j}|$ and $|\widehat{{\mathcal{S}}}|$.
+Figure 6: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for the averaged reference set size versus $|\widehat{{\mathcal{S}}}|$ (right), across $N=500$ independent runs of Vanilla_CP, BY, JOMI, and JOMI_rand.
+The $x$-axis of left and middle plots is the nominal level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
+6.2 Selecting small-sized prediction sets
+We then consider the second selection rule, where we first build preliminary conformal prediction intervals $\widehat{C}_{\alpha,n+j}^{\text{prelim}}$ via the score function $S(x,y)=|y-\widehat{\mu}(x)|/\widehat{\sigma}(x)$ (Lei et al.,, 2018); both the point prediction function $\widehat{\mu}(\cdot)$ and the conditional standard deviation are estimated via random forests.
+We then select those test units with $|\widehat{C}_{\alpha,n+j}^{\text{prelim}}|\leq 5$, i.e., the upper and lower bounds of the preliminary prediction intervals are less than $5$ days apart.
+This mimics the ideas in (Ren et al.,, 2023; Sokol et al.,, 2024) where small-sized prediction sets are “certified” as confident.
+After selection, we leverage the method in Section 4.3 to construct $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ for all selected test units.
+Since $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ is a superset of the exact output $\widehat{C}_{\alpha,n+j}$, we evaluate its empirical coverage to investigate whether it is over-conservative.
+Also, note from (22) that it is the union of three subsets; we also evaluate the number of disjoint segments in $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$.
+The miscoverage and length of prediction sets are reported in the left and middle plots in Figure 7.
+We observe that selectively certifying short prediction intervals can lead to under-coverage (orange curve), while JOMI achieves exact coverage (purple curve) by inflating the prediction sets, meaning that the superset $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ is effectively quite tight.
+The average number of disjoint segments in the right plot of Figure 7, which shows that $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ is almost always one single interval.
+Figure 7: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ (right), across $N=500$ runs of Vanilla_CP and JOMI when test units with short preliminary prediction sets are selected.
+The $x$-axis is nominal levels $\alpha\in\{0.1,0.2,\dots,0.9\}$.
+6.3 Selecting prediction sets below a threshold
+Finally, we study selection rule (iii) which is also based on a preliminary prediction set constructed in the same way as Section 6.2.
+We imagine that practitioners select a test unit $n+j$ if the upper bound of $\widehat{C}^{\text{pre}}_{\alpha,n+j}$ is below $6$, i.e., it appears the patient will stay in ICU less than 6 days.
+The miscoverage rate, length of prediction sets, and number of disjoint segments averaged over $N=500$ independent runs of JOMI (Section 4.3) and vanilla conformal prediction are summarized in Figure 8.
+We see that preliminary prediction sets with low upper bounds tend to under-cover for small $\alpha$, while JOMI achieves exact coverage despite that we construct a superset of $\widehat{C}_{\alpha,n+j}$.
+However, JOMI may produce multiple segments for large values of $\alpha$.
+Figure 8: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha,n+j}$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI.
+The nominal miscoverage levels on the $x$-axis are $\alpha\in\{0.1,0.2,\dots,0.9\}$.
+Finally, we present in Supplementary Section C.7 the tight empirical FCR control of JOMI in the above three tasks, since the size of the selection set is stable and nonzero.
+Reproducibility and data availability.
+The code for reproducing the numerical results in this paper is available at https://github.com/ying531/JOMI-paper.
+The data underlying this article are available online, with references given at appropriate places in the paper.
+Conflict of interest.
+The authors declare no conflict of interest.
+Acknowledgments.
+The authors thank the anonymous reviewers for their constructive comments.
+Z.R.
+acknowledges support from the National Science Foundation under grant DMS-2413135.
+Y.J.
+was partially supported by the Harvard Data Science Wojcicki-Troper Postdoctoral Fellowship.
+Appendix A Related works
+A.1 Literature overview
+Our method extends conformal prediction, which is a general framework for building marginally valid prediction intervals (Vovk et al.,, 2005; Lei et al.,, 2018; Romano et al.,, 2019, 2020).
+Compared with vanilla conformal prediction, we aim at selection-conditional coverage for a unit of interest, instead of marginal coverage for an “average” unit.
+Particularly related to our framework is Mondrian Conformal Prediction (Vovk et al.,, 2003, MCP), which uses a subset of calibration data to build a prediction set with coverage conditional on a pre-determined or permutation-equivariant class membership.
+Achieving selection-conditional coverage is a new goal with distinct motivations from MCP.
+Moreover, our techniques and scopes of application are significantly different from those of MCP (see Supplementary Section A.2).
+The selection-conditional guarantee we seek for is closely related to the post-selection inference (POSI) literature (Berk et al.,, 2013; Tibshirani et al.,, 2016; Chernozhukov et al.,, 2015) and in particular conditional selective inference (Lee et al.,, 2016; Markovic et al.,, 2017; Reid et al.,, 2017; Tibshirani et al.,, 2018; Kivaranovic and Leeb,, 2020; McCloskey,, 2024).
+In POSI, one important task is to build confidence intervals for selected model parameters conditional on a selection event, which is close to our selection-conditional coverage guarantee.
+Methods for this goal usually leverage specific problem structures such as linearity and distribution of the estimators (Zhong and Prentice,, 2008; Lee et al.,, 2016; Tian and Taylor,, 2017; Andrews et al.,, 2024; Liu,, 2023).
+As we focus on predictive inference (where the inferential targets are random variables instead of parameters), we leverage the exchangeability across units to achieve exact selection-conditional coverage, and thus our method substantially differs from POSI.
+Another related error notion in the selective inference literature is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), which is defined as the expected fraction of selected objects not covered by their confidence/prediction sets.
+Methods for achieving FCR control have been proposed for various settings (Benjamini and Yekutieli,, 2005; Weinstein et al.,, 2013; Weinstein and Ramdas,, 2020; Zhao and Cui,, 2020; Xu et al.,, 2022).
+In particular, Weinstein and Ramdas, (2020) introduce an online FCR-controlling method that can be applied to conformal prediction with selection.
+They consider sequentially arriving test units, and require the selection of a current unit to only depend on its prediction set and past selection decisions.
+In contrast, we allow the selection rule to simultaneously depend on all test units, such as top-K selection, optimization-based selection, and conformal p-value-based selection.
+Our approach to error control differs from theirs as well: while they achieve FCR control by adjusting the confidence level of marginal prediction sets, we leverage a judiciously chosen subset of calibration data to construct post-selection prediction sets with selection-conditional coverage.
+Another closely related work is Bao et al., (2024) who study prediction sets for selected units with (approximate) FCR control, yet they achieve exact guarantee only for a limited class of selection rules such as top-K selection.
+Later on, we will show that JOMI yields the same prediction sets as theirs when applied to top-K selection, and leads to exact guarantee for other problems they addressed with approximate FCR control.
+It also applies to many more settings with complicated selection rules that cannot be handled by their methods.
+Finally, a concurrent work of Gazin et al., (2024) (which was posted shortly after the first appearance of this work on arXiv) also studies the selective inference problem in conformal prediction.
+It specifically considers selection based on properties of prediction sets, but jointly calibrates the selection procedure and the construction of prediction sets while achieving FCR control.
+In contrast, we take a complementary perspective, decoupling the selection and uncertainty quantification steps, which allows the development of a versatile framework that is applicable to a broader range of selection rules.
+More broadly, many recent works have studied selective inference problems in conformal prediction, mostly focusing on multiple testing, i.e., selecting individuals that obey some properties with FDR control.
+This includes selecting outliers (Bates et al.,, 2023; Marandon et al.,, 2022; Liang et al.,, 2022; Bashari et al.,, 2024) and selecting high-quality data (Jin and Candès, 2023b, ; Jin and Candès, 2023a, ).
+The construction of conformal prediction sets is an orthogonal direction, and our methods also work for selection based on such procedures; see Section 4.2 for details.
+Under a similar name, Sarkar and Kuchibhotla, (2023) propose “post-selection inference for conformal prediction”, but they mainly focus on selecting the confidence level $\alpha\in(0,1)$ instead of selecting the units.
+In addition, they are closer to the “simultaneous inference” strand in POSI since they aim for valid inference simultaneously for all confidence levels, other than conditional on a selection event.
+For these reasons, our methods and guarantees are quite different.
+A.2 Connection to Mondrian Conformal Prediction
+Our framework is closely related to Mondrian Conformal Prediction (MCP) (Vovk et al.,, 2003) which applies to case of one test unit ($m=1$).
+Given $K$ categories $[K]$, MCP assumes access to a classifier ${\mathcal{K}}$ that gives the class label for all units: $(\kappa_{1},\kappa_{2},\dots,\kappa_{n+1})={\mathcal{K}}(Z_{1},Z_{2},\dots,Z_{n+1}),$ where $\kappa_{i}\in[K]$ is the class label for unit $i$.
+The classifier ${\mathcal{K}}$ is required to be equivariant to any permutation $\pi$ of $[n+1]$, i.e., $(\kappa_{\pi(1)},\kappa_{\pi(2)},\dots,\kappa_{\pi(n+1)})={\mathcal{K}}(Z_{\pi(1)},Z_{\pi(2)},\dots,Z_{\pi(n+1)}).$ The goal of MCP is to construct a prediction set for a test unit with category-wise coverage, i.e., $\mathbb{P}(Y_{n+1}\in\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}{\,|\,}\kappa_{n+1}=k)$ for each $k\in[K]$.
+This is achieved by finding a subset of calibration data that are categorized as $k$ for each hypothesized test outcome $y\in{\mathcal{Y}}$, i.e.,
+~$\widetilde{{\mathcal{R}}}(y)=\{i\in[n]:{\mathcal{K}}({\mathcal{D}}_{{\textnormal{calib}}},(X_{n+1},y))_{[i]}=k\},$ (25)
+and then constructing the prediction set as
+$\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+1},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widetilde{{\mathcal{R}}}(y)}\cup\{\infty\}\big{)}\Big{\}}.$
+While the setups seem different, JOMI effectively reduces to MCP when ${\mathcal{K}}$ is a binary classifier and ${\mathcal{S}}$ is equivariant under permutations of $[n+1]$.
+To see this, one can view the selection rule ${\mathcal{S}}$ as categorizing test points into two classes (selected or not selected), and selection-conditional coverage (3) as coverage given the “selected” category.
+In this way, one can check that our reference set ^$\widehat{{\mathcal{R}}}_{n+1}(y)$ coincides with ~$\widetilde{{\mathcal{R}}}(y)$ defined in (25), and thus the prediction sets also coincide.
+With a distinct goal of selection-conditional coverage, we additionally address the challenge that realistic selection rules are often asymmetric to calibration and (multiple) test points.
+To this end, JOMI generalizes MCP via a more delicate framework for finding exchangeable subgroups.
+As we jointly consider the calibration and test points when defining the reference set, we call our method JOint Mondrian Conformal Inference (JOMI).
+A.3 Comparison to the Benjamini-Yekutieli (BY) correction
+For post-selection inference, the Benjamini-Yekutieli (BY) procedure (Benjamini and Yekutieli,, 2005) is a natural solution, which is argued in prior works (Bao et al.,, 2024; Gazin et al.,, 2024) to heuristically control the FCR in practice.
+In the conformal prediction context, the BY prediction set is of the form
+$\widehat{C}_{\alpha,n+j}^{\textrm{BY}}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq\text{Quantile}\big{(}1-\alpha|\widehat{\mathcal{S}}|/m;\{V_{i}\}_{i\in{\mathcal{I}}_{{\textnormal{calib}}}})\cup\{\infty\}\Big{\}}.$
+which essentially produces conformal prediction sets at an adjusted level $\widetilde{\alpha}=\alpha|\widehat{\mathcal{S}}|/m$.
+As we demonstrate in our numerical experiments, the BY procedure indeed offers valid selection-conditional coverage empirically (though no formal theory is established in general).
+However, they are overly conservative in many cases, producing non-informative sets $\{0,1\}$ in binary classification problems, and prediction intervals with twice the lengths of JOMI.
+These results highlight the importance of leveraging the exchangeability structure unique in predictive inference for efficient inference.
+Apart from conservativeness, an issue with the BY prediction set is that it does not adapt to the nature of the selection.
+Note that given the covariate value of a test unit, $\widehat{C}_{\alpha,n+j}^{\textrm{BY}}$ depends on $\widehat{\mathcal{S}}$ only through its cardinality.
+Take top-K selection in a regression problem with $V(x,y)=|y-\widehat{\mu}(x)|$ as an example.
+In this case, no matter whether we select $K$ units with the highest predictions or the lowest predictions, the prediction set is always of the form $[\widehat{\mu}(x)-\widehat{\eta},\widehat{\mu}(x)+\widehat{\eta}]$ with $\widehat{\eta}$ being the $(1-\widetilde{\alpha})$-th quantile of all calibration scores.
+In contrast, by searching for “similar” calibration data in the reference set, our method provides more precise and targeted uncertainty quantification.
+Appendix B Additional results on covariate-dependent selection
+This subsection is devoted to stating the tailored solutions under covariate-dependent rules (1)-(3) in Section 4.1 of the main text.
+Throughout, we write $S_{i}=S(X_{i})$ and let ${\mathcal{T}}=\{S_{i}\colon i\in[n+m]\}$.
+B.1 Top-K selection
+In this example, we assume that $S_{n+j}$’s are distinct for all $j\in[m]$ almost surely for well-defined choice of top-K units. The top-K selection set is equivalently
+$\widehat{{\mathcal{S}}}_{\textnormal{topk}}=\{j\in[m]:S_{n+j}>T_{\textnormal{topk}}\},\quad\text{where}\quad T_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{j\in[m]}}\mathds{1}\{S_{n+j}\leq t\}\geq m-K\Big{\}}.$
+In this case, the reference set for each $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ takes a unified form:
+^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}\,:=\,\{i\in[n]:S_{i}>T_{\textnormal{topk}}\},$ (26)
+which is also agnostic to the choice of $\mathfrak{S}$.
+Intuitively, $T_{\textnormal{topk}}$ refers to the $(m-K)$-th smallest score among the test units; when $S_{n+j}$ is among the top $K$, replacing $S_{n+j}$ with another score above $T_{\textnormal{topk}}$ does not change the selection outcome.
+This intuition is formalized in the following proposition, with its proof deferred to Supplementary Section D.7.
+Proposition 6.
+With the top-K selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}\in 2^{[m]}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{topk}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ for all $y\in{\mathcal{Y}}$.
+Computing ^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ as a universal reference set reduces the computation complexity of the reference construction step in Algorithm 1 to $O(\max(m,n))$.
+B.2 Selection based on joint quantiles
+When selected units are those whose scores are above the $q$-th quantile of calibration and test scores, the selection set can be written as
+$\widehat{{\mathcal{S}}}_{\textnormal{jq}}=\{j\in[m]:S_{n+j}>T_{\textnormal{jq}}\},\quad\text{where}\quad T_{\textnormal{jq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{m+n}}\mathds{1}\{S_{i}\leq t\}\geq q(n+m)\Big{\}}$
+By definition, the selection threshold $T_{\textnormal{jq}}$ is invariant to the permutation of the calibration and test scores, and it is straightforward to check that for all $j\in\widehat{{\mathcal{S}}}_{\textnormal{jq}}$, the reference set is
+^$\widehat{{\mathcal{R}}}_{\textnormal{jq}}=\{i\in[n]:S_{i}>T_{\textnormal{jq}}\}.$
+Replacing ^$\widehat{{\mathcal{R}}}_{n+j}$’s with ^$\widehat{{\mathcal{R}}}_{jq}$ reduces the computation complexity of the reference set in Algorithm 1 to $O(\max(m,n))$.
+B.2.1 Selection based on calibration quantiles
+When selected units are those whose scores are greater than the $q$-th quantile of the calibration scores, the selection set is equivalently
+$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\big{\{}j\in[m]:S_{n+j}>T_{\textnormal{cq}}\big{\}},\quad\text{ where}\quad T_{\textnormal{cq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{n}}\mathds{1}\{S_{i}\leq t\}\geq qn\Big{\}}.$ (27)
+For each $j\in\widehat{{\mathcal{S}}}_{\textnormal{cq}}$, the reference set consists of all calibration units in the top $q$-th quantile, i.e.,
+^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\{i\in[n]:S_{i}>T_{\textnormal{cq}}\}.$ (28)
+The validity of ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ as the reference set is established below, whose proof is deferred to Supplementary Section D.8.
+Proposition 7.
+For the selection rule specified in (27), for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\widehat{{\mathcal{R}}}_{n+j}(y)$ defined in (8) for all $y\in{\mathcal{Y}}$.
+Similar to the previous two cases, directly computing ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ reduces the computation complexity of the reference set to $O(\max(m,n))$.
+As a side note, Bao et al., (2024) also considered the selection rule based on calibration quantiles (as an instance of “calibration-assisted selection”), for which they provide asymptotically FCR control guarantees under additional assumptions.
+Appendix C Additional experiment results
+This section collects additional experimental results that are omitted in Section 5 of the main text.
+Supplementary Section C.1 shows results for DPP with selection rules (ii) and (iii) with the APS score.
+Supplementary Section C.2 shows results for DPP with selection rules (i)-(iii) with the binary score.
+Supplementary Section C.5 shows results for DTI with selection rules (ii) and (iii).
+C.1 Additional results for APS score in drug property prediction
+In this part, we present results for DPP when test units are selected via (ii) top-K in both calibration and test data (Supplementary Section C.1.1), and (iii) only calibration data as reference (Supplementary Section C.1.2).
+C.1.1 Top-K selection in mixed sample
+We use the same values of $(\alpha,K)$ and the same scheme of repeated experiments as in Section 5.1 of the main text, with $\mathfrak{S}=2^{[m]}$.
+Figure 9 shows that vanilla conformal prediction incurs a large coverage gap for selected units, while JOMI and its randomized version both achieve valid selection-conditional coverage.
+The coverage gap due to discretization is eliminated again by randomization.
+Randomization also reduces the prediction set sizes on the right panel, but again, with $\alpha=0.8$, we see the concerning issue that it suffices for vanilla CP to construct empty prediction sets for selected units to achieve marginal coverage.
+Figure 9: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units whose $\widehat{\mu}(X_{n+j})$ are top-K (highest, solid line) or bottom-K (lowest, dashed line) among both calibration and test units.
+Details are otherwise the same as Figure 3.
+C.1.2 Calibration-referenced selection
+Now we select test units whose predicted affinities must be greater (or smaller) than the $K$-th largest (or smallest) predictions in the calibration set.
+We use the choices of $(\alpha,K)$ and the scheme of repeated experiments as in Section 5.1 of the main text with $\mathfrak{S}=2^{[m]}$.
+Figure 10: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by (iii) whose $\widehat{\mu}(X_{n+j})$ are greater than the $K$-th highest (solid line) or smaller than the $K$-th lowest (dashed line) in calibration units.
+Details are otherwise as Figure 3.
+Figure 10 shows that our proposed methods achieve valid coverage (exact with JOMI_rand) while vanilla conformal prediction fails to.
+Again, vanilla CP is either over-confident or under-confident for units with lowest or highest predicted affinities.
+C.2 Additional results for binary score in drug property prediction
+In Figures 1112, and 13, we show the empirical selection-conditional coverage and prediction set sizes for selection rules (i)-(iii) when the prediction sets are constructed with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$.
+We observe that while vanilla CP is either over-confident or under-confident, our proposed methods achieve valid coverage across all configurations.
+The coverage gap of the non-randomized JOMI is smaller with this specific score function compared with the APS score.
+Figure 11: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (i) top-K (or bottom-K) in test data.
+Details are otherwise as Figure 3.
+Figure 12: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (ii) top-K (or bottom-K) in both calibration and test data.
+Details are otherwise as Figure 9.
+Figure 13: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units that are (iii) greater than the $K$-th highest (or smaller than the $K$-th lowest) predicted affinities in the calibration data.
+Details are as Figure 10.
+C.3 Additional results for empirical FCR in drug property prediction
+In this part, we present empirical FCR results in drug property prediction, which complement the three tasks in Section 5 in the main text.
+Figure 14 plots the empirical FCR in DPP in three top-K selection tasks with two conformity scores (APS and binary) at nominal miscoverage level $\alpha=0.1$.
+Since the selection set size is fixed, by our theory, selection-conditional coverage implies FCR control, as evidence from the results.
+Figure 15 plots the empirical FCR when the units are selected by the conformal selection algorithm in Jin and Candès, 2023b .
+Here, the empirical FCR of JOMI is sill very close to the nominal conditional coverage level, because the selection set size is stable.
+Finally, Figure 16 plots the empirical FCR when the units are selected by running the constrained optimization program.
+While the empirical FCR of JOMI is close to the nominal conditional coverage level, the FCR of vanilla CP with both conformity scores is lower than the selection-conditional coverage.
+This is because the selection set can sometimes be empty, and FCR is considered on this event.
+This also shows the limitation of FCR as an overly “optimistic” error metric: even if the FCR is low, on the “interesting” event that the selection set is nonzero, the coverage might still be very low.
+Figure 14: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in top-K selection for drug property prediction.
+The $x$-axis is the nominal miscoverage level $\alpha=0.1$.
+Figure 15: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in drug property prediction when units are selected by FDR-controlled conformal selection.
+The $x$-axis is the nominal FDR level in the selection algorithm.
+Figure 16: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in drug property prediction when units are selected by running the optimization program.
+The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
+C.4 Application to drug-target-interaction prediction
+In this part, we include experiments in drug-target-interaction prediction tasks omitted in the main text.
 We use the DAVIS dataset in the DeepPurpose library (Huang et al.,, 2020).
@@
 We use the nonconformity score $V(x,y)=|y-\widehat{\mu}(x)|$ throughout.
-5.2.1 Conformalized selection for heterogeneous thresholds
-We then study the behavior of vanilla CP and our proposed methods for selected units by running conformalized selection (Jin and Candès, 2023b, ) to find test pairs $j\in{\mathcal{I}}_{\textnormal{test}}$ whose outcomes are greater than data-dependent thresholds $c_{j}$.
-Here, we set $c_{j}$ as the $0.7$-th quantile of activities for training pairs with the same target sequence as the $j$-th data, while controlling the FDR below $q\in\{0.2,0.3,\dots,0.9\}$.
+C.4.1 Conformalized selection for heterogeneous $c_{n+j}$’s
+We then study the behavior of vanilla CP and our proposed methods for selected units by running conformalized selection (Jin and Candès, 2023b, ) to find test pairs $j\in[m]$ whose outcomes are greater than data-dependent thresholds $c_{n+j}$.
+Here, we set $c_{n+j}$ as the $0.7$-th quantile of activities for training pairs with the same target sequence as the $j$-th data, while controlling the FDR below $q\in\{0.2,0.3,\dots,0.9\}$.
 The method of Bao et al., (2024) is not applicable to the heterogeneous thresholds here.
@@
 Again, we observe that the FCR of our methods is even lower than $\widehat{\textnormal{Miscov}}$, empirically yielding valid FCR control.
-Figure 5: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug-target-interaction for vanilla conformal prediction (Vanilla_CP) and JOMI (JOMI and JOMI_rand), at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for test units selected by conformalized selection at nominal FDR levels $q\in\{0.1,0.2,\dots,0.9\}$.
+Figure 17: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug-target-interaction for vanilla conformal prediction (Vanilla_CP) and JOMI (JOMI and JOMI_rand), at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for test units selected by conformalized selection at nominal FDR levels $q\in\{0.2,\dots,0.9\}$.
 Our method can sometimes lead to two separate prediction intervals.
@@
 Figure 18: Boxplot of the average number of intervals produced by JOMI and JOMI_rand among selected test units, over $N=1000$ independent runs, at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for various experiment configurations.
-5.2.2 Selection under budget constraint
+C.4.2 Selection under budget constraint
 Finally, we consider a covariate-dependent selection rule, where the scientist aims to maximize the predicted activities of selected drugs subject to a fixed budget for subsequent development.
@@
 Then, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ aims to solve the following optimization problem:
-$\mathop{\text{maximize}}_{S\subseteq\{n+1,\dots,m+m\}}$ $\quad\sum_{j\in S}\big{[}\widehat{\mu}(X_{j})-c_{j}\big{]}$ (25)
-subject to $\quad\sum_{j\in S}L_{j}\leq\bar{L},$
+$\mathop{\text{maximize}}_{S\subseteq[m]}$ $\quad\sum_{j\in S}\big{[}\widehat{\mu}(X_{n+j})-c_{n+j}\big{]}$ (29)
+subject to $\quad\sum_{j\in S}L_{n+j}\leq\bar{L},$
 where $\bar{L}=200$ is a budget limit.
-Again, as the dataset does not come with drug development costs, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
+Again, as the dataset does not come with drug development costs, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}$, where $\bar{\mu}=\max_{i\in{\mathcal{D}}_{\textnormal{train}}}|\widehat{\mu}(X_{i})|$, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
 The optimization problem (29) is known as the Knapsack problem which is NP-hard.
@@
 The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
-6 Application to health risk prediction
-Prediction machines are also widely used in healthcare for guiding clinical decision-making.
-These decisions may be determined by some complicated underlying processes such as clinical resource optimization (Ahmadi-Javid et al.,, 2017; Master et al.,, 2017) or preliminary uncertainty quantification (Olsson et al.,, 2022), and once made, would have substantial consequences.
-In this section, we demonstrate the application of our framework to health risk prediction settings where units are picked by some complicated underlying processes.
-We will consider three selection rules:
-  1. (1)
-    Selection with constraints.
-Clinical decision makers may optimize a performance measure subject to certain constraints such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014), and send patients to different care categories.
-In Section 6.1, we consider a stylized example where we minimize the total predicted ICU stay subject to total cost budget.
-  2. (2)
-    Selecting small conformal prediction sets.
-Based on preliminary conformal prediction, one may suggest human intervention for units with large set sizes (Olsson et al.,, 2022; Sokol et al.,, 2024) while leaving those with small prediction sets unattended.
-In Section 6.2, we re-calibrate predictive inference for those whose preliminary prediction sets are shorter than a threshold.
-  3. (3)
-    Selecting lower prediction bounds.
-Practitioners may also focus on units whose preliminary prediction intervals lie above a threshold (Levitskaya,, 2023).
-In Section 6.3, we study selected test units whose preliminary prediction intervals have lower bounds above a threshold.
-Among the above, (1) is covariate-dependent, and (2) and (3) depend on some preliminary uncertainty quantification criterion.
-All of them are efficiently tackled by the computation tricks in Section 4.
-Our experiments use the ICU-stay data in the MIMIC-IV dataset (Johnson et al.,, 2023).
-Data (and features in $X$) are pre-processed using Gupta et al., (2022), with the outcome $Y$ being the length of ICU stay.
-We use random forests in the scikit-learn Python package to train a point prediction model $\widehat{\mu}(\cdot)$ and two quantile regression models $\widehat{q}_{1-\beta/2}(\cdot)$ and $\widehat{q}_{\beta/2}(\cdot)$ using a heldout training set.
-Calibration and test data are randomly split with $n=3000$ and $m=2000$.
-6.1 Selection with constraints
-We first study the case where test units are selected by minimizing the total predicted ICU stay time subject to a budget constraint.
-Given the point predictor $\widehat{\mu}(\cdot)$, we solve a Knapsack problem similar to (25) using the same approximate algorithm, where we maximize the summation of $\widehat{\mu}(X_{j})$ in the selected subset, and the cost $L_{j}$ is generated in the same way as Section 5.2.2 and then taken the ceiling.
-Figure 8 shows the empirical miscoverage rate, length of prediction interval, and sizes of the selection set and reference sets.
-While vanilla conformal prediction is over-confident, our method achieves exact coverage for selected test units.
-We also see a slightly positive correlation of |^$|\widehat{\mathcal{R}}_{j}|$ and $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$.
-Figure 8: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=500$ independent runs of Vanilla_CP, JOMI, and JOMI_rand.
-The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
-6.2 Selecting small-sized prediction sets
-We then consider the second selection rule, where we first build preliminary conformal prediction intervals $\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})$ via the score function $S(x,y)=|y-\widehat{\mu}(x)|/\widehat{\sigma}(x)$ (Lei et al.,, 2018), and both point prediction $\widehat{\mu}(\cdot)$ and conditional standard deviation estimation are trained via random forests.
-We then select those test units with $|\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})|\leq 5$, i.e., the upper and lower bounds of the preliminary prediction intervals are less than $5$ days apart.
-This mimics the ideas in (Ren et al.,, 2023; Sokol et al.,, 2024) where small-sized prediction sets are “certified” as confident.
-After selection, we leverage the method in Section 4.3 to construct $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ for all selected test units $n+j$.
-Since $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is a superset of the exact output $\widehat{C}_{\alpha}(X_{n+j})$, we evaluate its empirical coverage to investigate whether it is over-conservative.
-Also, note from (23) that it is the union of three subsets; we also evaluate the number of disjoint segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$.
-The miscoverage and length of prediction sets are reported in the first two plots in Figure 9.
-We observe that selectively certifying short prediction intervals can lead to under-coverage (orange curve), while JOMI achieves exact coverage (purple curve) by inflating the prediction sets, meaning that the superset $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is quite tight.
-The average number of disjoint segments in the right plot of Figure 9, which shows that $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is always one single interval.
-Figure 9: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI when test units with short preliminary prediction sets are selected.
-The $x$-axis are nominal miscoverage levels $\alpha\in\{0.1,0.2,\dots,0.9\}$.
-6.3 Selecting high prediction sets
-Finally, we study selection rule (iii) which is also based on a preliminary prediction set constructed in the same away as Section 6.2.
-We imagine that practitioners select a test unit $n+j$ if the upper bound of $\widehat{C}^{\text{pre}}(X_{n+j})$ is below $6$, i.e., it appears the patient will stay in ICU less than 6 days.
-The miscoverage rate, length of prediction sets, and number of disjoint segments averaged over $N=500$ independent runs of JOMI (using the method in Section 4.3) and vanilla conformal prediction are summarized in Figure 10.
-We see that preliminary prediction sets with low upper bounds tend to under-cover for small $\alpha$, while JOMI achieves exact coverage despite that we construct a superset of $\widehat{C}_{\alpha}(X_{n+j})$.
-However, JOMI may produce multiple segments for large values of $\alpha$.
-Figure 10: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI.
-The nominal miscoverage levels on the $x$-axis are $\alpha\in\{0.1,0.2,\dots,0.9\}$.
-Appendix A Additional experiment results
-This section collects additional experimental results that are omitted in the main text.
-Appendix A.1 shows results for DPP with selection rules (ii) and (iii) with the APS score, omitted in Section 5.1.
-Appendix A.2 shows results for DPP with selection rules (i)-(iii) with the binary score.
-Appendix A.3 shows results for DTI with selection rules (ii) and (iii) that are omitted in Section 5.2.
-A.1 Additional results for APS score in drug property prediction
-In this part, we present results for DPP when test units are selected via (ii) top-K in both calibration and test data (Appendix A.1.1), and (iii) only calibration data as reference (Appendix A.1.2).
-A.1.1 Top-K selection in mixed sample
-We use the same values of $(\alpha,K)$ and the same scheme of repeated experiments as in Section 5.1.1, with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$.
-Figure 11 shows that vanilla conformal prediction incurs a large coverage gap for selected units, while JOMI and its randomized version both achieve valid selection-conditional coverage.
-The coverage gap due to discretization is eliminated again by randomization.
-Randomization also reduces the prediction set sizes on the right panel, but again, with $\alpha=0.8$, we see the concerning issue that it suffices for vanilla CP to construct empty prediction sets for selected units to achive marginal coverage.
-Figure 11: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units whose $\widehat{\mu}(X_{n+j})$ are top-K (highest, solid line) or bottom-K (lowest, dashed line) among both calibration and test units.
-Details are otherwise the same as Figure 2.
-A.1.2 Calibration-referenced selection
-Now we select test units whose predicted affinities must be greater (or smaller) than the $K$-th largest (or smallest) predictions in the calibration set.
-We use the choices of $(\alpha,K)$ and the scheme of repeated experiments as in Section 5.1.1 with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$.
-Figure 12: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by (iii) whose $\widehat{\mu}(X_{n+j})$ are greater than the $K$-th highest (solid line) or smaller than the $K$-th lowest (dashed line) in calibration units.
-Details are otherwise as Figure 2.
-Figure 12 shows that our proposed methods achieve valid coverage (exact with JOMI_rand) while vanilla conformal prediction fails to.
-Again, vanilla CP is either over-confident or under-confident for units with lowest or highest predicted affinities.
-A.2 Additional results for binary score in drug property prediction
-In Figures 1314, and 15, we show the empirical selection-conditional coverage and prediction set sizes for selection rules (i)-(iii) when the prediction sets are constructed with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$.
-We observe that while vanilla CP is either over-confident or under-confident, our proposed methods achieve valid coverage across all configurations.
-The coverage gap of the non-randomized JOMI is smaller with this specific score function compared with the APS score.
-Figure 13: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (i) top-K (or bottom-K) in test data.
-Details are otherwise as Figure 2.
-Figure 14: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units selected by (ii) top-K (or bottom-K) in both calibration and test data.
-Details are otherwise as Figure 11.
-Figure 15: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units that are (iii) greater than the $K$-th highest (or smaller than the $K$-th lowest) predicted affinites in the calibration data.
-Details are as Figure 12.
-A.3 Additional results for drug-target-interaction prediction
+The empirical FCR in the two tasks is reported in Supplementary Section C.6 showing similar messages as in drug property prediction tasks.
+While JOMI consistently controls the empirical FCR, it can in general be lower than the selection-conditional coverage when the selection set may be empty.
+C.5 Additional results for drug-target-interaction prediction
 In this section, we collect results for DTI with (1) covariate dependent top-K selection.
-We present results under selection rule (1-i) in Appendix A.3.1, (1-ii) in Appendix A.3.2 and (1-iii) in Appendix A.3.3.
-A.3.1 Top-K selection for fixed K
+We present results under selection rule (1-i) in Supplementary Section C.5.1, (1-ii) in Supplementary Section C.5.2 and (1-iii) in Supplementary Section C.5.3.
+C.5.1 Top-K selection for fixed K
 We consider the case where test units are selected via (i) top-K (or bottom-K) among predicted affinities in the test set.
-We apply the vanilla conformal prediction and our proposed methods with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$, so that selection-conditional coverage implies FCR control.
-Results under selection rules (ii) and (iii) demonstrate similar patterns, which we defer to Appendix A.3.
+We apply the vanilla conformal prediction and our proposed methods with $\mathfrak{S}=2^{[m]}$, so that selection-conditional coverage implies FCR control.
+Results under selection rules (ii) and (iii) demonstrate similar patterns, which are in Supplementary Section C.5.
 Figure 20 shows $\widehat{\textrm{Miscov}}$ (the left panel) and the average length of the prediction intervals (the right panel) for $K\in\{20,100,1000,2000,5000,10000\}$ and nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
-We see that vanilla conformal prediction is over-confident for pairs with the highest predicted affinites, while being under-confident for the lowest.
+We see that vanilla conformal prediction is over-confident for pairs with the highest predicted affinities, while being under-confident for the lowest.
 In contrast, our proposed methods achieve near-exact coverage under different nominal levels.
-By comparing the solid and dashed lines in the right panel, we observe that in this problem, the intrinstic uncertainty in higher-predicted-affinity pairs is larger, so a longer prediction interval is needed to achieve certain coverage.
-Figure 16: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), and randomized JOMI applied to test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units.
+By comparing the solid and dashed lines in the right panel, we observe that in this problem, the intrinsic uncertainty in higher-predicted-affinity pairs is larger, so a longer prediction interval is needed to achieve certain coverage.
+Finally, similar to DPP, BY is overly conservative despite valid coverage.
+The right panel reveals another drawback of BY: it doesn not adapt to the nature of selection.
+More specifically, by definition, it only depends on the $\alpha|\widehat{\mathcal{S}}|/m$-th quantile of calibration scores.
+As such, it only relies on the size of the selection set, neglecting the nature of the selection (i.e., whether we select highest predictions or lowerst predictions).
+In both cases, it returns a prediction set of the form $[\widehat{\mu}(X_{n+j})-\widehat{\eta},\widehat{\mu}(X_{n+j})+\widehat{\eta}]$ where $\widehat{\eta}$ remains the same between two ranking options.
+In contrast, by finding “similar” calibration data in the reference set, JOMI provides more precise and targeted uncertainty quantification.
+Figure 20: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for vanilla conformal prediction (Vanilla_CP), BY (BY), JOMI (JOMI), and randomized JOMI applied to test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units.
 The red dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
-A.3.2 Top-K selection in mixed sample
+C.5.2 Top-K selection in mixed sample
 Figure 21 shows the miscoverage and average length of prediction intervals when (ii) the focal test units are those whose predicted affinities are among the $K$ highest or lowest in both calibration and test data.
-We see that vanilla CP is over-confident for highest-prediction units, while under-confident for lowest-prediction units.
+We see that vanilla CP is over-confident for highest-prediction units, while under-confident for lowest-prediction units; BY is overly conservative in both cases.
 In contrast, our proposed methods always yield valid coverage.
@@
 Details are otherwise the same as Figure 20.
-A.3.3 Calibration-referenced selection
+C.5.3 Calibration-referenced selection
 Figure 22 shows the results with (iii) calibration-referenced selection, i.e., the selected test units are those whose predicted affinities are greater than (or smaller than) the $K$-th highest (or lowest) in the calibration data.
 Similar to rules (i) and (ii), vanilla CP is not calibrated for these focal units, while our methods yield near-exact coverage across the whole spectrum of $K$.
-Figure 18: Empirical FCR in drug-target interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ surpass the highest $K$ predicted affinities in the calibration data (or smaller than the lowest).
+As usual, BY is overly conservative.
+Figure 22: Empirical miscoverage (left) and prediction interval length (right) in drug-target interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ surpass the highest $K$ predicted affinities in the calibration data (or smaller than the lowest).
 Details are otherwise the same as Figure 20.
-Appendix B Technical proofs
-B.1 Proof of Proposition 1
-[…]
-$\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j}){\,|\,}j\in\widehat{\mathcal{S}})$ $=\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j})\,\big{|}\,j\in\widehat{\mathcal{S}},\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}\big{)}\cdot\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})$
-$\geq(1-\alpha)\cdot\sum_{\ell\in{\mathcal{L}}}\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})=1-\alpha,$
-where we use the fact that $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint and their union is the full set. Thus, it satisfies the weak selection-conditional coverage (2).
-B.2 Proof of Proposition 2
-Suppose $\widehat{C}_{\alpha}$ satisfies the strong selection-conditional coverage (3) for a class of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, it holds that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$.
+C.6 Additional results for empirical FCR in drug target interaction
+In this part, we collect results for the empirical FCR in drug-target interaction tasks in Section C.4.
+Figure 23 plots the empirical FCR when units are selected by running conformalized selection.
+We observe lower empirical FCR than selection-conditional coverage, demonstrating limitations of FCR as an error metric similar to the message in Section C.3.
+Figure 24 plots the empirical FCR when units are selected by constrained optimization.
+We observe tight empirical FCR below the nominal level, and in these experiments the selection set is always nonempty.
+Figure 23: Empirical FCR in drug-target interaction prediction when units are selected by conformal selection with heterogeneous $c_{n+j}$.
+The $x$-axis is the nominal FDR level in the selection program.
+Figure 24: Empirical FCR in drug-target interaction prediction when units are selected by constrained optimization. The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
+C.7 Additional results for empirical FCR in health risk prediction
+Figure 25 presents the empirical FCR for three tasks in health risk prediction. In these examples, the selection set sizes are stable and nonempty, hence the empirical FCR is also tightly controlled.
+Figure 25: Empirical FCR in health risk prediction when units are selected by constrained optimization (left), length of preliminary prediction interval (middle), and lower bound of preliminary prediction interval (right).
+The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
+Appendix D Technical proofs
+D.1 Proof of Proposition 1
+Suppose $\widehat{C}_{\alpha,n+j}^{(\ell)}$ satisfies $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and we let $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ whenever $j\in\widehat{\mathcal{S}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$.
+This is well-defined because $\mathfrak{S}_{\ell}$’s are disjoint.
+Then
+$\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}})$ =$\stackrel{{\scriptstyle\textrm{(a)}}}{{=}}\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)},\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\,\big{|}\,j\in\widehat{\mathcal{S}}\big{)}$
+$=\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}\,\big{|}\,j\in\widehat{\mathcal{S}},\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}\big{)}\cdot\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})$
+$\stackrel{{\scriptstyle\textrm{(b)}}}{{\geq}}(1-\alpha)\cdot\sum_{\ell\in{\mathcal{L}}}\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})=1-\alpha,$
+where in the first step we use the fact that $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint and their union is the full set, and the second step follows from the strong selection-conditional coverage.
+By definition, it satisfies the weak selection-conditional coverage (2).
+D.2 Proof of Proposition 2
+Suppose $\widehat{C}_{\alpha,n+j}^{(\ell)}$ satisfies the strong selection-conditional coverage (3) for a class of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, it holds that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$.
 By definition, and since $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint, we have
-FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}(\widehat{{\mathcal{S}}}),\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}$ (26)
-$=\sum_{r=1}^{m}\sum_{\ell\in{\mathcal{L}}}\mathbb{E}\bigg{[}\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\cdot\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}.$ (27)
+FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}$ (30)
+$=\sum_{\ell\in{\mathcal{L}}}\mathbb{E}\bigg{[}\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\cdot\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}^{(\ell)}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}.$ (31)
 Noting that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ hence $|\widehat{{\mathcal{S}}}|=r(\ell)$ when $\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}$, we can rewrite the above as
-$\sum_{\ell\in{\mathcal{L}}}\sum^{m}_{j=1}\frac{1}{r(\ell)}\mathbb{E}\big{[}\mathds{1}\big{\{}\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell},j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\big{\}}\bigg{]}$ (28)
+$\sum_{\ell\in{\mathcal{L}}}\sum^{m}_{j=1}\frac{1}{r(\ell)\vee 1}\mathbb{E}\big{[}\mathds{1}\big{\{}\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell},j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}^{(\ell)}\big{\}}\big{]}$
 […]
@@
 where the last inequality follows from the strong selection-conditional coverage. Finally, the FCR control proof is completed by noting that
-(28)$=\sum_{\ell\in{\mathcal{L}}}\frac{\alpha}{r(\ell)}\mathbb{E}\Big{[}|\widehat{{\mathcal{S}}}|\cdot\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\Big{]}=\alpha\cdot\mathbb{P}\big{(}\widehat{{\mathcal{S}}}\neq\varnothing\big{)}\leq\alpha.$
-B.3 Proof of Proposition 3
-Consider the following example with $m=2$ test units and ${\mathcal{Y}}=\mathbb{R}$. Suppose $\alpha\in(0,1)$. The selection rule returns
+(D.2)$=\sum_{\ell\in{\mathcal{L}}}\frac{\alpha}{r(\ell)\vee 1}\mathbb{E}\Big{[}|\widehat{{\mathcal{S}}}|\cdot\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\Big{]}=\alpha\cdot\mathbb{P}\big{(}\widehat{{\mathcal{S}}}\neq\varnothing\big{)}\leq\alpha.$
+D.3 Proof of Proposition 3
+Consider the following example with $m=2$ test units and ${\mathcal{Y}}=\mathbb{R}$. Suppose the target miscoverage level $\alpha\in(0,1)$, and the selection rule returns
 $\widehat{{\mathcal{S}}}=\begin{cases}\{1,2\}&\text{with probability }1-\frac{2\alpha}{1+\alpha},\\ \{1\}&\text{with probability }\frac{\alpha}{1+\alpha},\\ \{2\}&\text{with probability }\frac{\alpha}{1+\alpha}.\end{cases}$
-Consider a prediction set $\widehat{C}_{\alpha}$ such that
-$\widehat{C}_{\alpha}\big{(}X_{n+1};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+1};\{1\}\big{)}=\varnothing,\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{2\}\big{)}=\varnothing.$
-Since the construction is symmetric in $1$ and $2$, we focus on $j=1$ without loss of generality.
+Consider prediction sets $\widehat{C}_{\alpha,n+1}$ and $\widehat{C}_{\alpha,n+2}$ such that
+  • when $\widehat{{\mathcal{S}}}=\{1,2\}$, $\widehat{C}_{\alpha,n+1}=\mathbb{R},\widehat{C}_{\alpha,n+2}=\mathbb{R}$;
+  • when $\widehat{{\mathcal{S}}}=\{1\}$, $\widehat{C}_{\alpha,n+1}=\varnothing$;
+  • when $\widehat{{\mathcal{S}}}=\{2\}$, $\widehat{C}_{\alpha,n+2}=\varnothing$.
+We can check that the weak selection-conditional coverage is satisfied for $j=1$:
 […]
-For the FCR, we have
+By symmetry, the weak selection-conditional coverage is also satisfied for $j=2$. Meanwhile, one can check that
 $\text{FCR}=\frac{\alpha}{1+\alpha}+\frac{\alpha}{1+\alpha}=\frac{2\alpha}{1+\alpha}>\alpha.$
-Therefore, the constructed prediction set $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage but violates the FCR control.
-B.4 Proof of Theorem 1
-Proof of (a).
-To prove (11), it suffices to show that
-$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{j}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S}\Big{)}\geq 1-\alpha.$ (32)
-Recall that ${\mathcal{D}}_{j}={\mathcal{D}}_{{\textnormal{calib}}}\cup\{Z_{n+j}\}$, and ${\mathcal{D}}_{j}^{c}=\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}}$ is the set of test data without the $j$-th one; ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$ and ${\mathcal{I}}_{j}^{c}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$.
+Therefore, the constructed prediction sets $\widehat{C}_{\alpha,n+1}$ and $\widehat{C}_{\alpha,n+2}$ satisfy the weak selection-conditional coverage but violate the FCR control.
+D.4 Proof of Theorem 1
+Recall that ${\mathcal{D}}_{j}={\mathcal{D}}_{{\textnormal{calib}}}\cup\{Z_{n+j}\}$ and ${\mathcal{D}}_{j}^{c}=\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$.
+For notational convenience, we also denote the index sets ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$ and ${\mathcal{I}}_{j}^{c}=[m]\backslash\{j\}$.
 Let $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ be the unordered set of ${\mathcal{D}}_{j}$, and we denote the realized value of the unordered set as $[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$, i.e., $z_{i}$’s are fixed realized values of $Z_{i}$’s, but we do not know which values in $[d_{j}]$ correspond to the calibration set and which to the test point.
-We also write $z_{i}=(x_{i},y_{i})$ and $v_{i}=V(x_{i},y_{i})$.
-The exchangeability of $\{Z_{i}\}_{i\in{\mathcal{I}}_{j}}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$ implies that, given knowledge of the unordered realized data values in $[d_{j}]$, the probability of any assignment of these values to data points is equal.
+We also write $z_{i}=(x_{i},y_{i})$ and $v_{i}=V(x_{i},y_{i})$, with $x_{i}$ and $y_{i}$ being the realized values of $X_{i}$ and $Y_{i}$.
+We then prove the two statements separately.
+Proof of (a). To prove (11), it suffices to show that
+$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S}\Big{)}\geq 1-\alpha.$ (33)
+The exchangeability of data in ${\mathcal{D}}_{j}$ given ${\mathcal{D}}_{j}^{c}$ implies that, given knowledge of the unordered realized data values in $[d_{j}]$, the probability of any assignment of these values to the data points is equal.
 That is, for any permutation $\pi$ of ${\mathcal{I}}_{j}$,
 $\mathbb{P}\big{(}Z_{1}=z_{\pi(1)},\dots,Z_{n}=z_{\pi(n)},Z_{n+j}=z_{\pi(n+j)}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{1}{(n+1)!}.$
-Next, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we define the following subset of ${\mathcal{I}}_{j}$:
+Next, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we define the following subset of ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$:
 $\widetilde{R}=\big{\{}i\in{\mathcal{I}}_{j}\colon j\in{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i}),\,{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\in\mathfrak{S}\big{\}},$
-where $[D_{\textnormal{calib}}^{i}]=[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ is an unordered set, and $D_{\textnormal{test}}^{i}=(X_{n+1},\dots,x_{i},\dots,X_{n+m})$.
-Note that the collection $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, so $\widetilde{R}$ is also fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
-Recall that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})=\{i\in{\mathcal{I}}_{\textnormal{calib}}:j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\}$.
-We let ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})$ if $j\notin\widehat{\mathcal{S}}$ and ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}$ if $j\in\widehat{\mathcal{S}}$.
+where $[D_{\textnormal{calib}}^{i}]=[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ is an unordered set ($[d_{j}]$ with $z_{i}$ removed) and $D_{\textnormal{test}}^{i}=(X_{n+1},\dots,x_{i},\dots,X_{n+m})$, i.e., the test data with $X_{n+j}$ replaced by $x_{i}$.
+Also, $[D_{\textnormal{calib}}^{n+j}]=[z_{1},\dots,z_{n}]$.
+We note that a set is always unordered.
+By definition, the unordered collection of $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
+That is, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, no matter which value in $[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ the test point $Z_{n+j}$ takes on, the unordered set of $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ remains the same.
+As a result, $\widetilde{R}$ is also fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
+In addition, the set of scores $\{v_{i}\colon i\in\widetilde{R}\}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, which remains the same no matter which value in $[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ the test point $Z_{n+j}$ takes on.
+Recall that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})=\{i\in[n]:j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j})\in\mathfrak{S}\}$.
+We let ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})$ if $j\notin\widehat{\mathcal{S}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$ and ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}$ otherwise.
 We claim that given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
 $\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}=\{v_{i}\colon i\in\widetilde{R}\},$ (34)
-where repetition of elements is allowed on both sides. To see this, let $\pi$ denote the permutation on ${\mathcal{I}}_{j}$ such that $V_{i}=v_{\pi(i)}$ and $V_{n+j}=v_{\pi(n+j)}$. Then
+where repetition of elements is allowed on both sides.
+In words, (34) means that no matter which value in $[z_{1},\dots,z_{n},z_{n+j}]$ is the test point $Z_{n+j}$, the left-handed side always equals $\{v_{i}\colon i\in\widetilde{R}\}$, which is determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
+To see this, let $\pi$ denote the permutation on ${\mathcal{I}}_{j}$ such that $V_{i}=v_{\pi(i)}$ and $V_{n+j}=v_{\pi(n+j)}$.
+On this event,
 $\widehat{\mathcal{S}}={\mathcal{S}}\big{(}\underbrace{[z_{\pi(1)},z_{\pi(2)},\dots,z_{\pi(n)}]}_{\textrm{calibration data}},\underbrace{X_{n+1},\dots,x_{\pi(n+j)},\dots,X_{n+m}}_{\textrm{test data}}\big{)}={\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(n+1)}],D_{\textnormal{test}}^{\pi(n+1)}\big{)},$ (35)
@@
 $\qquad\qquad={\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}$
-for all $i\in\{1,\dots,n\}$. Denoting $\widehat{\mathcal{S}}=\widehat{\mathcal{S}}^{{\textnormal{swap}(n+j,j)}}$ for notational simplicity, we have,
-$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}$ =$\overset{\rm(a)}{=}\big{\{}V_{i}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\big{\}}$
-=$\overset{\rm(b)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}\big{\}}$
+for all $i\in\{1,\dots,n\}$. Denoting $\widehat{\mathcal{S}}=\widehat{\mathcal{S}}^{{\textnormal{swap}(n+j,j)}}(Y_{n+j})$ for notational simplicity, we have,
+$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}$ =$\overset{\rm(a)}{=}\big{\{}V_{i}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j})\in\mathfrak{S}\big{\}}$
+=$\overset{\rm(b)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(Y_{n+j})\in\mathfrak{S}\big{\}}$
 =$\overset{\rm(c)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\pi(i)}],D_{{\textnormal{test}}}^{\pi(i)}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}\in\mathfrak{S}\big{\}}$
-=$=\big{\{}v_{\ell}\colon\ell\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\ell}],D_{{\textnormal{test}}}^{\ell}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}\in\mathfrak{S}\big{\}}$
+=$=\big{\{}v_{\ell}\colon\ell\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\ell}],D_{{\textnormal{test}}}^{\ell}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\ell}],D_{\textnormal{test}}^{\ell}\big{)}\in\mathfrak{S}\big{\}}$
 =$\overset{\rm(d)}{=}\{v_{i}\colon i\in\widetilde{R}\}.$
-Above, step (a) follows from the definition of ^$\widehat{\mathcal{R}}_{n+j}^{+}$; steps (b) and (c) are due to (34), and step (d) follows from the definition of ${\widetilde{R}}$.
-We thus completed the proof of Eqn.
-(33).
-Returning to the proof of (32), we have by Bayes’ rule that
+Above, step (a) follows from the definition of ^$\widehat{\mathcal{R}}_{n+j}^{+}$; steps (b) and (c) are due to (LABEL:eq:representation), and step (d) follows from the definition of ${\widetilde{R}}$.
+We thus completed the proof of Equation (34).
+Returning to the proof of (33), we have that
 $\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S},\leavevmode\nobreak\ [d_{j}],{\mathcal{D}}_{j}^{c}\Big{)}$ (36)
@@
 […]
+where the last step is due to Bayes’ rule.
 Again, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, the only randomness is the assignment of $[z_{1},\dots,z_{n},z_{n+j}]$ to $(Z_{1},\dots,Z_{n},Z_{n+j})$.
-By definition and the exchangeability of ${\mathcal{D}}_{j}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$, it is straightforward that
-$\mathbb{P}\big{(}j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}$ $=\sum_{i\in{\mathcal{I}}_{j}}\mathbb{P}\big{(}Z_{n+j}=z_{i}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}\cdot\mathds{1}\{i\in{\widetilde{R}}\}$ (36)
-$=\sum_{i\in{\mathcal{I}}_{j}}\frac{\mathds{1}\{i\in\widetilde{R}\}}{n+1}=\frac{|\widetilde{R}|}{n+1}.$ (37)
-On the other hand, by (33),
+By (34),
 $\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\Big{(}1-\alpha;\{V_{i}\}_{i\in{\mathcal{R}}_{n+j}^{+}(Y_{n+j})}\Big{)},j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\Big{)}$
@@
 […]
-Since ${\widetilde{R}}$ is determistic conditional on $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
+Since ${\widetilde{R}}$ is deterministic conditional on $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we know that
 ($=\leavevmode\nobreak$ $\sum_{\ell\in{\widetilde{R}}}\mathbb{P}\big{(}V_{n+1}=v_{\ell}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}\cdot\mathds{1}\{v_{\ell}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{v_{i}\colon i\in\widetilde{R}\}\}$ (39)
@@
 (40)$\geq\frac{(1-\alpha)|\widetilde{R}|}{n+1}.$
-Combining the above and (37), we have thus verified (35). Then by the tower property, we conclude (32), and thus (11).
-When the elements in ${\widetilde{R}}$ are distinct, we additionally have the following determistic inequality following from the definition of the quantile function:
+On the other hand, it is straightforward that
+$\mathbb{P}\big{(}j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}$ $=\sum_{i\in{\mathcal{I}}_{j}}\mathbb{P}\big{(}Z_{n+j}=z_{i}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}\cdot\mathds{1}\{i\in{\widetilde{R}}\}$ (41)
+$=\sum_{i\in{\mathcal{I}}_{j}}\frac{\mathds{1}\{i\in\widetilde{R}\}}{n+1}=\frac{|\widetilde{R}|}{n+1}.$ (42)
+Combining the above and (42), we have thus verified (LABEL:eq:equiv_cond_cov3). Then by the tower property, we conclude (33), and thus (11).
+When the elements in ${\widetilde{R}}$ are distinct, we additionally have the following deterministic inequality following from the definition of the quantile function:
 (40)$\leq\frac{(1-\alpha)|{\widetilde{R}}|+1}{n+1},$
@@
 $\mathbb{P}\big{(}Z_{n+j}=z_{[i]}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{1}{n+1}.$
-Following the proof of part (a) up to Equation (33), we know that
+Following the proof of part (a), we know that
 $\mathbb{P}\big{(}j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}=\frac{|\widetilde{R}|}{n+1}.$
-On the other hand, due to (33) and the fact that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}=\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})$. when $j\in\widehat{\mathcal{S}}$, for any $t\in(0,1)$,
-$\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{i}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{i}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
+On the other hand, due to (34) we have for any $t\in(0,1)$,
+$\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}<V_{i}\}+U_{j}(1+\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}=V_{i}\})}{1+|\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
+$=$ $\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}<V_{i}\}+U_{j}(\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}^{+}(Y_{n+j})}\mathds{1}\{V_{n+j}=V_{i}\})}{|\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
 $=$ $\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{[i]}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{[i]}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
@@
 where we used the take-out property of conditional probability/expectation in the first step; the probability in the second line is only with respect to $U_{j}\sim\textnormal{Unif}[0,1]$, and the last equality is a standard result that is straightforward to check with exchangeability (see, e.g., Vovk et al., (2005, Proposition 2.4)).
-Putting them togther and letting $t=1-\alpha$, we have
+Putting them together and letting $t=1-\alpha$, we have
 $\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{\textnormal{rand}}\,\big{|}\,j\in\widehat{\mathcal{S}},[d_{j}],{\mathcal{D}}_{d}^{c}\big{)}=\frac{\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j},\leavevmode\nobreak\ j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}}{\mathbb{P}\big{(}j\in\widehat{\mathcal{S}}\,\big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\big{)}}=1-\alpha,$
 which concludes the proof of statement (b).
-B.5 Proof of Proposition 4
+D.5 Proof of Proposition 4
+Fix $j\in\widehat{{\mathcal{S}}}$.
+The proof proceeds as follows: we consider the cases of $y\leq c_{n+j}$ and $y>c_{n+j}$ separately.
+In each case, we show ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$ by proving that ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$ and ^$\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$.
+For notational simplicity, let $A^{\textnormal{swap}(i,j)}_{s}(y)$ and $B^{\textnormal{swap}(i,j)}_{s}(y)$ denote the partial sums resulting from swapping units $i$ and $n+j$ in $A_{s}$ and $B_{s}$, and $\tau^{\textnormal{swap}(i,j)}(y)$ is the corresponding stopping time.
+D.5.1 Case 1: $y\leq c_{n+j}$.
+Showing ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$.
+Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+We now show that $i\in\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$.
+By definition, we have $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)$ and
+$\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}(y)\geq\tau^{\textnormal{swap}(i,j)}\}\in\mathfrak{S}.$
+  1. (a)
+    Suppose $Y_{i}\leq c_{i}$. By the definition of $\tau^{(j)}(1,0)$, we have that
+    $\mathds{1}\{\tau^{(j)}(1,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(1,0)\}_{s\leq\tau^{\textnormal{swap}(i,j)}},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}}).$
+    Since $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)$, for any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, $\widehat{S}_{i}\geq s$ and
+    $A_{s}^{(j)}(1,0)$ $=\mathds{1}\{\widehat{S}_{n+j}\geq s\}+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+1=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    As a result,
+    $\begin{split}\mathds{1}\{\tau^{(j)}(1,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}&=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(1,0)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)})\\ &=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{\textnormal{swap}(i,j)}},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{\textnormal{swap}(i,j)}})\\ &=\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1,\end{split}$ (43)
+    where we use the definition of $\tau^{\textnormal{swap}(i,j)}(y)$.
+Above, we have shown $\tau^{(j)}(1,0)\leq\tau^{\textnormal{swap}(i,j)}(y)$, and so that $\widehat{S}_{i}\geq\tau^{(j)}(1,0)$.
+Conversely, by the definition of $\tau^{\textnormal{swap}(i,j)}(y)$, there is
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,0)\}=f_{\tau^{(j)}(1,0)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,0)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,0)}).$
+    Now, observe that for any $s\leq\tau^{(j)}(1,0)$,
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+1$
+    $=\mathds{1}\{\widehat{S}_{n+j}\geq s\}+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}=A_{s}^{(j)}(1,0),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Using an argument similar to (43), we can draw the conclusion that $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,0)\}=1$.
+Combining the two results, we have that $\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(1,0)$.
+Therefore, $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(1,0)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}(y)\geq\tau^{\textnormal{swap}(i,j)}\}\in\mathfrak{S}.$
+    That is, $i\in\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}$(y).
+  2. (b)
+    Suppose that $Y_{i}>c_{i}$. Then by the definition of $\tau^{(j)}(1,1)$, we have that
+    $\mathds{1}\{\tau^{(j)}(1,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(1,1)\}_{s\leq\tau^{\textnormal{swap}(i,j)}},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}}).$
+    For any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$,
+    $A_{s}^{(j)}(1,1)$ $=1+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    As before, we can conclude that $\tau^{(j)}(1,1)\leq\tau^{\textnormal{swap}(i,j)}(y)$, and therefore $\widehat{S}_{i}\geq\tau^{(j)}(1,1)$.
+Conversely, by the definition of $\tau^{\textnormal{swap}(i,j)}(y)$, there is
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,1)\}=f_{\tau^{(j)}(1,1)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,1)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,1)}).$
+    For any $s\leq\tau^{(j)}(1,1)$,
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=1+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s\}=A^{(j)}(1,1),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Then $\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,1)$, and therefore $\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(1,1)$.
+Consequently, $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(1,1)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(1,1)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}\in\mathfrak{S}.$
+Combining (a) and (b), we have proved that ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}^{\textnormal{cp}}(y)$.
+Showing ^$\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$.
+Now suppose $i\in\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$. We shall show that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+  1. (c)
+[…]
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,0)\}$ $=f_{\tau^{(j)}(1,0)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,0)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,0)}).$
+    For any $s\leq\tau^{(j)}(1,0)$, we have
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s\}=A_{s}^{(j)}(1,0),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Then as in (43), we have $\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,0)$.
+We also have that $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)$.
+Conversely, using the definition of $\tau^{(j)}(1,0)$, we can write that
+    $\mathds{1}\{\tau^{(j)}(1,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(1,0)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}).$
+    For any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, we can check that
+    $A_{s}^{(j)}(1,0)$ $=\mathds{1}\{\widehat{S}_{n+j}\geq s\}+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+1=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    Then $\mathds{1}\{\tau^{(j)}(1,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$.
+Collectively, we arrive at $\tau^{(j)}(1,0)=\tau^{\textnormal{swap}(i,j)}(y)$, implying that $\widehat{S}_{i}\geq\tau^{(j)}(1,0)=\tau^{\textnormal{swap}(i,j)}(y)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(1,0)\}\in\mathfrak{S},$
+    which implies that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+  2. (d)
+    Suppose $Y_{i}>c_{i}$. In this case, we have $\widehat{S}_{i}\geq\tau^{(j)}(1,1)$. The definition of $\tau^{\textnormal{swap}(i,j)}(y)$ leads to
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,1)\}=f_{\tau^{(j)}(1,1)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,1)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(1,1)}).$
+    For any $s\leq\tau^{(j)}(1,1)$, we have
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=1+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s\}=A_{s}^{(j)}(1,1),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Then we have $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(1,1)\}=1$, and therefore $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)$.
+Conversely, by the definition of $\tau^{(j)}(1,1)$, we can write that
+    $\mathds{1}\{\tau^{(j)}(1,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(1,1)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}).$
+    For any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, there is
+    $A_{s}^{(j)}(1,1)$ $=1+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    Then we have $\mathds{1}\{\tau^{(j)}(1,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$.
+Combining the two results, we have that $\tau^{(j)}(1,1)=\tau^{\textnormal{swap}(i,j)}(y)$, and therefore $\widehat{S}_{i}\geq\tau^{(j)}(1,1)=\tau^{\textnormal{swap}(i,j)}(y)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(1,1)\}\in\mathfrak{S},$
+    which implies that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+Combining (c) and (d), we have proved that ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$, which finishes the proof for case 1.
+D.5.2 Case 2: $y>c_{n+j}$.
+As in case 1, we prove that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$ by showing that ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$.
+Showing ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$.
+Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. We proceed to show that $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$.
+  1. (a)
+    Suppose $Y_{i}\leq c_{i}$. Using the definition of $\tau^{(j)}(0,0)$, we write
+    $\mathds{1}\{\tau^{(j)}(0,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(0,0)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}).$
+    We have for any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$ that
+    $A_{s}^{(j)}(0,0)$ $=\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    We then have $\mathds{1}\{\tau^{(j)}(0,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$.
+Similarly, we can use the definition of $\tau^{\textnormal{swap}(i,j)}(y)$ and obtain that
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,0)\}=f_{\tau^{(j)}(0,0)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,0)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,0)}).$
+    Again for any $s\leq\tau^{(j)}(0,0)$, we have
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}=A_{s}^{(j)}(0,0),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Then we have $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,0)\}=\mathds{1}\{\tau^{(j)}(0,0)\leq\tau^{(j)}(0,0)\}=1$.
+Combining the two results, we have that $\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(0,0)$, and therefore $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(0,0)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(0,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}\in\mathfrak{S},$
+    which implies that $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$.
+  2. (b)
+    Suppose $Y_{i}>c_{i}$. We use the definition of $\tau^{(j)}(0,1)$:
+    $\mathds{1}\{\tau^{(j)}(0,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(0,1)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}).$
+    For any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, we have
+    $A_{s}^{(j)}(0,1)$ $=1+\sum_{k\in[n]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    Then we have $\mathds{1}\{\tau^{(j)}(0,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$.
+Similarly, we can use the definition of $\tau^{\textnormal{swap}(i,j)}(y)$ and obtain that
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,1)\}=f_{\tau^{(j)}(0,1)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,1)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,1)}).$
+    For any $s\leq\tau^{(j)}(0,1)$, we have $\widehat{S}_{i}\geq s$ and
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=1+\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}=A_{s}^{(j)}(0,1),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Then we have $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,1)\}=1$, and therefore $\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(0,1)$.
+This implies that $\widehat{S}_{i}\geq\tau^{\textnormal{swap}(i,j)}(y)=\tau^{(j)}(0,1)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(0,1)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}\in\mathfrak{S}.$
+    Then we have $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$.
+Combining (a) and (b), we have proved that ^$\widehat{{\mathcal{R}}}_{n+j}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)$ when $y>c_{n+j}$. We proceed to show the reverse inclusion.
+Showing ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$.
+Suppose that $i\in\widehat{{\mathcal{R}}}^{\textnormal{cp}}_{n+j}(y)$.
+  1. (c)
+    When $Y_{i}\leq c_{i}$, $\widehat{S}_{i}\geq\tau^{(j)}(0,0)$. As before, we have
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,0)\}=f_{\tau^{(j)}(0,0)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,0)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,0)}).$
+    For any $s\leq\tau^{(j)}(0,0)$, observe that
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\ \geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\ \geq s,Y_{k}\leq c_{k}\}=A_{s}^{(j)}(0,0),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)},$
+    which leads to $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,0)\}=1$. Conversely, we have
+    $\mathds{1}\{\tau^{(j)}(0,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{{\textnormal{swap}(i,j)}}(y)}(\{A_{s}^{(j)}(0,0)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}),$
+    and for any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, we have
+    $A_{s}^{(j)}(0,0)$ $=\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\ \geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    Then we have $\mathds{1}\{\tau^{(j)}(0,0)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$, which implies that $\tau^{(j)}(0,0)=\tau^{\textnormal{swap}(i,j)}(y)$.
+We then have $\widehat{S}_{i}\geq\tau^{(j)}(0,0)=\tau^{\textnormal{swap}(i,j)}(y)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(0,0)\}\in\mathfrak{S},$
+    implying that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+  2. (d)
+    When $Y_{i}>c_{i}$, $\widehat{S}_{i}\geq\tau^{(j)}(0,1)$. We have
+    $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,1)\}=f_{\tau^{(j)}(0,1)}(\{A_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,1)},\{B_{s}^{\textnormal{swap}(i,j)}(y)\}_{s\leq\tau^{(j)}(0,1)}).$
+    For any $s\leq\tau^{(j)}(0,1)$, there is
+    $A_{s}^{\textnormal{swap}(i,j)}(y)$ $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\ \geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}$
+    $=1+\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\ \geq s,Y_{k}\leq c_{k}\}=A_{s}^{(j)}(0,1),$
+    $B_{s}^{\textnormal{swap}(i,j)}(y)$ $=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{(j)}.$
+    Therefore, we have $\mathds{1}\{\tau^{\textnormal{swap}(i,j)}(y)\leq\tau^{(j)}(0,1)\}=1$. Conversely, we have
+    $\mathds{1}\{\tau^{(j)}(0,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=f_{\tau^{\textnormal{swap}(i,j)}(y)}(\{A_{s}^{(j)}(0,1)\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)},\{B_{s}^{(j)}\}_{s\leq\tau^{\textnormal{swap}(i,j)}(y)}),$
+    where for any $s\leq\tau^{\textnormal{swap}(i,j)}(y)$, we have
+    $A_{s}^{(j)}(0,1)$ $=1+\sum_{k\in[m]}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}$
+    $=1+\sum_{k\neq i}\mathds{1}\{\widehat{S}_{k}\geq s,Y_{k}\leq c_{k}\}+\mathds{1}\{\widehat{S}_{n+j}\geq s,y\leq c_{n+j}\}=A_{s}^{\textnormal{swap}(i,j)}(y),$
+    $B_{s}^{(j)}$ $=1+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=\mathds{1}\{\widehat{S}_{i}\geq s\}+\sum_{\ell\neq j}\mathds{1}\{\widehat{S}_{n+\ell}\geq s\}=B_{s}^{\textnormal{swap}(i,j)}(y).$
+    Then we have $\mathds{1}\{\tau^{(j)}(0,1)\leq\tau^{\textnormal{swap}(i,j)}(y)\}=1$, which implies that $\tau^{(j)}(0,1)=\tau^{\textnormal{swap}(i,j)}(y)$.
+Therefore, $\widehat{S}_{i}\geq\tau^{(j)}(0,1)=\tau^{\textnormal{swap}(i,j)}(y)$ and
+    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq\tau^{(j)}(0,1)\}\in\mathfrak{S},$
+    implying that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
+Combining (c) and (d), we have proved that ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{cp}}(y)\subseteq\widehat{{\mathcal{R}}}_{n+j}(y)$ when $y>c_{n+j}$. The whole proof is now complete.
+D.6 Proof of Proposition 5
+Recall that the generic form of the JOMI prediction set is
+$\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(y)}\cup\{\infty\},1-\alpha\big{)}\Big{\}},$
+with ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ being the reference set.
+We categorize $y$ into three cases according to the value of $S(X_{n+j},y)$: (1) $S(X_{n+j},y)<\eta^{-}$, (2) $S(X_{n+j},y)>\eta^{+}$, and (3) $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$.
+The case where $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$ is obvious, so we discuss the remaining two cases in the following.
+In what follows, we let $\eta^{\textnormal{swap}(i,j)}(y)$ denote the $K$-th smallest elements in $\{S^{\textnormal{swap}(i,j)}_{\ell}(y)\}_{\ell\in[n]}$.
+Case 1: $S(X_{n+j},y)<\eta^{-}$.
+Consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)<\eta^{-}$ and $y\in\widehat{C}_{\alpha,n+j}$. We claim that in this case
+^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}\,:=$$\,\{i\in[n]:S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1,\{\ell\in[m]:{\mathcal{L}}(X^{\textnormal{swap}(i,j)}_{n+\ell},\eta)=1\}\in\mathfrak{S}\}$
+$\cup\{i\in[n]:S_{i}>\eta^{-},{\mathcal{L}}(X_{i},\eta^{-})=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{-})=1\}\in\mathfrak{S}\}.$
+To see this, note that when $S_{i}\leq\eta^{-}$, we have $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$. Then ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta)$ and
+$\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta)=1\}=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{{\textnormal{swap}(i,j)}}(y))=1\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
+When $S_{i}>\eta^{-}$, $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{-}$, ${\mathcal{L}}(X_{i},\eta^{-})={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$ and
+$\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{-})=1\}=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{{\textnormal{swap}(i,j)}}(y))=1\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
+As a result, we have that ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}$ and therefore
+$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}:i\in{\widetilde{R}}_{n+j}^{(1)}\}\cup\{\infty\}\big{)}$
+Case 2: $S(X_{n+j},y)>\eta^{+}$.
+We now consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)>\eta^{+}$ and $y\in\widehat{C}_{\alpha,n+j}$. In this case, we instead have
+^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(2)}\,:=$$\,\{i\in[n]:S_{i}\leq\eta,{\mathcal{L}}(X_{i},\eta^{+})=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{+})=1\}\in\mathfrak{S}\}$
+$\cup\{i\in[n]:S_{i}>\eta,{\mathcal{L}}(X_{i},\eta)=1,\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta)=1\}\in\mathfrak{S}\}.$
+This is because when $S_{i}\leq\eta$, $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{+}$ and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta^{+})$, as well as
+$\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{{\textnormal{swap}(i,j)}}(y))=1\}=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{+})=1\}.$
+When $S_{i}>\eta$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$. Therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta)$, and
+$\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta^{{\textnormal{swap}(i,j)}}(y))=1\}=\{\ell\in[m]:{\mathcal{L}}(X_{n+\ell}^{\textnormal{swap}(i,j)},\eta)=1\}.$
+Consequently,
+$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}:i\in{\widetilde{R}}_{n+j}^{(2)}\}\cup\{\infty\}\big{)}.$
+With the above, we have completed the proof of Proposition 5.
+D.7 Proof of Proposition 6
 We start by defining an alternative stopping time that is determined solely by ${\mathcal{D}}_{j}^{c}$:
-${\widetilde{T}}_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\geq m-K\Big{\}}.$
-Abusing the notation a bit, we call $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$ and ${\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ the corresponding stopping times defined with $({\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}$, ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)})$.
-Since ${\widetilde{T}}_{\textnormal{topk}}$ is agnostic to the swapping of unit $i$ and $n+j$, ${\widetilde{T}}^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
-We claim that when $S_{n+j}>T_{\textnormal{topk}}$, $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, To see this, note first that for any $t\in{\mathcal{T}}$,
-$\sum_{i\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{S_{i}\leq t\}\geq\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ {\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}.$
+${\widetilde{T}}_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq t\}\geq m-K\Big{\}}.$
+Abusing the notations a bit, we call $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$ and ${\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ the corresponding stopping times defined with $({\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}$, ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)})$.
+By definition, it is straightfoward to check that ${\widetilde{T}}^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
+We claim that $S_{n+j}>T_{\textnormal{topk}}$ implies $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$. To see this, note first that for any $t\in{\mathcal{T}}$,
+$\sum_{\ell\in[m]}\mathds{1}\{S_{n+\ell}\leq t\}\geq\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq t\}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ {\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}.$
 On the other hand, one can check that
-$\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}\overset{\rm(a)}{=}\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}+\mathds{1}\{S_{n+j}\leq T_{\textnormal{topk}}\}\overset{\rm(b)}{\geq}m-K.$ (42)
+$\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}\}\overset{\rm(a)}{=}\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}\}+\mathds{1}\{S_{n+j}\leq T_{\textnormal{topk}}\}\overset{\rm(b)}{\geq}m-K.$ (44)
 Above, step (a) is because $S_{n+j}>T_{\textnormal{topk}}$ and step (b) follows from the choice of $T_{\textnormal{topk}}$.
 With the definition of ${\widetilde{T}}_{\textnormal{topk}}$, we can then conclude from (44) that $T_{\textnormal{topk}}\geq{\widetilde{T}}_{\textnormal{topk}}$, and hence $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
-Similarly, we have ${\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ by construction. When $S_{i}>T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$,
-$\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}=\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}+\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}\geq m-K.$
-By the definition of $T_{\textnormal{topk}}$, we then have $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\geq T_{\textnormal{topk}}$, and therefore $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}=T_{\textnormal{topk}}$.
-Fix $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$, we proceed to show that ^$\widehat{{\mathcal{R}}}_{\textnormal{fixed}}=\widehat{{\mathcal{R}}}_{n+j}$.
-For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{topk}}$, there is $S_{i}>T_{\textnormal{topk}}$.
-We then have $S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$, since $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}\geq T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$.
+We can similarly show that $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$ when $S_{i}>T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$.
+By construction, we have ${\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$.
+Next,
+$\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}=\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}+\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}\geq m-K.$
+By the definition of ${\widetilde{T}}_{\textnormal{topk}}$, we then have $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\geq{\widetilde{T}}_{\textnormal{topk}}$, and therefore $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}={\widetilde{T}}_{\textnormal{topk}}$.
+We now proceed to show that ^$\widehat{{\mathcal{R}}}_{\textnormal{fixed}}=\widehat{{\mathcal{R}}}_{n+j}$.
+For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{topk}}$, it holds that $S_{i}>T_{\textnormal{topk}}$.
+Then since $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}\geq T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$, we know that $S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$.
 Therefore, $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, and
@@
 wher step (a) follows from the fact that $S_{n+j}^{\textnormal{swap}(i,j)}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{topk}}$.
-As a result, $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$.
-On the other hand, for any $i\in\widehat{{\mathcal{R}}}_{n+j}$, we have that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$.
-Then there is
-$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{topk}}\Rightarrow S_{i}>T_{\textnormal{topk}}.$ (43)
+As a result, $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$, i.e., $i\in\widehat{{\mathcal{R}}}_{n+j}$.
+On the other hand, for any $i\in\widehat{{\mathcal{R}}}_{n+j}$, it holds that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$.
+Then,
+$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}=T_{\textnormal{topk}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ S_{i}>T_{\textnormal{topk}}.$ (45)
 Combining all the above, we complete the proof.
-B.6 Proof of Proposition 5
-Fix $j\in\widehat{{\mathcal{S}}}$. Define the proxy stopping time that is agnostic to the swapping of units $i$ and $n+j$:
+D.8 Proof of Proposition 7
+Fix $j\in\widehat{{\mathcal{S}}}_{\textnormal{cq}}$. Define the proxy stopping time that is agnostic to the swapping of units $i$ and $n+j$:
 ${\widetilde{T}}_{\textnormal{cq}}=\inf\bigg{\{}t\in{\mathcal{T}}:\sum_{i\in{\mathcal{I}}_{j}}\mathds{1}\{S_{i}\leq t\}\geq qn\bigg{\}}.$
@@
 To see this, note that
-$\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$ (44)
-where the last inequality follows from the definition of ${\widetilde{T}}_{\textnormal{cq}}$.
+$\sum_{i\in[n]}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{i\in[n]}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$ (46)
+where the first step is because $S_{n+j}>{\widetilde{T}}_{\textnormal{cq}}$ and the second step follows from the definition of ${\widetilde{T}}_{\textnormal{cq}}$.
 By the definition of $T_{\textnormal{cq}}$, (46) implies that ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}$ and therefore $T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
-Similarly, we have ${\widetilde{T}}_{\textnormal{cq}}\leq T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}$ by construction. When $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$,
-$\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$
-with the last step following from the definition of ${\widetilde{T}}_{\textnormal{cq}}$.
+Similarly, we claim that $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$ when $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$.
+By construction, we have ${\widetilde{T}}_{\textnormal{cq}}\leq T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}$.
+When $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$,
+$\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{\ell\in[n]}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$
+with the first step following from $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$ and the second step from the definition of ${\widetilde{T}}_{\textnormal{cq}}$.
 We then have ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}^{\textnormal{swap}(i,j)}$ and therefore $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
@@
 $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ =$=\{\ell\in[m]:S_{n+\ell}^{\textnormal{swap}(i,j)}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\}{=}\{\ell\in[m]:S_{n+\ell}^{\textnormal{swap}(i,j)}>{\widetilde{T}}_{\textnormal{cq}}\}$
-=$\overset{\rm(a)}{=}\{j\}\cup\{\ell\in[m]\backslash\{j\}:S_{n+\ell}>{\widetilde{T}}_{\textnormal{cq}}\}=\widehat{{\mathcal{S}}}_{\textnormal{cq}},$
+=$\overset{\rm(a)}{=}\{j\}\cup\{\ell\in[m]\backslash\{j\}:S_{n+\ell}>T_{\textnormal{cq}}\}=\widehat{{\mathcal{S}}}_{\textnormal{cq}},$
 where step (a) follows from the fact that $S^{\textnormal{swap}(i,j)}_{n+j}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$.
-The above implies that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and and we have $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$.
-Reversely, for $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$,
-$S_{n+j}^{\textnormal{swap}(i,j)}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{cq}}\Rightarrow S_{i}>T_{\textnormal{cq}},$
-completing the proof.
-B.7 Proof of Proposition 6
-To start, we claim that the $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ is equivalent to
-${\widetilde{S}}_{\textnormal{fixed}}=\{j\in[m]:\widehat{{\mathcal{S}}}_{n+j}\geq T\},\text{ where }T=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}\geq q(n+1)\bigg{\}}.$
-To see this, suppose $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$, then $p_{j}\leq q$. We can check
-$\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}=p_{j}\leq q.$
-By the definition of $T$, $\widehat{S}_{n+j}\geq T$. Conversely, if $\widehat{S}_{n+j}\geq T$, then
-$p_{j}=\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}\leq\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T,Y_{i}\leq c_{i}\}\Big{)}\leq q\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}.$
-We then have $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}={\widetilde{S}}_{\textnormal{fixed}}$. We shall work with ${\widetilde{S}}_{\textnormal{fixed}}$ in the following.
-Fix $j\in{\widetilde{S}}_{\textnormal{fixed}}$ and a selection taxonamy $\mathfrak{S}$ such that ${\widetilde{S}}_{\textnormal{fixed}}\in\mathfrak{S}$.
-We prove the proposition for $y\leq c_{n+j}$ and $y>c_{n+j}$ separately.
-For notational simplicity, we define the swapped stopping time for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ and $y\in{\mathcal{Y}}$ as
-$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
-Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$
-Case 1: $y\leq c_{n+j}$.
-The goal is to show that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$. We prove this for $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
-  • $Y_{i}\leq c_{i}$.
-    Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$, i.e., $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
-    ${1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i},y\leq c_{n+j}\}}\leq q(n+1).$ (45)
-    Since $y\leq c_{n+j}$, the above inequality is equivalent to
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}\leq q(n+1).$ (46)
-    Since $\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{i},Y_{i}\leq c_{i}\}=1$, we further have
-    ${\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}}\leq q(n+1).$
-    The above and the definition of $T_{j}^{\textnormal{fixed}}(1,0)$ implies that $\widehat{{\mathcal{S}}}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)$. Next, note that
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{{\mathcal{S}}}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$
+The above implies that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and that $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$.
+Conversely, for $i\in[n]$ such that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$,
+$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}=T_{\textnormal{cq}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ S_{i}>T_{\textnormal{cq}},$
+which completes the proof.
+Appendix E Auxiliary results
+E.1 Connection between selection-conditional coverage and FCR control
+Lemma 1 establishes the asympototic equivalence between the selection-conditional coverage and the false coverage rate (FCR) when the selection rule is asymptotically close to a threshold function.
+Lemma 1.
+Suppose $\{Z_{i}\}_{i=1}^{m+n}$ are i.i.d., and there exists some fixed function $f:{\mathcal{X}}\mapsto\mathbb{R}$ and fixed constant $\tau\in\mathbb{R}$ such that the selection set $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ obeys $|\widehat{\mathcal{S}}\Delta\widehat{\mathcal{S}}^{*}|/|\widehat{\mathcal{S}}|{\to}0$ in probability as $n,m\rightarrow\infty$, where $\widehat{\mathcal{S}}^{*}=\{j\colon f(X_{n+j})\geq\tau\}$ and $\mathbb{P}(f(X_{n+1})\geq\tau)>0$.
+Assume that $\widehat{S}$ is equivariant to the permutation of the units in ${\mathcal{D}}_{\textnormal{test}}$ and that the prediction set takes the form $\widehat{C}_{\alpha,n+j}=\{y\colon V(X_{n+j},y)\leq\widehat{q}_{m,n}\}$, where $\widehat{q}_{m,n}\stackrel{{\scriptstyle\textnormal{p}}}{{\to}}q^{*}\in\mathbb{R}$, as $m,n\to\infty$.
+Assume also that the function $G(t):=\mathbb{P}(f(X_{n+1})\geq\tau,V_{n+1}>t)$ is continuous in $t$.
+Then,
 […]
-    where the first step is because $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$; the second step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$.
-Eqn.
-(B.7) implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{{\textnormal{swap}(i,j)}}(y)$.
-Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{{\textnormal{swap}(i,j)}}(y),Y_{i}\leq c_{i}\}=1$, we have
-    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}$
-[…]
-    following the definition of $T^{\textnormal{swap}(i,j)}(y)$.
-We then conclude that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{fixed}}_{j}(1,0)$ and therefore $T_{j}^{(1,0)}=T^{{\textnormal{swap}(i,j)}}(y)$.
-As a result,
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}.$
-    We have therefore verified that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
-    Next, given $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, our goal now is to show that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. By definition, there is
-    $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)\text{ and }\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
-    Since $y\leq c_{n+j}$ and $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,0),Y_{i}\leq c_{i}\}=1$,
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{{\mathcal{S}}}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$
-[…]
-    where the last step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$.
-The above implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$, and therefore $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$.
-Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$, and $y\leq c_{n+j}$,
-    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}$
+That is, the asymptotic FCR coincides with the asymptotic selection-conditional miscoverage.
+Proof.
+By the definition of FCR, we have
+FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
+$=\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
+$\qquad-\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*}\setminus\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}+\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}\setminus\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}.$ (47)
+Note that
+$\Bigg{|}-\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*}\setminus\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\Bigg{]}+\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}\setminus\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}\bigg{|}$
 […]
 […]
-    following the definition of $T^{\textnormal{swap}(i,j)}(y)$.
-As a result, $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(1,0)\Rightarrow T_{j}^{\textnormal{fixed}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$.
-The swapped selection set
-    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)$ =$=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
-    It follows that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-  • $Y_{i}>c_{i}$.
-    Note that for any $t\in{\mathcal{T}}$,
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t\}$
-[…]
-    becuase $y\leq c_{n+j}$ and $Y_{i}>c_{i}$.
-Inspecting the definition of $T_{j}^{\textnormal{fixed}}(1,1)$ and $T^{\textnormal{swap}(i,j)}(y)$, we can conclude that $T_{j}^{\textnormal{fixed}}(1,1)=T^{\textnormal{swap}(i,j)}(y)$.
-As a result, we have $\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
-    $\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,1)\}=\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}.$
-    Therefore $i\in\widehat{{\mathcal{R}}}_{n+y}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed}}(y)$.
-Combing the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$ for $y\leq c_{n+j}$.
-Case 2: $y>c_{n+j}$.
-We again consider the case of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$ separately.
-  • $Y_{i}\leq c_{i}$.
-    Fix $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-By definition, there is $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$.
-Since $y>c_{n+j}$,
-    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$
+which goes to zero by assumption and the dominated convergence theorem. For the first term in Equation (E.1), we further have
+$\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
 […]
 […]
-    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$.
-We thus have $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$.
-Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$,
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
+The absolute value of the second term above is bounded by $\mathbb{E}[|\widehat{\mathcal{S}}^{*}\Delta\widehat{\mathcal{S}}|/|\widehat{\mathcal{S}}|]$, and therefore goes to zero by assumption.
+Next, by the strong law of large numbers,
 […]
-    by the definition of $T_{j}^{\textnormal{fixed}}(0,0)$. As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)=T^{\textnormal{swap}(i,j)}(y)$, and
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(0,0)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
-    It then follows that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
-    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$ Note that since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
+Meanwhile,
+$\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}-G(\widehat{q}_{m,n})\bigg{|}$ $\leq\sup_{t\in\mathbb{R}}\leavevmode\nobreak\ \bigg{|}\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>t\}-G(t)\bigg{|}$
 […]
-    following the definition of $T_{j}^{\textnormal{fixed}}(0,0)$.
-As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$.
-Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$,
-    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$
+The last step follows from the uniform law of large numbers. Since $G(t)$ is continuous in $t$ and that $\widehat{q}_{m,n}\stackrel{{\scriptstyle p}}{{\to}}q^{*}$, there is
+$\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}\stackrel{{\scriptstyle\text{p}}}{{\to}}G(q^{*}).$ (48)
+Then, applying the continuous mapping theorem and the dominated convergence theorem yields
+$\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}^{*}|}\bigg{]}\rightarrow\mathbb{P}\big{(}V(X_{n+1},Y_{n+1})>q^{*}{\,|\,}f(X_{n+1})\geq\tau\big{)}.$
+This establishes the expression of the asymptotic FCR.
+In addition, due to the exchangeability of $\{Z_{n+j}\}_{j=1}^{m}$ and that $\widehat{S}$ is equivariant to the permutation of the test points, we have
 […]
-    by the choice of $T^{\textnormal{swap}(i,j)}(y)$.
-The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)\Rightarrow T^{\textnormal{swap}(i,j)}=T_{j}^{\textnormal{fixed}}(0,0)$.
-Consequently, we have
-    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$
-    We have thus shown that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-  • $Y_{i}>c_{i}$.
-    Since $y>c_{n+j}$ and $Y_{i}>c_{i}$, for any $t\in{\mathcal{T}}$ there is
-    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}=1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}.$
-    implying that $T_{j}^{\textnormal{fixed}}(0,1)=T^{\textnormal{swap}(i,j)}(y)$.
-Consequently, $\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
-    Then there is $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-Combining the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}=\widehat{{\mathcal{R}}}_{n+j}(y)$, and therefore ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$, for any $y>c_{n+j}$.
-The proof is now complete.
-B.8 Proof of Proposition 7
-We start from making the observation that $\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\in[m]:\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$, where
-$T^{\textnormal{BH}}=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq t\})\vee 1}\frac{m}{1+n}\leq q\bigg{\}}.$
-To see this, let $j^{*}=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$ and $\widehat{S}_{n+(1)},\widehat{S}_{n+(2)},\ldots,\widehat{S}_{n+(m)}$ denote the order statistics of $\{\widehat{S}_{n+j}\}_{j\in[m]}$ in an decending order.
-It then suffices to show that $j^{*}=k^{*}$.
-When $k^{*}\geq 1$,
-$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq\widehat{S}_{n+(k^{*})},Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq\widehat{S}_{n+(k^{*})}\})\vee 1}\frac{m}{1+n}=p_{(k^{*})}\frac{m}{k^{*}}\leq q,$ (48)
-where the last inequality follows from the definition of $k^{*}$.
-From the definition of $T^{\textnormal{BH}}$, we can conclude that $\widehat{S}_{n+(k^{*})}\geq T^{\textnormal{BH}}$; since $\widehat{S}_{n+(j^{*})}$ is the smallest element in $\{\widehat{S}_{n+j}\}_{j\in[m]}$ that is no less than $T^{\textnormal{BH}}$, we also have that $\widehat{S}_{n+(k^{*})}\geq\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$.
-Conversely,
-$p_{(j^{*})}\frac{m}{j^{*}}$ =$\overset{\rm(a)}{=}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+(j^{*})},Y_{i}\leq c_{i}\}}{n+1}\cdot\frac{m}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}$
-$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}},Y_{i}\leq c_{i}\}}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}\cdot\frac{m}{n+1}\overset{\rm(c)}{\leq}q.$
-Above, step (a) uses the definition of $j^{*}$, step (b) is because $\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$ and step (c) is by the definition of $T^{\textnormal{BH}}$.
-From the form of the BH procedure, we have that $j^{*}\geq k^{*}$, and therefore $j^{*}=k^{*}$.
-When $k^{*}=0$, for any $t\in{\mathcal{T}}$, we let $j(t)=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\}$. If $j(t)>0$, then
-$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{n+1}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{j(t)}\frac{m}{n+1}$
-[…]
-where the last step is because $k^{*}=0$. The above implies that $T^{\textnormal{BH}}>\widehat{S}_{n+(1)}\Rightarrow j^{*}=0$. The equivalence is therefore established.
-As before, we define for any $i\in[n]$, $j\in[m]$, and $y\in{\mathcal{Y}}$ the swapped stopping time:
-$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
-Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$.
-We are now ready to prove the proposition. We consider the two cases of $y\leq c_{n+j}$ and $y>c_{n+j}$ separately.
-Case 1: $y\leq c_{n+j}$.
-We discuss the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
-  • $Y_{i}\leq c_{i}$.
-Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$.
-Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),T_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$,
-    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{1+n}$
-[…]
-[…]
-    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$.
-By the choice of $T^{\textnormal{BH}}_{j}(1,0)$, there is $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(1,0)$.
-As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$.
-Conversely,
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
-[…]
-    following from the definition of $T^{\textnormal{BH}}_{j}(1,0)$.
-We therefore have $T^{\textnormal{BH}}_{j}(1,1)\geq T^{\textnormal{swap}(i,j)}(y)$, and then $T^{\textnormal{BH}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$.
-The above also implies that
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
-    By definition, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
-    On the other hand, suppose $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH}}(y)$, we have $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}$.
-We can check that
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
-[…]
-    by the definition of $T^{\textnormal{BH}}_{j}(1,0)$; the definition of $T^{\textnormal{swap}(i,j)}(y)$ implies that $T^{\textnormal{BH}}_{j}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$.
-Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$, there is
-    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
-[…]
-    and therefore $T^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\Rightarrow T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(1,0)$.
-We then have that $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
-    ^$\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}.$
-    That is, $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-We have thus shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}\leq c_{i}$.
-  • If $Y_{i}>c_{i}$, then for any $t\in{\mathcal{T}}$,
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
-[…]
-    By their definition, we have $T^{\textnormal{BH}}_{j}(1,1)=T^{\textnormal{swap}(i,j)}$. As a consequence,
-    $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\iff j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
-    and
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,1)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
-    As a result, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}>c_{i}$.
-Concluding the above two cases, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)$ when $y\leq c_{n+j}$.
-Case 2: $y>c_{n+j}$.
-We again prove ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$ for $y>c_{n+j}$ in two cases.
-  • $Y_{i}\leq c_{i}$.
-Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$.
-Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$,
-    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
-[…]
-    by the construction of $T^{\textnormal{swap}(i,j)}(y)$; therefore, $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)$.
-Meanwhile, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{i}\leq c_{i}\}=1$,
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}}\frac{m}{n+1}$
-[…]
-    following from the definition of $T^{\textnormal{BH}}_{j}(0,0)$.
-Therefore $T^{\textnormal{BH}}_{j}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$.
-As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,0)$ and
-    $\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
-    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
-    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0)$ and therefore
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
-[…]
-    which is a result of the definition of $T^{\textnormal{BH}}(0,0)$. The above also implies that $T^{\textnormal{BH}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Next,
-    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}\}}\frac{m}{n+1}$
-[…]
-    where the last step is a result of $T^{\textnormal{swap}(i,j)}(y)$’s definition.
-The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$.
-We then have
-    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}\in\mathfrak{S}.$
-    It follows from the definition that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
-Combining the above, we have shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)$ for any $i$ such that $Y_{i}\leq c_{i}$.
-  • $Y_{i}>c_{i}$. In this case, we have for any $t\in{\mathcal{T}}$ that
-    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
-[…]
-    By their definitions, there is $T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(0,1)$, and therefore
-    $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$
-    and
-    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,1)\}.$
-    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i$ such that $Y_{i}>c_{i}$.
-Combining the two cases, we have shown that ^$\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $y>c_{n+j}$. The proof is now complete.
-B.9 Proof of Proposition 8
-Recall that
-$\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(y)}\cup\{\infty\},1-\alpha\big{)}\Big{\}},$
-where ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ ia the reference set.
-We consider the following three cases: $S(X_{n+j},y)<\eta^{-}$, $S(X_{n+j},y)>\eta^{+}$ and $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$.
-The case where $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$ is obvious, so we discuess the remaining two cases in the following.
-As before, we define $\eta^{\textnormal{swap}(i,j)}(y)$ to be the $K$-th smallest elements in $\{\widehat{S}^{\textnormal{swap}(i,j)}_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}$.
-Case 1: $S(X_{n+j},y)<\eta^{-}$.
-Consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)<\eta^{-}$ and $y\in\widehat{C}_{\alpha,n+j}$. We claim that in this case
-^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta^{-},{\mathcal{L}}(X_{i},\eta^{-})=1\}.$
-To see this, note that when $S_{i}\leq\eta^{-}$, we have $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$.
-and therefore ${\mathcal{L}}(X_{i},\eta)={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$; when $S_{i}>\eta^{-}$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{-}$, and therefore ${\mathcal{L}}(X_{i},\eta^{-})={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$.
-As a result, we have that ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}$ and therefore
-$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(1)}}\cup\{\infty\},1-\alpha\big{)}$
-Case 2: $S(X_{n+j},y)>\eta^{+}$.
-We now consider $y\in{\mathcal{Y}}$ such that $S(X_{n+j},y)>\eta^{+}$ and $y\in\widehat{C}_{\alpha,n+j}$. In this case, we instead have
-^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(2)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta,{\mathcal{L}}(X_{i},\eta^{+})=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta,{\mathcal{L}}(X_{i},\eta)=1\}.$
-This is because when $S_{i}\leq\eta$, $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{+}$ and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta^{+})$; when $S_{i}>\eta$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$, and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta)$.
-Consequently,
-$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(2)}}\cup\{\infty\},1-\alpha\big{)}.$
-Combining the three cases, we finish the proof.
-Appendix C Auxiliary results
-C.1 The conformal p-values
-The conformal p-value and its random counterpart defined in Jin and Candès, 2023b are
-$\bar{p}_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}\text{ and }$
-$\bar{p}^{\textnormal{rand}}_{j}=\frac{U_{n+j}(1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}=\widehat{S}_{n+j}\})+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}>\widehat{S}_{n+j}\}}{n+1}.$
+$=\frac{\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{V(X_{n+j},Y_{n+j})>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}]}{\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{\mathcal{S}}\}]}$
+For the numerator, we have
+$\Big{|}\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}\big{]}-\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}^{*}\}\big{]}\Big{|}$
+$\leq\frac{1}{m}\mathbb{E}\big{[}|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|\big{]}=\mathbb{E}\bigg{[}\frac{|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|}{|\widehat{{\mathcal{S}}}^{*}|}\frac{|\widehat{{\mathcal{S}}}^{*}|}{m}\bigg{]}\to 0$
+where the last step is because $|\widehat{\mathcal{S}}\Delta\widehat{\mathcal{S}}^{*}|/|\widehat{\mathcal{S}}|{\to}0$ in probability and the dominated convergence theorem. Using (48), we then have
+$\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}\big{]}\to G(q^{*}).$
+Similarly, for the denominator,
+$\frac{1}{m}\Big{|}\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{\mathcal{S}}\}]-\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{{\mathcal{S}}}^{*}\}]\Big{|}\leq\frac{1}{m}\mathbb{E}\bigg{[}\frac{|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|}{|\widehat{{\mathcal{S}}}^{*}|}|\widehat{{\mathcal{S}}}^{*}|\bigg{]}\to 0.$
+Therefore, we obtain the same expression for the asymptotic selection-conditional coverage. ∎
+E.2 The conformal p-values
+Recall that in Section 4.2 of the main text, we introduce the conformal p-value for each $j\in[m]$ as
+$p_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1},$
+where $\widehat{S}_{i}=S(X_{i},c_{i})$, and $S\colon{\mathcal{X}}\times\mathbb{R}\to\mathbb{R}$ is the pre-specified score function.
+We can construct sharper p-values by introducing extra randomness.
+Letting $U_{n+1},U_{n+2},\ldots,U_{n+m}$ be i.i.d.
+uniform random variables on $[0,1]$ that are independent of everything else, we can construct the randomized conformal p-values as
+$p_{j}^{\textnormal{rand}}=\frac{U_{n+j}(1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}=\widehat{S}_{n+j},Y_{i}\leq c_{i}\}+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}>\widehat{S}_{n+j},Y_{i}\leq c_{i}\})}{n+1}.$
+The conformal p-values and their random counterparts defined in Jin and Candès, 2023b are
+$\bar{p}_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}\text{ and }$
+$\bar{p}^{\textnormal{rand}}_{j}=\frac{U_{n+j}(1+\sum_{i\in[n]}\mathds{1}\{S_{i}=\widehat{S}_{n+j}\})+\sum_{i\in[n]}\mathds{1}\{S_{i}>\widehat{S}_{n+j}\}}{n+1}.$
+The validity of $p_{j}$ and $p_{j}^{\textnormal{rand}}$, as well as their relation to $\bar{p}_{j}$ and $\bar{p}_{j}^{\textnormal{rand}}$, is established in the following lemma.
 Lemma 2.
 Under the condition of Theorem 1, for any $j\in[m]$, the conformal p-values $p_{j}$ and $p_{j}^{\textnormal{rand}}$ defined in (15) satisfy that
-$\mathbb{P}\big{(}p_{j}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\text{ and }\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,$
-for any $t\in[0,1]$.
-Furthermore, there is $p_{j}\leq\bar{p}_{j}$ and $p_{j}^{\textnormal{rand}}\leq\bar{p}_{j}^{\textnormal{rand}}$ deterministically, where the equalities hold for the nonformity score $S(x,y)=\widehat{\mu}(X)-\big{(}\sup_{x}2|\widehat{\mu}(x)|\big{)}\cdot y$.
-Proof.
-For any $j\in[m]$ and $t\in[0,1]$, it can be checked that when $Y_{n+j}\leq c_{n+j}$,
+$\mathbb{P}\big{(}p_{j}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\quad\text{and}\quad\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}\leq t,\quad\forall t\in[0,1].$
+Furthermore, there is $p_{j}\leq\bar{p}_{j}$ and $p_{j}^{\textnormal{rand}}\leq\bar{p}_{j}^{\textnormal{rand}}$ deterministically.
+If the selection score $S(x,y)=\widehat{\mu}(X)-\big{(}\sup_{x}2|\widehat{\mu}(x)|\big{)}\cdot y$, where $\widehat{\mu}$ is regression function fitted on ${\mathcal{D}}_{\textnormal{calib}}$, then the equalities hold.
+Proof of Lemma 2.
+For any $j\in[m]$ and $t\in(0,1)$, it can be checked that when $Y_{n+j}\leq c_{n+j}$,
 $p_{j}\leq t\iff\widehat{{\mathcal{S}}}_{n+j}\text{ is among the }k^{*}\text{-th largest element in }{\mathcal{N}}_{j}:=\{\widehat{S}_{i}:i\in{\mathcal{I}}_{j},Y_{i}\leq c_{i}\},$
@@
 Let $T_{k^{*}}$ denote the value of the $k^{*}$-th largest element in ${\mathcal{N}}_{j}$.
-Then apparently, both $k^{*}$ and $T_{k^{*}}$ are invariant to the permutation on ${\mathcal{I}}_{j}$.
+Apparently, both $k^{*}$ and $T_{k^{*}}$ are invariant to the permutation on ${\mathcal{I}}_{j}$.
 Consequently,
@@
 $E_{i}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}=\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\},\leavevmode\nobreak\ G_{i}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}\leq\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\},\leavevmode\nobreak\ G_{i}^{+}\,:=\,\sum_{\ell\in{\mathcal{I}}_{j}}\mathds{1}\{\widehat{S}_{i}<\widehat{S}_{\ell},Y_{\ell}\leq c_{\ell}\}.$
-Using these notations,
-$\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}=\mathbb{P}\big{(}U_{n+j}E_{n+j}+G_{n+j}^{+}\leq(n+1)t,Y_{n+j}\leq c_{n+j}\big{)}$ (49)
+Using these notations, we can write
+$\mathbb{P}\big{(}p_{j}^{\textnormal{rand}}\leq t,Y_{n+j}\leq c_{n+j}\big{)}=\mathbb{P}\big{(}U_{n+j}E_{n+j}+G_{n+j}^{+}\leq(n+1)t,Y_{n+j}\leq c_{n+j}\big{)}$
 […]
 […]
-where the second step follows from exchangeability.
+where the first step follows from exchangeability.
 Let $i^{*}$ denote the index of the largest element in $\{\widehat{S}_{i}:i\in{\mathcal{N}}_{j},G_{i}^{+}<(n+1)t\}$ (with ties broken arbitrarily).
 Then
-($=\leavevmode\nobreak$ $\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}|{\mathcal{N}}_{j}|+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot\Big{(}G_{i^{*}}^{+}+E_{i^{*}}\cdot\mathds{1}\Big{\{}U_{n+j}\leq\frac{(n+1)t-G_{i^{*}}^{+}}{E_{i^{*}}}\Big{\}}\Big{)}\Big{]}$
-$=\frac{1}{n+1}\mathbb{E}\Big{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}(n+1)t+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot(n+1)t\Big{]}$
-$\leq t.$
-By definition,
-$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1}$ $\overset{\rm(a)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{Y_{i}\leq c_{i},S_{i}\geq\widehat{S}_{n+j}\}}{n+1}$
-$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}=\bar{p}_{j}.$
+($=\leavevmode\nobreak$ $\frac{1}{n+1}\mathbb{E}\Bigg{[}\mathds{1}\big{\{}|{\mathcal{N}}_{j}|\leq(n+1)t\big{\}}\cdot|{\mathcal{N}}_{j}|+\mathds{1}\{|{\mathcal{N}}_{j}|>(n+1)t\}\cdot\bigg{(}G_{i^{*}}^{+}+E_{i^{*}}\cdot\mathds{1}\Big{\{}U_{n+j}\leq\frac{(n+1)t-G_{i^{*}}^{+}}{E_{i^{*}}}\Big{\}}\bigg{)}\Bigg{]}$
+[…]
+We have therefore proved the validity of $p_{j}$’s and $p_{j}^{\textnormal{rand}}$’s.
+Next, we have
+$p_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1}$ $\overset{\rm(a)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},S_{i}\geq\widehat{S}_{n+j}\}}{n+1}$
+$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}=\bar{p}_{j}.$
 Above, the step (a) is because when $Y_{i}\leq c_{i}$, $S_{i}\geq\widehat{S}_{i}$; when $S(x,y)=\widehat{\mu}(x)-(\sup_{x}2|\widehat{\mu}(x)|)\cdot\mathds{1}\{y>c\}$, the $\leq$” is replaced by “$=$” in step (a) and (b).
+A similar argument leads to that results corresponding to $p_{j}^{\textnormal{rand}}$.
 ∎
+E.3 Additional results on the conformal p-value-based procedures
+Below, Lemma 3 and Lemma 4 show that that fixed-threshold and BH-threshold selection rules are special instances of the general selection rule described in Section 4.2 of the main text.
+Lemma 3.
+Let $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:p_{j}\leq q\}$.
+Then $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ is equivalent to ${\widetilde{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:\widehat{{\mathcal{S}}}_{n+j}\geq T\}$, where
+$T=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}\geq q(n+1)\bigg{\}}.$
+Proof of Lemma 3.
+Suppose $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$, then $p_{j}\leq q$. By the definition of $p_{j}$,
+$\frac{1}{n+1}\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}=p_{j}\leq q.$
+By the definition of $T$, $\widehat{S}_{n+j}\geq T$, and therefore $j\in{\widetilde{\mathcal{S}}}_{\textnormal{fixed}}$. Conversely, if $\widehat{S}_{n+j}\geq T$, then
+$p_{j}=\frac{1}{n+1}\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}\leq\Big{(}1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq T,Y_{i}\leq c_{i}\}\Big{)}\leq q,$
+where the last step is by the definition of $T$.
+As a result, $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$.
+We then have $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}={\widetilde{\mathcal{S}}}_{\textnormal{fixed}}$.
+∎
+Lemma 4.
+Let $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ denote the selection set obtained by the BH procedure applied to the conformal p-values.
+Then $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ is equivalent to ${\widetilde{\mathcal{S}}}_{\textnormal{BH}}:=\{j\in[m]:\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$, where
+$T^{\textnormal{BH}}=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{1+\sum_{i\in[n]}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum_{j\in[m]}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{1+n}\leq q\bigg{\}}.$
+Proof of Lemma 4.
+Let $j^{*}=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$ and $\widehat{S}_{n+(1)},\widehat{S}_{n+(2)},\ldots,\widehat{S}_{n+(m)}$ denote the order statistics of $\{\widehat{S}_{n+j}\}_{j\in[m]}$ in a descending order.
+It then suffices to show that $j^{*}=k^{*}$.
+When $k^{*}\geq 1$,
+$\frac{1+\sum_{i\in[n]}\{\widehat{S}_{i}\geq\widehat{S}_{n+(k^{*})},Y_{i}\leq c_{i}\}}{(\sum_{j\in[m]}\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{n+(k^{*})}\})\vee 1}\frac{m}{1+n}=p_{(k^{*})}\frac{m}{k^{*}}\leq q,$ (50)
+where the last inequality follows from the definition of $k^{*}$.
+From the definition of $T^{\textnormal{BH}}$, we can conclude that $\widehat{S}_{n+(k^{*})}\geq T^{\textnormal{BH}}$; since $\widehat{S}_{n+(j^{*})}$ is the smallest element in $\{\widehat{S}_{n+j}\}_{j\in[m]}$ that is no less than $T^{\textnormal{BH}}$, we also have that $\widehat{S}_{n+(k^{*})}\geq\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$.
+Conversely,
+$p_{(j^{*})}\frac{m}{j^{*}}$ =$\overset{\rm(a)}{=}\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+(j^{*})},Y_{i}\leq c_{i}\}}{n+1}\cdot\frac{m}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}$
+$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}},Y_{i}\leq c_{i}\}}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}\cdot\frac{m}{n+1}\overset{\rm(c)}{\leq}q.$
+Above, step (a) uses the definition of $j^{*}$, step (b) is because $\widehat{S}_{n+(j^{*})}\geq T^{\textnormal{BH}}$ and step (c) is by the definition of $T^{\textnormal{BH}}$.
+From the form of the BH procedure, we have that $j^{*}\geq k^{*}$, and therefore $j^{*}=k^{*}$.
+When $k^{*}=0$, for any $t\in{\mathcal{T}}$, we let $j(t)=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\}$. If $j(t)>0$, then
+$\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{n+1}=\frac{1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{j(t)}\frac{m}{n+1}$
+[…]
+where the last step is because $k^{*}=0$. The above implies that $T^{\textnormal{BH}}>\widehat{S}_{n+(1)}\Rightarrow j^{*}=0$. The equivalence is therefore established. ∎
+E.4 Prediction set construction by further sample splitting
+In this section, we discuss the possibility of simplifying the prediction set construction by further splitting the calibration set.
+Since we are primarily interested in the case where we do not have the flexibiility to change the selection rule, it turns out that—in most of the cases—the sample-splitting idea does not help with finding the exchangeable reference set without substantially changing the nature of selection.
+There is one class of selection rules, however, that can benefit from the sample-splitting idea with only minor modifications to the original selection rule.
+To be concrete, consider the selection rule such that $\mathds{1}\{j\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}},X_{n+j})$ for some function $f$, i.e., whether $j$ is selected depends solely on the calibration set ${\mathcal{D}}_{\textnormal{calib}}$ and the feature $X_{n+j}$.
+Examples of such selection rules include thresholding conformal p-values with a fixed threshold, and selecting units based on properties of the preliminary prediction set.
+For such selection rules, we could split the calibration set into two parts, ${\mathcal{D}}_{\textnormal{calib}}={\mathcal{D}}_{\textnormal{calib}}^{(1)}\cup{\mathcal{D}}_{\textnormal{calib}}^{(2)}$.
+If we allow ourselves to determine the selection set via $\mathds{1}\{j\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}}^{(1)},X_{n+j})$, then we can construct the prediction set as
+$\begin{split}&\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}:V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}}\cup\{\infty\},1-\alpha\big{)}\Big{\}},\\ &\text{where }\widehat{{\mathcal{R}}}:=\{Z_{i}\in{\mathcal{D}}_{\textnormal{calib}}^{(2)}:f({\mathcal{D}}_{\textnormal{calib}}^{(1)},X_{i})=1\}.\end{split}$ (51)
+The following proposition shows that the prediction set constructed in (51) is a valid $(1-\alpha)$ selection-conditional prediction set.
+Proposition 8.
+Under the condition of Theorem 1, and suppose that for $\forall j\in[m]$, the selection rule is such that $\mathds{1}\{\ell\in\widehat{\mathcal{S}}\}=f({\mathcal{D}}_{\textnormal{calib}},X_{n+\ell})$ for some function $f$.
+Then the prediction set $\widehat{C}_{\alpha,n+j}$ constructed in (51) is a valid $(1-\alpha)$ prediction set.
+Proof.
+Treating ${\mathcal{D}}_{\textnormal{calib}}^{(2)}$ as the calibration set in Theorem 1, it is straight forward to check that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}$, and therefore by Theorem 1, the prediction set $\widehat{C}_{\alpha,n+j}$ is a valid $(1-\alpha)$ prediction set.
+∎
+E.5 Analysis of calibration sample size
+A practical concern of our proposed method is that accounting for the selection event may reduce the number of calibration samples, which has a direct impact on the size and stability of the prediction set.
+The degree of effect is high dependent on the selection rule, which determines the fundamental difficulty of the problem: for example, if the selection rule picks out only the test unit with the largest predicted value, then conditional inference essentially becomes inference on the tail of a distribution; it is not surprising to see that only a limited number of calibration samples can be obtained.
+In this section, we provide an asymptotic analysis of a class of selection rules that are relatively “mild”.
+To be concrete, we consider the selection rule yields a reference set in the form of $\{i\in[n]:h(X_{i})\geq\widehat{q}\}$, where $h$ is a function of the features and $\widehat{q}$ is a threshold that potentially dependent on $({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ but converges to a constant as $n,m$ goes to infinity.
+The following proposition characterizes the size of such reference set asymptotically.
+Proposition 9.
+Suppose the reference set ^=$\widehat{{\mathcal{R}}}=\{i\in[n]:h(X_{i})\geq\widehat{q}\}$, where $h$ is a fixed function of the features and $\widehat{q}$ is a data-dependent threshold such that $\lim_{n,m\rightarrow}\widehat{q}=q^{*}$ a.s.
+For $\forall t\in\mathbb{R}$, let $G(t)=\mathbb{P}(h(X)\geq t)$.
+Assume that $G(t)$ is continuous in $t$.
+Then the size of the reference set satisfies that
+[…]
+Proof.
+By definition, we can write |^$|\widehat{{\mathcal{R}}}|=\sum_{i=1}^{n}\mathds{1}\{h(X_{i})\geq\widehat{q}\}$. By the strong law of large numbers, we have that
+[…]
+Since $\lim_{n,m\rightarrow\infty}\widehat{q}=q^{*}$, and $G(t)$ is continuous in $t$, we have by the continuous mapping theorem that $\lim_{n,m\rightarrow\infty}|G(\widehat{q})-G(q^{*})|=0$.
+By triangular inequality, we complete the proof.
+∎
+Many selection rules we have considered in this paper fall into the class.
+For example, for the quantile based selection rule, the reference set is $\{i\in[n]:S(X_{i})\geq\widehat{q}\}$, where $\widehat{q}$ is either the joint quantile or the calibration quantile, which converges to the popuplation quantile as $n,m$ goes to infinity.
+The class of conformal p-value-based selection rules considered in Section 4.2 can also be characterized by the proposition.
+To see this, consider the fixed threshold selection rule or the BH-threshold selection rule.
+For simplicity suppose that $y\leq c_{n+j}$ and $\mathfrak{S}=2^{[m]}$, then
+^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in[n]:Y_{i}\leq c_{i},\widehat{S}_{i}\geq\tau^{(j)}(1,0)\big{\}}\cup\big{\{}i\in[n]:Y_{i}>c_{i},\widehat{S}_{i}\geq\tau^{(j)}(1,1)\big{\}}.$
+Above, $\tau^{(j)}(1,0)$ and $\tau^{(j)}(1,1)$ converge to the same constant for both variants (c.f. Jin and Candès, 2023b ).